### v1

# Hyperfocal subalgebras of hyperfocal abelian Frobenius blocks

Hyperfocal subalgebras of hyperfocal abelian Frobenius blocks
Xueqin Hu1, Kun Zhang2 and Yuanyang Zhou1
($1.$ School of Mathematics and Statistics, Central China Normal University, Wuhan, 430079, China
$2.$ Faculty of Mathematics and Statistics, Hubei University, Wuhan, 430062, China
[…]
Mathematics Subject Classification 20C15; 20C20
1 Introduction
In this paper, we always assume that $p$ is a prime. Let $(\mathcal{K},\mbox{$\mathcal{O}$},k)$ be a complete $p$-modular system consisting of a complete discrete valuation ring $\mathcal{O}$ with a residue field $k$ of prime characteristic $p$ and a quotient field $\mathcal{K}$ of characteristic 0. Moreover, we always assume that the residue field $k$ is algebraically closed and the quotient field $\mathcal{K}$ is large enough for all finite groups considered below.
[…]
In the light of Rouquier’s conjecture, some recent work focuses on the investigation of the structure of the block algebra $\mbox{$\mathcal{O}$}Gb$ with some special abelian hyperfocal subgroup. In [33] and [7], the numbers of ordinary irreducible characters and irreducible Brauer characters were obtained for the blocks with cyclic and Klein four hyperfocal subgroup, respectively. In [8], both authors constructed an isotypy, which is a character-theoretic shadow of Rouquier’s conjecture, between the blocks $b$ and $c$ when the hyperfocal subgroup $D$ is cyclic and the quotient group $P/D$ is abelian. These results only reflect the influences of the structure of hyperfocal subgroups on the structure of characters of block algebras. It may be more interesting and meaningful if the algebraic structure of the block algebras with some special abelian hyperfocal subgroups can be determined.
[…]
Motivated by this, in this paper, we will focus on the structure of the hyperfocal subalgebra of a block with a special class of hyperfocal subgroups, which includes the two classical cases where the hyperfocal subgroup is a Klein four group and cyclic, respectively. To make this explicit, we need some notation. Let $\mathbb{A}$ be a hyperfocal subalgebra of the block $b$. Let $D_{\delta}$ be a local pointed group of $D$ on the block algebra $\mbox{$\mathcal{O}$}Gb$ and denote by $N_{G}(D_{\gamma})$ its stabilizer under the conjugation action of $G$ (see the paragraph 2.9 below). As it is shown in Proposition 3.7 below, the quotient group $N_{G}(D_{\gamma})/DC_{G}(D)$ is $p$-nilpotent. Its normal $p$-complement is denoted by $E_{\mathfrak{hyp}}$. Since the order of $E_{\mathfrak{hyp}}$ is coprime to $p$, we can view it as a subgroup of the automorphism group $\mathrm{Aut}(D)$ of $D$. Then the block $b$ is called with Frobenius hyperfocal inertial quotient if $E_{\mathfrak{hyp}}$ acts freely on $D-\{1\}$, namely the set of all nontrivial elements of $D$ (see the paragraph 3.8 below). This is equivalent to requiring that the semidirect product $D\rtimes E_{\mathfrak{hyp}}$ is a Frobenius group. Now we can state the main result of this paper as follows.
Theorem 1.1. (see also Theorem 5.14) Keep the notation as above. Suppose that the hyperfocal subgroup $D$ is abelian and the block $b$ is with Frobenius hyperfocal inertial quotient. Then there is a stable equivalence of Morita type between the hyperfocal subalgebra $\mathbb{A}$ and the group algebra $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{hyp}})$.
Throughout this paper, a block with abelian hyperfocal subgroup and Frobenius hyperfocal inertial quotient is called a hyperfocal abelian Frobenius block for simplicity. By definition, the hyperfocal abelian Frobenius block is an analogous version of the block with abelian defect group and Frobenius inertial quotient, which is called an abelian Frobenius block here for simplicity, at the hyperfocal level. So the main theorem above can be viewed as a hyperfocal version of [17, Theorem 10.5.1], which is due to Puig (see [23, 6.8]). Just like abelian Frobenius blocks including the blocks with Klein four defect groups and cyclic defect groups, the hyperfocal abelian Frobenius blocks include the blocks with Klein four hyperfocal subgroups and cyclic hyperfocal subgroups (see Remark 3.9 below). Thus we can apply this main theorem above to the blocks with Klein four hyperfocal subgroups and cyclic hyperfocal subgroups and obtain the structures of the corresponding hyperfocal subalgebras as follows, which are similar to the structures of the blocks with Klein four defect groups and cyclic defect groups, respectively.
[…]
Theorem 1.3. (see also Theorem 6.20) Assume that the hyperfocal subgroup $D$ is a cyclic group. Then the hyperfocal subalgebra $\mathbb{A}$ is Rickard equivalent to the group algebra $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{hyp}})$.
Theorem 1.3 is a generalization of [9, Theorem 1.2] since we don’t need the defect group being abelian and then gives a complete description of the structure of the hyperfocal subalgebra with cyclic hyperfocal subgroup. This result also gives a positive answer to a question whether the hyperfocal subalgebra with nontrivial cyclic hyperfocal subgroup is a Brauer tree algebra proposed by Professor Linckelmann some years ago.
Furthermore, by using Clifford theoretic relationship between the representation theory of the source algebra and its hyperfocal subalgebra, we can also obtain some information about the characters of the blocks with Klein four hyperfocal subgroups and cyclic hyperfocal subgroups in terms of source algebras. We refer to Propositions 6.15 and 6.34 for detailed descriptions of these structures. A key step towards these descriptions is to determine the relationship between the dimenisons of simple $\mathcal{K}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}$-modules and the prime number $p$. At the same time, this relates a conjecture, denoted by KLN cojecture for simplicity, proposed by Kessar and Linckelmann and Navarro in [13]. We recall this conjecture in the following.
Denote by $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})$ a set of representatives of isomorphism classes of simple $\mathcal{K}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}$-modules and by $\mathrm{Irr}_{\mathcal{K},0}(\mathbb{A})$ the subset corresponding to simple modules of dimensions coprime to $p$. In [13], Kessar and Linckelmann and Navarro proposed the following conjecture.
KLN Conjecture. ([13, Conjecture 1.2]) With the notation above, assume that $\mathcal{K}$ is a splitting field of $\mathbb{A}$. Then $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})=\mathrm{Irr}_{\mathcal{K},0}(\mathbb{A})$ if and only if the hyperfocal subgroup $D$ is abelian.
In [13], the KLN conjecture is viewed as a ‘hyperfocal height zero’ version of Brauer’s height zero conjecture. In this paper, we can verify the forward direction of the KLN conjecture for the cases where the hyperfocal subgroup $D$ is a Klein four group and cyclic.
Proposition 1.4. (see also Proposition 6.9) When the hyperfocal subgroup $D$ is a Klein four group, then $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})=\mathrm{Irr}_{\mathcal{K},0}(\mathbb{A})$.
Proposition 1.5. (see also Proposition 6.28) When the hyperfocal subgroup $D$ is cyclic, then $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})=\mathrm{Irr}_{\mathcal{K},0}(\mathbb{A})$.
Now let us illustrate the construction of the stable equivalence of Morita type in Theorem 1.1 briefly. This construction is similar to the one in [17, Theorem 10.5.1] for abelian Frobenius blocks. The main ingredients used for the construction in [17, Theorem 10.5.1] are the descriptions of local structures of an abelian Frobenius block and Puig’s Gluing Theorem for indecomposable endopermutation modules of abelian $p$-groups. Let us make it more explicit. For an abelian Frobenius block with defect group $P$, its Brauer correspondents in the centralizers of nontrivial subgroups of $P$ are nilpotent in the sense of [2]. By the structure theorem of the source algebras of nilpotent blocks, due to Puig [22], these local structures can yield a family of indecomposable endopermutation modules (see the paragraph 2.14 below) of quotient groups of the abelian $p$-group $P$ satisfying some compatibility assumptions. Then by Puig’s Gluing Theorem, these endopermutation modules can be glued together to an indecomposable endopermutation module of $P$, which plays an essential role in the construction of the stable equivalence of Morita type for abelian Frobenius blocks.
Analogously, this method can be carried over to the hyperfocal algebra of the hyperfocal abelian Frobenius block. Instead of describing the structures of the Brauer correspondents in the centralizers of nontrivial subgroups of the defect group $P$, we need to determine the structures of the Brauer correspondents in the centralizers of nontrivial subgroups of the hyperfocal subgroup $D$, which are shown to be also nilpotent by Proposition 3.12 below. Naturally, the next step should be applying the structure theorem of the source algebras of nilpotent blocks to these blocks to yield a family of indecomposable endopermutation modules of quotient groups of $D$. However, there is an obstruction to this application because the Brauer quotients of the source algebra at nontrivial subgroups of $D$ may not be a source algebra of the block of the corresponding centralizer when the defect group $P$ is nonabelian. Due to this obstruction, instead of considering the source algebra and the hyperfocal subalgebra, we will focus on a so-called embedded algebra $\mathbb{A}_{\delta}$ of the pointed group $D_{\delta}$ of the hyperfocal subalgebra $\mathbb{A}$ (see the paragraph 2.10 below), which can be viewed as a subalgebra (not unitary in general) of the hyperfocal subalgebra $\mathbb{A}$. The embedded algebra $\mathbb{A}_{\delta}$ has some properties similar to the source algebra, such as the existence of a Morita equivalence between $\mathbb{A}_{\delta}$ and the hyperfocal subalgebra $\mathbb{A}$ (see Corollary 3.4 below) and an $\mbox{$\mathcal{O}$}(D\times D)$-module structure similar to the $\mbox{$\mathcal{O}$}(P\times P)$-module structure of the source algebra (see Proposition 4.6 below). These properties will play essential roles in the descriptions of the structures of the Brauer quotients of $\mathbb{A}_{\delta}$ at nontrivial subgroups of $D$ and eventually the construction of the stable equivalence of Morita type we need.
The paper is organized as follows. In Section $2$, some necessary notation and definitions are collected. In Section $3$, the definition of the hyperfocal abelian Frobenius block is introduced after we show that the quotient group $N_{G}(D_{\delta})/DC_{G}(D)$ is $p$-nilpotent. Then the local structures of this block associated with nontrivial subgroups of the hyperfocal subgroup are determined. Also, some basic properties of the embedded algebra $\mathbb{A}_{\delta}$ are proved in this section. Section $4$ is devoted to determining the structures of the embedded algebra $\mathbb{A}_{\delta}$, including its structure as an $\mbox{$\mathcal{O}$}(D\times D)$-module and the algebraic structure of its Brauer quotient at every nontrivial subgroup of the hyperfocal subgroup, which are essential to the construction of a stable equivalence of Morita type done in Section $5$. Besides this construction, in Section $5$, we also investigate an equivariant property of the bimodule inducing this stable equivalence, which is necessary when we want to obtain the information about the source algebra. In Section $6$, the structures of the hyperfocal algebras of blocks with Klein four hyperfocal subgroups and cyclic hyperfocal subgroups are determined as applications of the main result obtained in Section $5$. Some information about characters of these two blocks is also obtained in Section $6$.
2 Preliminaries
In this section, we collect some necessary notation and definitions.
2.1.
Let $S$ be a finite set. We denote by $|S|$ the cardinality of $S$. Let $H$ be a finite group and $\mathfrak{b}$ a block of $H$ over $\mathcal{O}$. Denote by $\mathrm{Irr}(H)$ and $\mathrm{Irr}(\mathfrak{b})$ the set of ordinary irreducible characters of $H$ and the set of ordinary irreducible characters of $\mathfrak{b}$, respectively. Similary, we denote by $\mathrm{IBr}(H)$ and $\mathrm{IBr}(\mathfrak{b})$ the set of modular irreducible characters of $H$ and the set of modular irreducible characters of $\mathfrak{b}$, respectively. If no confusion arises, we denote by $1_{H}$ the trivial character of $H$. Let $H_{1}$ and $H_{2}$ be two subgroups of $H$. For any $\xi\in\mathrm{Irr}(H_{1})$, we denote by $\mathrm{Ind}_{H_{1}}^{H}(\xi)$ the induced character of $\xi$. We use the similar notation for the induced module. Analogously, we denote the restriction by $\mathrm{Res}_{H_{1}}^{H}$. For any $\mathcal{O}H_{1}$-module $M$ and any $h\in H$, we define the twisted $\mathcal{O}(hH_{1}h^{-1})$-module $_{h}M$ by setting $_{h}M=M$ as an $\mathcal{O}$-module, with $hh_{1}h^{-1}\in hH_{1}h^{-1}$ acting on $m\in{{}_{h}M}$ as $h_{1}$ on $m\in M$. Furthermore, if $M$ is an $\mathcal{O}H_{1}$-$\mathcal{O}H_{2}$-bimodule, then by preserving the right $\mathcal{O}H_{2}$-module structure, $_{h}M$ is an $\mathcal{O}(hH_{1}h^{-1})$-$\mathcal{O}H_{2}$-bimodule.
$2.$
Throughout this paper, an $\mathcal{O}$-algebra always satisfies the property that it is free and has finite rank as an $\mathcal{O}$-module. An $\mathcal{O}$-algebra $C$ is called an indecomposable algebra if it is indecomposable as a $C$-$C$-bimodule, which is equivalent to saying that its center $Z(C)$ is local. Let $\tilde{C}$ be a subalgebra of $C$. We denote by $\mathrm{Res}_{\tilde{C}}(M)$ the restriction of $M$ to $\tilde{C}$ for any left $C$-module $M$. Denote by $1_{C}$ (or 1 if no confusion arises), $J(C)$ and $C^{*}$ the unit element of $C$, the Jacobson radical of $C$, the set of invertible elements of $C$, respectively. We denote by $C^{\circ}$ the opposite algebra of $C$. Let $C^{\prime}$ be another $\mathcal{O}$-algebra. Let $\varphi$ be an isomorphism from $C$ to $C^{\prime}$ and $M$ a left $C$-module. We use ${{}_{\varphi}}M$ to denote the left $C^{\prime}$-module with being equal to $M$ as $\mathcal{O}$-modules and the left action of $c^{\prime}\in C^{\prime}$ on $_{\varphi}C$ being equal to the left action of $\varphi^{-1}(c^{\prime})$ on $M$. For any $CC^{\prime}-$bimodule $M$, we can define $(x\mathop{\otimes}\limits_{\mathcal{O}}x^{\prime})m=xmx^{\prime}$ for any $x\in C$, $x^{\prime}\in(C^{\prime})^{\circ}$ and $m\in M$. In this way, it can be viewed as a $C\mathop{\otimes}\limits_{\mathcal{O}}(C^{\prime})^{\circ}$-module and vice versa. Throughout this paper, we identify $CC^{\prime}-$bimodules with $C\mathop{\otimes}\limits_{\mathcal{O}}(C^{\prime})^{\circ}$-modules through this way and vice versa.
2.3.
We denote the Heller operator of $C$ by $\Omega_{C}$ for an $\mathcal{O}$-algebra $C$. This operator can induce an equivalence on the $\mathcal{O}$-stable category of $\mathrm{mod}(C)$, of which the definition we recall as follows. Recall that a $C$-module $U$ is called relative $\mathcal{O}$-projective, if it is isomorphic to a direct summand of $C\mathop{\otimes}\limits_{\mathcal{O}}V$ for some $\mathcal{O}$-module $V$ (see [16, Definition 2.6.11]). Note that if the $C$-module $U$ is relative $\mathcal{O}$-projective and free of finite rank, then $U$ is a projective $C$-module. Denote by $\mathrm{mod}(C)$ the category of finitely generated left $C$-modules. The quotient category of $\mathrm{mod}(C)$ by the subcategory of relatively $\mathcal{O}$-projective modules is called $\mathcal{O}$-stable category of $\mathrm{mod}(C)$ (see [14, 11.1]). We denote it by $\overline{\mathrm{mod}}(C)$.
2.4.
Let us recall the definitions of the stable equivalence of Morita type and Rickard equivalence. Let $C$ and $C^{\prime}$ be two $\mathcal{O}$-algebras. A $CC^{\prime}-$bimodule $M$ and a $^{\prime}-CC$-bimodule $M^{\prime}$ are said to induce a stable equivalence of Morita type between $C$ and $C^{\prime}$ ([1]), if $M$ and $M^{\prime}$ are projective both as left and right modules, and if $M\mathop{\otimes}\limits_{C^{\prime}}M^{\prime}\cong C\oplus X$ as $C$-$C$-bimodules and $M^{\prime}\mathop{\otimes}\limits_{C}M\cong C^{\prime}\oplus X^{\prime}$ as $^{\prime}-CC^{\prime}$-bimodules, where $X$ is a projective $C$-$C$-bimodule and $X^{\prime}$ is a projective $^{\prime}-CC^{\prime}$-bimodule. In this case, the functors $M\mathop{\otimes}\limits_{C^{\prime}}$- and $M^{\prime}\mathop{\otimes}\limits_{C}$- induce mutually inverse equivalences between the $\mathcal{O}$-stable categories $\overline{\mathrm{mod}}(C^{\prime})$ and $\overline{\mathrm{mod}}(C)$.
2.5.
Keep the notation above. Denote $\mathrm{Hom}_{\mathcal{O}}(C,\mathcal{O})$ by $C^{*}$. It is the dual module of the $C$-$C$-bimodule $C$. The $\mathcal{O}$-algebra $C$ is called symmetric over $\mathcal{O}$ if $C$ is finitely generated projective as an $\mathcal{O}$-module and $C\cong C^{*}$ as $C$-$C$-bimodules (see [16, Definition 2.11.1]). Take two symmetric $\mathcal{O}$-algebras $C$ and $C^{\prime}$. Let $M_{\cdot}$ be a bounded complex of finitely generated $CC^{\prime}-$bimodules which are projective as left $C$-modules and right $C^{\prime}$-modules. Set $M_{\cdot}^{*}$ to be the dual $\mathrm{Hom}_{\mathcal{O}}(M_{\cdot},\mathcal{O})$ of the complex $M_{\cdot}$. We say that the complex $M_{\cdot}$ induces a Rickard equivalence between $C^{\prime}$ and $C$, if there is a $CC^{\prime}-$bimodule $F$ and a $^{\prime}-CC$-bimodule $F^{\prime}$ such that $F$ and $F^{\prime}$ are homotopy equivalent to 0 as complexes of bimodules, and if $M_{\cdot}\mathop{\otimes}\limits_{C^{\prime}}M_{\cdot}^{*}\cong C\oplus F$ as complexes of $C$-$C$-bimodules and $M_{\cdot}^{*}\mathop{\otimes}\limits_{C}M_{\cdot}\cong C^{\prime}\oplus F^{\prime}$ as complexes of $^{\prime}-CC^{\prime}$-bimodules (see [27, Definitions 2.2.1]). The complex $M_{\cdot}$ is called a Rickard complex. It is clear that if $C$ and $C^{\prime}$ are Rickard equivalent, then they are derived equivalent, namely their bounded derived categories $\mathcal{D}^{b}(C)$ and $\mathcal{D}^{b}(C^{\prime})$ are equivalent to each other as triangulated categories.
In the following two paragraphs, we will collect some notation from [12, §2-§3] which will be used later in the descriptions of characters of the source algebras and the hyperfocal subalgebras in Section $6$. Throughout the following $2$ paragraphs, we fix a symmetric $\mathcal{O}$-algebra $C$ such that $\mathcal{K}\mathop{\otimes}_{\mathcal{O}}C$ is a split semi-simple $\mathcal{K}$-algebra, which means that $\mathcal{K}\mathop{\otimes}_{\mathcal{O}}C$ is isomorphic to a direct product of matrix algebras over $\mathcal{K}$.
2.6.
Denote $\mbox{$\mathcal{K}$}\mathop{\otimes}\limits_{\mathcal{O}}C$ and $k\mathop{\otimes}\limits_{\mathcal{O}}C$ by $\hat{C}$ and $\bar{C}$, respectively. We have two Grothendieck groups $R_{\mathcal{K}}(C)$ and $R_{k}(C)$ associated with these two algebras $\hat{C}$ and $\bar{C}$, respectively. This means that $R_{\mathcal{K}}(C)$ and $R_{k}(C)$ are the Grothendieck groups of finite dimensional $\hat{C}-$modules and $\bar{C}$-modules, respectively. Furthermore, denote by $\mathrm{Pr}_{\mathcal{O}}(C)$ the subgroup of $R_{\mathcal{K}}(C)$ generated by the images of modules of the form $\hat{C}\mathfrak{i}$ for some idempotent $\mathfrak{i}$ of $C$. For any finite dimensional module $U$ over $\hat{C}-$modules or $\bar{C}$-modules, we denote by $[U]$ the image of $U$ in $R_{\mathcal{K}}(C)$ or $R_{k}(C)$, respectively. It is clear that the set of images of simple $\hat{C}-$modules, denoted by $\mathrm{Irr}_{\mathcal{K}}(C)$, is a $\mathbb{Z}$-basis of $R_{\mathcal{K}}(C)$. Set $\mathrm{k}(C)=|\mathrm{Irr}_{\mathcal{K}}(C)|$. Similarly, the set of images of simple $\bar{C}$-modules, denoted by $\mathrm{Irr}_{k}(C)$, is a $\mathbb{Z}$-basis of $R_{k}(C)$. Set $l(C)=|\mathrm{Irr}_{k}(C)|$.
2.7.
Furthermore, for any finite dimensional $\bar{C}$-module $X$, there exists a finitely generated $C$-module $Y$ which is free of finite rank as an $\mathcal{O}$-module such that $\mathcal{K}\mathop{\otimes}\limits_{\mathcal{O}}Y$ is isomorphic to $X$. This induces the so-called decomposition map $d_{C}$ from $R_{\mathcal{K}}(C)$ to $R_{k}(C)$, sending $[X]$ to $k\mathop{\otimes}\limits_{\mathcal{O}}Y]$. Its kernel is denoted by $L^{0}(C)$. The nonnegative integers occurring in the matrix of $d_{C}$ with respect to the canonical bases $\mathrm{Irr}_{\mathcal{K}}(C)$ and $\mathrm{Irr}_{k}(C)$ are called decomposition numbers. It is well known that there is a bilinear form $\langle,\rangle_{C}$ on $R_{\mathcal{K}}(C)$ such that $\langle[U],[V]\rangle_{C}=\mathrm{dim}_{\mathcal{K}}(U,V)$ for any two finite dimensional $\hat{C}-$modules $U$ and $V$. Set $L^{0}(C)^{\bot}$ to be the subgroup of $R_{\mathcal{K}}(\mathbb{A})$ consisting of elements orthogonal to every elements of $L^{\circ}(C)$ with respect to this bilinear form.
Next, we will recall some terminology and results about pointed groups which are due to Puig. We refer to [19, 20, 21] for details. With these notation, we will give the definitions of hyperfocal subgroups and hyperfocal subalgebras introduced by Puig.
2.8.
We fix a finite group $H$ in the remaining paragraphs of this section. An $\mathcal{O}$-algebra $\mathbb{B}$ is called an $H$-algebra if there is a group homomorphism $H\longrightarrow\mathrm{Aut}_{\mathcal{O}}(\mathbb{B})$. If this group homomorphism is induced by a group homomorphism $\theta:H\longrightarrow\mathbb{B}^{*}$, the $H$-algebra $\mathbb{B}$ is called an $H$-interior. Let $\mathbb{B}$ be an $H$-interior algebra. Clearly, $\mathbb{B}$ becomes an $\mbox{$\mathcal{O}$}H$-$\mbox{$\mathcal{O}$}H$-bimodule via this group homomorphism. We simply write $\theta(h^{-1})(a)\theta(h)$ as $a^{h}$ and $\varphi(x)a\varphi(y)$ as xayxay for any $h,x,y\in H$ any $a\in\mathbb{B}$. Let $\mathbb{B}^{\prime}$ be another $H$-interior algebra. An $\mathcal{O}$-algebra homomorphism $f:\mathbb{B}\rightarrow\mathbb{B}^{\prime}$ is called an $H$-interior algebra homomorphism if $f$ preserves the $H$-interior algebra structures on $\mathbb{B}$ and $\mathbb{B}^{\prime}$. Here, $f$ needs not be unitary. We say it is an embedding if $\text{Ker}(f)=\{0\}$ and $\text{Im}(f)=f(1)\mathbb{B}^{\prime}f(1)$. For any subgroup $Z$ of $H$, denote by $\mathbb{B}^{Z}$ the subalgebra of all elements of $\mathbb{B}$ fixed by the $Z$-conjugation. The $H$-interior algebra $\mathbb{B}$ is called primitive if $\mathbb{B}^{H}$ is a local $\mathcal{O}$-algebra.
2.9.
A point $\alpha$ of $Z$ on $\mathbb{B}$ is a $(\mathbb{B}^{Z})^{*}$-conjugacy class of primitive idempotents in $\mathbb{B}^{Z}$. The pair $(Z,\alpha)$, denoted by $Z_{\alpha}$, is called a pointed group on $\mathbb{B}$. We call a $Z$-interior algebra $\mathbb{B}_{\alpha}$ an embedded algebra of $Z_{\alpha}$ if there is an embedding of $Z$-interior algebras $f_{\alpha}:\mathbb{B}_{\alpha}\longrightarrow\mathbb{B}$ such that $f_{\alpha}(1_{\mathbb{B}_{\alpha}})\in\alpha$. By [20, 1.6], embedded algebras of $Z_{\alpha}$ are unique up to isomorphism. A typical choice of embedded algebras of $Z_{\alpha}$ is $a\mathbb{B}a$ and then the embedding $f_{\alpha}:a\mathbb{B}a\longrightarrow\mathbb{B}$ is just induced by inclusion. Here, $a$ is an element of $\alpha$ and $\mathbb{B}_{\alpha}$ is a $Z$-interior algebra with the group homomorphism $Z\rightarrow(\mathbb{B}_{\alpha})^{*},z\mapsto za$. We denote its image by $Za$. For two pointed groups $Z_{\alpha}$ and $Z^{\prime}_{\alpha^{\prime}}$ on $\mathbb{B}$, we say that $Z_{\alpha}$ is contained in $Z^{\prime}_{\alpha^{\prime}}$ if $Z\leq Z^{\prime}$ and there are $e\in\alpha$ and $e^{\prime}\in\alpha^{\prime}$ such that $ee^{\prime}=e=e^{\prime}e$. It is clear that the subgroup $N_{H}(Z)$ acts on the set $\{Z_{\alpha}~|~\alpha~\mathrm{runs~over~all~points~of~}Z~\mathrm{on~}\mathbb{B}\}$ by conjugation. Denote by $N_{H}(Z_{\alpha})$ the stabilizer of the pointed group $Z_{\alpha}$ under this action.
2.10.
For any subgroup $K$ of $Z$, we have the relative trace map $\mathrm{Tr}_{K}^{Z}:\mathbb{B}^{K}\rightarrow\mathbb{B}^{Z},a\mapsto\sum\limits_{z}a^{z}$, where $z$ runs over a set of representatives of the right cosets of $K$ in $Z$. Denote by $\mathbb{B}_{K}^{Z}$ the image of the map $\mathrm{Tr}_{K}^{Z}$. The sum $\sum\limits_{K<Z}\mathbb{B}_{K}^{Z}+J(\mathcal{O})\mathbb{B}^{Z}$ is an ideal of $\mathbb{B}^{Z}$. Denote by $\mathbb{B}(Z)$ the quotient of $\mathbb{B}^{Z}$ by this ideal and by $\mathrm{Br}_{Z}^{\mathbb{B}}$ the canonical homomorphism $\mathbb{B}^{Z}\rightarrow\mathbb{B}(Z)$. The quotient $\mathbb{B}(Z)$ is called the Brauer quotient of $\mathbb{B}$ at $Z$ and the associated homomorphism $\mathrm{Br}_{Z}^{\mathbb{B}}$ is called the Brauer homomorphism. If no confusion arises, we denote $\mathrm{Br}_{Z}^{\mathbb{B}}$ by $\mathrm{Br}_{Z}$. By [29, Lemma 11.7], $\mathbb{B}(Z)\neq 0$ only if $Z$ is a $p$-group. When $\mathbb{B}$ is the group algebra $\mathcal{O}L$ for some finite group $L$, the inclusion $\mathcal{O}C_{L}(Z)\subseteq(\mathcal{O}L)^{Z}$ induces a $k$-algebra isomorphism $(\mathcal{O}L)(Z)\cong kC_{L}(Z)$. We identify these two $k$-algebras through the isomorphism. More generally, there is a similar construction for any $\mathcal{O}Z$-module $M$. Set $M^{K}=\{m\in M\mid x\cdot m=m,\forall x\in M\}$. We can define a relative trace map $\mathrm{Tr}_{K}^{Z}:M^{K}\rightarrow M^{Z}$ by a similar way and denote its image by $M_{K}^{Z}$. Similarly, we set $M(Z)$ to be the quotient $\mathcal{O}$-module $M^{Z}/(\sum\limits_{K<Z}M_{K}^{Z}+J(\mathcal{O})M^{Z})$ and also denote the canonical homomorphism $M^{Z}\rightarrow M(Z)$ by $\mathrm{Br}_{Z}^{M}$, or $\mathrm{Br}_{Z}$ for simplicity if no confusion arises.
2.11.
Recall that a pointed group $R_{\gamma}$ on $\mathbb{B}$ is local, if $\mathrm{Br}_{R}(\gamma)\neq 0$. For any local pointed group $R_{\gamma}$, it is clear that there is a unique primitive idempotent $e$ of $Z(\mathbb{B}(R))$ such that $e\cdot\mathrm{Br}_{R}(\gamma)=\mathrm{Br}_{R}(\gamma)$. We call this primitive idempotent $e$ is associated with $R_{\gamma}$. A pointed group $R_{\gamma}$ on $\mathbb{B}$ is a defect pointed group of a pointed group $Z_{\alpha}$, if $R_{\gamma}$ is a maximal local pointed group contained in $Z_{\alpha}$. Let $R_{\gamma}$ be a defect pointed group of $K_{\beta}$ on $\mathbb{B}$. Then the embedded algebra $\mathbb{B}_{\gamma}$ of $R_{\gamma}$ is called a source algebra of the embedded algebra $\mathbb{B}_{\beta}$ of $K_{\beta}$ (see [19]). Note that source algebras of the embedded algebra $\mathbb{B}_{\beta}$ of $K_{\beta}$ are unique up to $K$-conjugations and isomorphisms (see [29, §18]).
2.12.
Let $R_{\gamma}$ be a local point on $\mathbb{B}$. Take an element $i\in\gamma$ and set $\mathbb{B}_{\gamma}=i\mathbb{B}i$. Denote the subgroup $\{a\in\mathbb{B}_{\gamma}^{*}\mid{{}^{a}}(Ri)=Ri\}$ of $\mathbb{B}_{\gamma}^{*}$ by $N_{\mathbb{B}_{\gamma}}(R)$. It has three normal subgroups $Ri$ and $(\mathbb{B}_{\gamma}^{R})^{*}$ and $1+J(\mathbb{B}_{\gamma}^{R})$. Denote the two quotient groups $N_{\mathbb{B}_{\gamma}}(R)/Ri\cdot(1+J(\mathbb{B}_{\gamma}^{R}))$ and $N_{\mathbb{B}_{\gamma}}(R)/Ri\cdot(\mathbb{B}_{\gamma}^{R})^{*}$ by $\bar{F}_{\mathbb{B}}(R)$ and $F_{\mathbb{B}}(R)$, respectively. It is well-known that $(\mathbb{B}_{\gamma}^{R})^{*}$ is isomorphic to $k^{*}\times(1+J(\mathbb{B}_{\gamma}^{R}))$. We can get the following short exact sequence
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${k^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bar{F}_{\mathbb{B}}(R)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{\mathbb{B}}(R)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
When $\mathbb{B}$ is the group algebra $\mbox{$\mathcal{O}$}L$ for some finite group $L$, by [20, Theorem 3.1], there is a group isomorphism between $N_{L}(R_{\gamma})/RC_{L}(R)$ and $F_{\mathcal{O}L}(R)$ which preserves the actions on $R$ up to inner automorphisms.
2.13.
For any element $x$ in $H$ and any subgroup $K$ of $H$, we denote by $[x,K]$ the subgroup of $H$ generated by elements $xux^{-1}u^{-1}$, where $u$ runs over $K$. Let $\mathfrak{b}$ be a block of $H$. Then $H_{\{\mathfrak{b}\}}$ is a pointed group on the $H$-interior algebra $\mathcal{O}H$. Let $Q_{\delta}$ be a defect pointed group of the block $\mathfrak{b}$ and $A$ a source algebra of $\mathcal{O}H\mathfrak{b}$. Then $Q$ can act on $A$ through the conjugation. Recall that the hyperfocal subgroup $\mathfrak{hyp}(Q_{\delta})$ of $Q_{\delta}$ is the subgroup of $Q$ generated by the subgroups $[x,T]$ where $T_{\epsilon}$ runs over the set of local pointed groups on $\mathcal{O}H$ such that $T_{\epsilon}\subseteq Q_{\delta}$, and $x$ over the set of $p^{\prime}$-elements of $N_{H}(T_{\epsilon})$ (see [25, 1.7.2]). Sometimes, we also call $\mathfrak{hyp}(Q_{\delta})$ a hyperfocal subgroup of $Q$ for simplicity. It is easy to see that $\mathfrak{hyp}(Q_{\delta})$ is a normal subgroup of $Q$. Denote by $\mathcal{S}$ a set of representatives for $Q/\mathfrak{hyp}(Q_{\delta})$ in $Q$. By [25, Theorem 1.8], there is a $Q$-stable $\mathcal{O}$-subalgebra $\mathbb{A}$ of $A$, unique up to $(A^{Q})^{*}$-conjugation, containing the image of $\mathfrak{hyp}(Q_{\delta})$ and fulfilling $A=\mathop{\bigoplus}\limits_{v\in\mathcal{S}}\mathbb{A}\cdot v$. It is clear that $\mathbb{A}$ is an $\mathfrak{hyp}(Q_{\delta})$-interior algebra inherited from the $Q$-interior algebra structure of $A$. The $\mathfrak{hyp}(Q_{\delta})$-interior algebra $\mathbb{A}$ is called a hyperfocal subalgebra of $A$. Sometimes, we also call it a hyperfocal subalgebra of the block $\mathfrak{b}$.
In the following paragraph, we will collect some notation and definition about endopermutation modules over a $p$-group which is introduced by Dade [4, 5].
2.14.
Let $P$ be a $p$-group. An $\mathcal{O}P$-module $V$ is called an endopermutation $\mathcal{O}P$-module if the endomorphism algebra $\mathrm{End}_{\mathcal{O}}(V)$ has a $P$-stable $\mathcal{O}$-basis under the $P$-conjugation action. Let $V$ be an indecomposable endopermutation $\mbox{$\mathcal{O}$}P$-module with vertex $P$, which means that $S(P)\neq 0$. Here, $S=\mathrm{End}_{\mathcal{O}}(V)$, which is called a Dade $P$-interior algebra. Then by [17, Proposition 7.3.7], for any subgroup $Q$ of $P$, there is up to isomorphism a unique endopermutation $k(N_{P}(Q)/Q)$-module $W_{Q}$ such that $S(Q)\cong\mathrm{End}_{k}(W_{Q})$ as $N_{P}(Q)/Q$-algebras. It is clear that $W_{Q}$ is an indecomposable endopermutation $k(N_{P}(Q)/Q)$-module. Moreover, let $R$ be a subgroup of $P$ such that $Q$ is normal in $R$. We denote by $\emph{Defres}_{R/Q}^{P}(V)$ the endopermutation $k(R/Q)$-module $\mathrm{Res}_{R/Q}^{N_{P}(Q)/Q}(W_{Q})$. In particular, when $Q$ is normal in $P$, $\mathrm{Defres}_{P/Q}^{P}(V)$ is just the indecomposable endopermutation $kP/Q$-module $W_{Q}$.
2.15.
Finally, let us recall the definitions of twisted group algebra and relative separability. Let $\alpha\in Z^{2}(H;k^{*})$ be a $2$-cocycle of $H$. Since $k$ is perfect, we have a canonical isomorphism $\mbox{$\mathcal{O}$}^{*}\cong k^{*}\times(1+J(\mbox{$\mathcal{O}$}))$ (see [28]). So we can view $k^{*}$ as a subgroup of $\mbox{$\mathcal{O}$}^{*}$. The twisted group algebra of $H$ by $\alpha$, denoted by $\mbox{$\mathcal{O}$}_{\alpha}H$, is the $\mathcal{O}$-algebra with an $\mathcal{O}$-basis consisting of all elements of $H$ and the multiplication induced by the equality $\lambda_{x}x\cdot\lambda_{y}y=(\lambda_{x}\lambda_{y}\alpha(x,y))xy$ for all $x,y\in H$ and all $\lambda_{x},\lambda_{y}\in\mbox{$\mathcal{O}$}$. An $H$-interior algebra $\mathbb{B}$ is called relatively $\mbox{$\mathcal{O}$}H$-separable if $\mathbb{B}$ is isomorphic to a direct summand of $\mathbb{B}\mathop{\otimes}\limits_{\mathcal{O}H}\mathbb{B}$ as $\mathbb{B}$-$\mathbb{B}$-bimodules (see [16, §4.8]). This is equivalent to requiring that the canonical $\mathbb{B}$-$\mathbb{B}$-bimodule homomorphism $\mathbb{B}\mathop{\otimes}\limits_{\mathcal{O}H}\mathbb{B}\longrightarrow\mathbb{B}$ induced by multiplication in $\mathbb{B}$ splits by [16, Theorem 2.6.10].
3 Hyperfocal abelian Frobenius blocks
In this section, we will give the definition of the hyperfocal abelian Frobenius blocks and investigate their local structures.
3.1.
Let $G$ be a finite group and $b$ a block of $G$ over $\mathcal{O}$ with defect pointed group $P_{\gamma}$. We denote the hyperfocal subgroup of $P_{\gamma}$ by $D$. Then by [25, Propositions 4.2 and 3.3], there is a unique local point $\delta$ of $D$ on $\mbox{$\mathcal{O}$}Gb$ such that $D_{\delta}\leq P_{\gamma}$. Take an element $i$ of $\gamma$ and set $A=i\mbox{$\mathcal{O}$}Gi$. Then $A$ is a source algebra of $\mbox{$\mathcal{O}$}Gb$ and let $\mathbb{A}$ be a hyperfocal subalgebra of $A$. We take an element $j$ of $\delta$ such that $ij=j=ji$. By [25, Proposition 2.8], we can assume that $j$ belongs to $\mathbb{A}$. We can view $D_{\delta}$ as a local pointed group of $D$ on both $A$ and $\mathbb{A}$. Set $A_{\delta}=jAj$ and $\mathbb{A}_{\delta}=j\mathbb{A}j$. Clearly, $A_{\delta}$ and $\mathbb{A}_{\delta}$ are embedded algebras of the local pointed group $D_{\delta}$ on $A$ and $\mathbb{A}$, respectively. By the uniqueness of $D_{\delta}$, there is an element $a_{u}$ in $(\mathbb{A}^{D})^{*}$ such that $a_{u}ja_{u}^{-1}=uju^{-1}$ for any $u\in P$. Denote $a_{u}^{-1}uj$ by $a_{u,\delta}$ for any $u\in P$. Then $a_{u,\delta}$ belongs to $A_{\delta}^{*}$ fulfilling $a_{u,\delta}(vj)a_{u,\delta}^{-1}=(uvu^{-1})j$ for any $v\in D$. Moreover, we have $A_{\delta}=\mathop{\bigoplus}\limits_{uD\in P/D}\mathbb{A}_{\delta}a_{u,\delta}$ as $\mathcal{O}$-modules.
3.2.
It is well-known that the source algebra $A$ is closely related to the block algebra $\mbox{$\mathcal{O}$}Gb$. For example, they are both relatively $\mbox{$\mathcal{O}$}P$-separable $P$-interior algebras; there is a canonical Morita equivalence between the source algebra $A$ and the block algebra $\mbox{$\mathcal{O}$}Gb$ induced by the $A$-$\mbox{$\mathcal{O}$}Gb$-bimodule $i\mbox{$\mathcal{O}$}G$. Next, we will show that these properties also hold for the hyperfocal subalgebra $\mathbb{A}$ and its embedded subalgebra $\mathbb{A}_{\delta}$ by the following general result.
Lemma 3.3.
Let $Q$ be a $p$-subgroup and $C$ an indecomposable relatively $\mbox{$\mathcal{O}$}Q$-separable $Q$-interior algebra. Assume that $C$ has a unique local pointed group $Q_{\epsilon}$ and as $\mbox{$\mathcal{O}$}(Q\times Q)$-modules, the indecomposable direct summand of $C$ has the form $\mathrm{Ind}_{\Delta_{\varphi}(T)}^{Q\times Q}(\mbox{$\mathcal{O}$})$ for some subgroup $T$ of $Q$ and some injective homomorphism $\varphi:T\longrightarrow Q$. Here, $\Delta_{\varphi}(T)=\{(u,\varphi(u))\mid u\in T\}$. Moreover, suppose that $\frac{\mathrm{rank}_{\mathcal{O}}(C)}{|Q|}$ is coprime to $p$. Take an element $l$ of $\epsilon$. Then the $Q$-interior algebra $lCl$ is a relatively $\mbox{$\mathcal{O}$}Q$-separable $Q$-interior algebra and the lCllCl-$C$-bimodule lClC induces a Morita equivalence between $lCl$ and $C$.
Proof.
Since the indecomposable direct summand of the $\mbox{$\mathcal{O}$}(Q\times Q)$-module $C$ has the form $\mathrm{Ind}_{\Delta_{\varphi}(T)}^{Q\times Q}(\mbox{$\mathcal{O}$})$, $C$ is a direct sum of some copies of $\mbox{$\mathcal{O}$}Q$ as left $\mbox{$\mathcal{O}$}Q$ and right $\mbox{$\mathcal{O}$}Q$-modules. In particular, $C$ is free as left and right $\mbox{$\mathcal{O}$}Q$-modules. Hence, the structural map $Q\rightarrow C^{*}$ is injective and we can view $\mbox{$\mathcal{O}$}Q$ as a subalgebra of $C$. Obviously, every finitely generated projective $C$-module has $\mathcal{O}$-rank divided by $|Q|$.
By the definition of relatively $\mbox{$\mathcal{O}$}Q$-separable $Q$-interior algebras, $C$ is a direct summand of $C\mathop{\otimes}\limits_{\mathcal{O}Q}C$ as $C$-$C$-bimodules. Since $C$ is an indecomposable algebra, there are some primitive idempotents $e,f$ of $C^{Q}$ such that $C$ is a direct summand of $Ce\mathop{\otimes}\limits_{\mathcal{O}Q}fC$. Suppose that $\mathrm{Br}_{Q}(e)=0$. By [29, Theorem (23.1)], there is a proper subgroup $Q_{0}$ of $Q$ and a primitive idempotent $e_{0}$ of $C^{Q_{0}}$ such that $e=\mathrm{Tr}_{Q_{0}}^{Q}(j_{0})$ and ${{}^{u}}e_{0}\cdot e=0$ for any $u\in Q-Q_{0}$. Then the map sending $c\in Ce$ to $\sum\limits_{u\in Q/Q_{0}}cue_{0}\mathop{\otimes}\limits_{\mathcal{O}Q_{0}}u^{-1}$ induces a $C$-$\mbox{$\mathcal{O}$}Q$-bimodule isomorphism from $C_{e}$ to $Ce_{0}\mathop{\otimes}\limits_{\mathcal{O}Q_{0}}\mbox{$\mathcal{O}$}Q$. Therefore, $C$ is a direct summand of $Ce_{0}\mathop{\otimes}\limits_{\mathcal{O}Q_{0}}fC$ as $C$-$C$-bimodules. Denote by $\bar{C}$ the $k$-algebra $C/J(\mbox{$\mathcal{O}$})C$. Then for any simple $C$-module $W$, we have $W$ is a direct summand of $\bar{C}\mathop{\otimes}\limits_{kQ_{0}}U$ for some indecomposable $kQ_{0}$-module $U$. By assumption and the Mackey decomposition formula, the indecomposable direct summand of $\mathrm{Res}_{Q\times Q_{0}}^{Q\times Q}(\bar{C})$ has the form $kQ\mathop{\otimes}\limits_{kT}({i}kQ_{0})$ for some proper subgroup $T$ of $Q$ and injective homomorphism $\psi:T\longrightarrow Q_{0}$. By Green’s indecomposability theorem, this implies that the dimension of every simple $C$-module has dimension divided by $p$. Let $P_{W}$ be the projective cover of the simple $C$-module $W$. We have $\mathrm{dim}(\bar{C})=\sum\limits_{W}\mathrm{dim}(P_{W})\mathrm{dim}(W)$ which is divided by $p|Q|$. This is contradict the assumption that $\frac{\mathrm{rank}_{\mathcal{O}}(C)}{|Q|}$ is coprime to $p$.
We conclude that $\mathrm{Br}_{Q}(e)\neq 0$. Similarly, we can also get $\mathrm{Br}_{Q}(f)\neq 0$. Since $\epsilon$ is the unique local point of $Q$ on $C$, we have $e$ and $f$ both belongs to $\epsilon$. In particular, we have $Ce$ is isomorphic to $Cl$ as $C$-$\mbox{$\mathcal{O}$}Q$-bimodules and $fC$ is isomorphic to $lC$ as $\mbox{$\mathcal{O}$}Q$-$C$-bimodules. So $C$ is a direct summand of $Cl\mathop{\otimes}\limits_{\mathcal{O}Q}lC$ as $C$-$C$-bimodules and then $lCl$ is a direct summand of $lCl\mathop{\otimes}\limits_{\mathcal{O}Q}lCl$ as lCllCl-lCllCl-bimodules. This means that $lCl$ is a relatively $\mbox{$\mathcal{O}$}Q$-separable $Q$-interior algebra.
Since $C$ is a direct summand of $Cl\mathop{\otimes}\limits_{\mathcal{O}Q}lC$, $W$ is a direct summand of $Cl\mathop{\otimes}\limits_{\mathcal{O}Q}l\cdot W$ for any simple $C$-module $W$. In particular, $l\cdot W$ is not equal to 0 for any simple $C$-module $W$. By [16, Theorem 2.8.7], the lCllCl-$C$-bimodule lClC induces a Morita equivalence between $lCl$ and $C$. We are done. ∎
Corollary 3.4.
The embedded subalgebra $\mathbb{A}_{\delta}$ of $\mathbb{A}$ is a relatively $\mbox{$\mathcal{O}$}D$-separable $D$-interior algebra and the $\mathbb{A}_{\delta}$-$\mathbb{A}$-bimodule $j\mathbb{A}$ induces a Morita equivalence between $\mathbb{A}_{\delta}$ and $\mathbb{A}$.
Proof.
By [32, Theorem 1], the hyperfocal subalgebra $\mathbb{A}$ is a relatively $\mbox{$\mathcal{O}$}D$-separable $D$-interior algebra. By [9, Lemma 3.2], $\mathbb{A}$ is an indecomposable $\mathcal{O}$-algebra. From the paragraph 3.1, we can get $D$ has a unique local point on $\mathbb{A}$. Since $A=\mathop{\bigoplus}\limits_{uD\in P/D}\mathbb{A}u$, we have $\frac{\mathrm{rank}_{\mathcal{O}}(\mathbb{A})}{|D|}=\frac{\mathrm{rand}_{\mathcal{O}}(A)}{|P|}$ which is coprime to $p$ by [29, Corollary (44.8)]. Since $\mathbb{A}$ is a direct summand of $A$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules, the statements in the corollary follow from [17, Theorem 8.7.1] and Lemma 3.3 above. ∎
3.5.
Let us consider the quotient group $N_{G}(D_{\delta})/DC_{G}(D)$. Different from the inertial quotient $N_{G}(P_{\gamma})/PC_{G}(P)$ being a $p^{\prime}$-group, $N_{G}(D_{\delta})/DC_{G}(D)$ may have a nontrivial Sylow $p$-subgroup $PC_{G}(D)/DC_{G}(D)$ in general. However, we can show that it is a $p$-nilpotent group, which means that it has a normal $p$-complement. We call this normal $p$-complement a hyperfocal inertial quotient of the block $b$ and denote it by $E_{\mathfrak{h}}$ throughout this paper. Hence, $E_{\mathfrak{h}}$ is a $p^{\prime}$-group and $N_{G}(D_{\delta})/DC_{G}(D)=PC_{G}(D)/DC_{G}(D)\cdot E_{\mathfrak{h}}$. To show this result, we need the following lemma about the structures of the Brauer quotients $A_{\delta}(D)$ and $\mathbb{A}_{\delta}(D)$ of $A_{\delta}$ and $\mathbb{A}_{\delta}$ at $D$.
Lemma 3.6.
We have $A_{\delta}(D)\cong kC_{P}(D)$ as $Z(D)$-interior algebras such that its restriction to $\mathbb{A}_{\delta}(D)$ gives a $Z(D)$-interior algebra isomorphism $\mathbb{A}_{\delta}(D)\cong kZ(D)$. In particular, $(j\mathbb{A}uj)(D)\cong kZ(D)u$ when $u$ is in $C_{P}(D)$ and $(j\mathbb{A}uj)(D)=0$ when $u$ is not in $DC_{P}(D)$.
Proof.
Let $\bar{b}_{D}$ be the block of $C_{G}(D)$ over $k$ associated with the local pointed group $D_{\delta}$. It is clear that $\bar{b}_{D}$ is $P$-stable and nilpotent by [25, Proposition 3.3]. Then $\bar{b}_{D}$ can be viewed as a block of $PC_{G}(D)$ and we denote it by $\bar{c}_{D}$ when it is viewed as a block of $PC_{G}(D)$. So the block $\bar{c}_{D}$ is also nilpotent by [15, Proposition 6.5]. It is easy to check that the block $\bar{c}_{D}$ has a defect group $P$ and then $C_{P}(D)$ is a defect group of the block $\bar{b}_{D}$.
By [29, Lemma (40.2)], we have $\mathrm{Br}_{D}(i)$ belongs to a local point of $P$ on $kC_{G}(D)$. In particular, $\mathrm{Br}_{D}(i)kPC_{G}(D)\mathrm{Br}_{D}(i)$ is a source algebra of the block $\bar{c}_{D}$. Note that $A(D)=\mathrm{Br}_{D}(i)kC_{G}(D)\mathrm{Br}_{D}(i)$ is a primitive $P$-algebra. Then by [17, Theorem 8.12.3], there is a primitive Dade $P$-interior algebra $T_{D}$ over $k$ such that we have an isomorphism of $P$-interior algebras
$\mathrm{Br}_{D}(i)kPC_{G}(D)\mathrm{Br}_{D}(i)\cong T_{D}\mathop{\otimes}\limits_{k}kP$
which restricts to a $C_{P}(D)$-interior algebra isomorphism
$A(D)\cong T_{D}\mathop{\otimes}\limits_{k}kC_{P}(D).$ (3.6.1)
Since $D$ is normal in $PC_{G}(D)$, we have $v\cdot 1_{T_{D}}=1_{T_{D}}$ for any $v\in D$. By the isomorphism (3.6.1), we have a $Z(D)$-interior algebra isomorphism $A_{\delta}(D)\cong kC_{P}(D)$. Obviously, the image of $\mathbb{A}_{\delta}(D)$ under this isomorphism contains $kZ(D)$.
Borrowing the notation from the paragraph 3.1, for any $u\in P$, there is an element $a_{u,\delta}$ of $A_{\delta}^{*}$ such that ${{}^{a_{u,\delta}}}(v\cdot j)=({{}^{u}}v)\cdot j$ for any $v\in D$ and $A_{\delta}=\mathop{\bigoplus}\limits_{uD\in P/D}\mathbb{A}_{\delta}a_{u,\delta}$. It is clear that $\mathbb{A}_{\delta}a_{u,\delta}$ is a $D$-stable $\mathcal{O}$-module for any $u\in P$ and then $A_{\delta}(D)=\mathop{\bigoplus}\limits_{uD\in P/D}(\mathbb{A}_{\delta}a_{u,\delta})(D)$. When $u$ is in $C_{P}(D)$, $a_{u,\delta}$ is in $A_{\delta}^{D}$ and then $(\mathbb{A}_{\delta}a_{u,\delta})(D)=\mathbb{A}_{\delta}(D)\mathrm{Br}_{D}(a_{u,\delta})$ and in particular they have the same dimension. Therefore, $\mathop{\bigoplus}\limits_{uZ(D)\in C_{P}(D)/Z(D)}(\mathbb{A}_{\delta}a_{u,\delta})(D)$ has dimension greater than or equal to $|C_{P}(D)|$. Then we are done. ∎
Now we can show that the quotient group $N_{G}(D_{\delta})/DC_{G}(D)$ is a $p$-nilpotent group.
Proposition 3.7.
The quotient group $N_{G}(D_{\delta})/DC_{G}(D)$ is a $p$-nilpotent group.
Proof.
Let us borrow the notation from the paragraphs 2.12 and 3.1. Set $P(A_{\delta}^{D})^{*}=(A_{\delta}^{D})^{*}\cdot\{a_{u,\delta}\mid u\in P\}$. Then $P(A_{\delta}^{D})^{*}$ is a subgroup of $N_{A_{\delta}}(D)$.
By [25, Proposition 2.8], we have
$F_{A}(D)=(P(A_{\delta}^{D})^{*})\cdot N_{\mathbb{A}_{\delta}}(D)/D(A_{\delta}^{D})^{*}$
and $N_{\mathbb{A}_{\delta}}(D)(A_{\delta}^{D})^{*}/D(A_{\delta}^{D})^{*}$ is normal in $F_{A}(D)$. Take an element $\mathfrak{a}_{\delta}$ of $P(A_{\delta}^{D})^{*}\cap N_{\mathbb{A}_{\delta}}(D)$. So there is some $u\in P$ such that ${{}^{\mathfrak{a}_{\delta}}}vj={{}^{u}}vj$ for any $v\in D$. Then $\mathfrak{a}_{\delta}^{-1}u$ is in $(j\mathbb{A}ju)^{D}$. By the proof of Lemma 3.6, there is some $\mathfrak{a}\in(\mathbb{A}^{D})^{*}$ such that $\mathfrak{a}^{-1}j\mathfrak{a}=u^{-1}ju$. Therefore, $j\mathbb{A}ju=j\mathbb{A}u\mathfrak{a}^{-1}j\mathfrak{a}=j\mathbb{A}uj\mathfrak{a}$. Combining these facts together, we have $(j\mathbb{A}uj)(D)\neq 0$. By Lemma 3.6, this forces $u\in DC_{P}(D)$. Then the element $\mathfrak{a}_{\delta}$ has to belong to $D(\mathbb{A}_{\delta}^{D})^{*}$. Hence, $(P(A_{\delta}^{D})^{*}/D(A_{\delta}^{D})^{*})\cap(N_{\mathbb{A}_{\delta}}(D)(A_{\delta}^{D})^{*}/D(A_{\delta}^{D})^{*})=1$.
On the other hand, $F_{A}(D)$ is isomorphism to $N_{G}(D_{\delta})/DC_{G}(D)$ with $PC_{G}(D)/DC_{G}(D)$ being a Sylow $p$-subgroup. It is clear $P(A_{\delta}^{D})^{*}/D(A_{\delta}^{D})^{*}$ is the preimage of $PC_{G}(D)/DC_{G}(D)$ under this isomorphism. In conclusion, $F_{A}(D)$ is a $p$-nilpotent group. So is $N_{G}(D_{\delta})/DC_{G}(D)$. ∎
3.8.
Now we can define the hyperfocal abelian Frobenius block. Keep the notation above. By the Schur-Zassenhaus Theorem, we can identify $E_{\mathfrak{h}}$ with a subgroup of $\mathrm{Aut}(D)$ and then it can act on $D$ by conjugation. Then the block $b$ is called with Frobenius hyperfocal inertial quotient if $E_{\mathfrak{h}}$ is nontrivial and acts freely on $D-\{1\}$, equivalently, the corresponding semidirect product $D\rtimes E_{\mathfrak{h}}$ is a Frobenius group. Moreover, a block $b$ with Frobenius hyperfocal inertial quotient is called a hyperfocal abelian Frobenius block if a hyperfocal subgroup of the block $b$ is abelian.
Remark 3.9.
There are some examples of hyperfocal abelian Frobenius blocks. The first well-known example is the block with cyclic hyperfocal subgroup. The second example is the block with Klein four hyperfocal subgroup. In this case, by [7, Proposition 2.3], $N_{G}(D_{\delta})/C_{G}(D)$ is isomorphic to a cyclic group of order $3$, or the symmetric group $\mathrm{S}_{3}$ of degree 3, depending on $C_{P}(D)=P$ or not. In both cases, they are both hyperfocal abelian Frobenius blocks.
In the following, we investigate the local structures of hyperfocal abelian Frobenius blocks.
3.10.
For any element $x$ of $\mbox{$\mathcal{O}$}G$, we use $\bar{x}$ to denote the image of $x$ in $kG$ under the natural surjective homomorphism. There is an analoguous notation for the subset of $\mbox{$\mathcal{O}$}G$. Denote by $b_{P}$ the block of $C_{G}(P)$ over $\mathcal{O}$ such that the block $\bar{b}_{P}$ of $kC_{G}(P)$ is associated with the local pointed group $P_{\gamma}$. Hence, the pair $(P,\bar{b}_{P})$ is a maximal $b$-Brauer pair. It is well-known that for any subgroup $R$ of $P$ there is a unique block $b_{R}$ of $C_{G}(R)$ over $\mathcal{O}$ with $(R,\bar{b}_{R})\leq(P,\bar{b}_{P})$ and the pair $(R,\bar{b}_{R})$ is called a $b$-Brauer pair. Denote by $\mathcal{B}_{G}(b)_{\leq(P,\bar{b}_{P})}$ the Brauer category of the block $b$ with objects consisting of all $b$-Brauer pairs contained in $(P,\bar{b}_{P})$. Let $H$ be a subgroup of $G$. $H$ is said to control fusion in $\mathcal{B}_{G}(b)_{\leq(P,\bar{b}_{P})}$, or control fusion of the block $b$ for simplicity, if the morphisms of the Brauer category $\mathcal{B}_{G}(b)_{\leq(P,\bar{b}_{P})}$ can be induced by elements of $H$ (see [29, §49]).
3.11.
Let us recall the definitions of some special local pointed groups. A self-centralizing pointed group $Q_{\epsilon}$ on $\mbox{$\mathcal{O}$}G$ is a local pointed group on $\mbox{$\mathcal{O}$}G$ such that $Z(Q)$ is a defect group of the block $b(\epsilon)$ of $C_{G}(Q)$ associated with $Q_{\epsilon}$. Furthermore, a self-centralizing pointed group $Q_{\epsilon}$ is called an essential pointed group if the quotient $N_{G}(Q_{\epsilon})/QC_{G}(Q)$ contains a proper subgroup $M$ such that $p$ divides $|M|$ but does not divide $|M\cap M^{x}|$ for any $x\in(N_{G}(Q_{\epsilon})/QC_{G}(Q))-M$. In particular, $N_{G}(Q_{\epsilon})/QC_{G}(Q)$ has no nontrivial normal $p$-subgroup. For an essential pointed group $Q_{\epsilon}$, it is clear that $N_{G}(Q_{\epsilon})=N_{G}(Q,\bar{b}(\epsilon))$ and we also call the corresponding $b$-Brauer pair $N_{G}(Q,\bar{b}(\epsilon))$ an essential $b$-Brauer pair.
Proposition 3.12.
Suppose that the block $b$ is a hyperfocal abelian Frobenius block. Then for any nontrivial subgroup $Q$ of $D$, the block $b_{Q}$ of $C_{G}(Q)$ over $\mathcal{O}$ is nilpotent.
Proof.
Let $Q$ be a nontrivial subgroup of $D$. Suppose that $C_{P}(Q)$ is a defect group of the block $b_{Q}$ of $C_{G}(Q)$. We denote $C_{P}(Q)$ by $Q_{c}$. We first prove this proposition for this special case. For any subgroup $R$ of $C_{P}(Q)$, $(R,\bar{b}_{QR})$ is a $b_{Q}$-Brauer pair. In particular $(Q_{c},\bar{b}_{Q_{c}})$ is a maximal $b_{Q}$-Brauer pair. In order to show that the block $b_{Q}$ is nilpotent, it suffices to prove that the hyperfocal subgroup of the block $b_{Q}$ is trivial. By [25, 1.7.1], it reduces to calculate that $N_{C_{G}(Q)}(R,\bar{b}_{QR})/C_{G}(QR)$ is a $p$-group when $(R,\bar{b}_{QR})$ is an essential $b_{Q}$-Brauer pair and $(R,\bar{b}_{QR})=(Q_{c},\bar{b}_{Q_{c}})$.
It is clear that $N_{C_{G}(Q)}(R,\bar{b}_{QR})=N_{G}(QR,\bar{b}_{QR})\cap C_{G}(Q)$. By [33, Theorem 2], $N_{G}(QR,\bar{b}_{QR})=(N_{G}(D,\bar{b}_{D})\cap N_{G}(QR,\bar{b}_{QR}))C_{G}(QR)$. Therefore,
$N_{C_{G}(Q)}(R,\bar{b}_{QR})=C_{G}(QR)(N_{G}(D,\bar{b}_{D})\cap N_{G}(QR,\bar{b}_{QR})\cap C_{G}(Q))$
and then
$N_{C_{G}(Q)}(R,\bar{b}_{QR})/C_{G}(QR)\cong(N_{G}(D,\bar{b}_{D})\cap N_{G}(QR,\bar{b}_{QR})\cap C_{G}(Q))/(N_{G}(D,\bar{b}_{D})\cap C_{G}(QR)).$
Obviously, $N_{G}(D,\bar{b}_{D})/C_{G}(D)$ is isomorphic to $N_{G}(D_{\delta})/C_{G}(D)$. Since $E_{\mathfrak{h}}$ acts freely on $D-\{1\}$, we have $N_{G}(D,\bar{b}_{D})\cap C_{G}(Q)=Q_{c}C_{G}(D)$.
Suppose that $D\leq R$. Then
$C_{G}(QR)=C_{G}(R)~\text{and}~N_{C_{G}(Q)}(R,\bar{b}_{R})=N_{Q_{c}C_{G}(D)}(R,\bar{b}_{R})~\text{and}~N_{G}(D,\bar{b}_{D})\cap C_{G}(R)=C_{G}(R).$
In this case, the pair $(R,\bar{b}_{R})$ is also a $b_{D}$-Brauer pair when $b_{D}$ is viewed as a block of $Q_{c}C_{G}(D)$ over $\mathcal{O}$. Since by [25, Proposition 3.3] the block $b_{D}$ of $Q_{c}C_{G}(D)$ is nilpotent with defect group $Q_{c}$, we have
$N_{C_{G}(Q)}(R,\bar{b}_{R})=N_{Q_{c}C_{G}(D)}(R,\bar{b}_{R})=N_{Q_{c}}(R)C_{G}(R).$
Hence,
$N_{C_{G}(Q)}(R,\bar{b}_{R})/C_{G}(R)\cong N_{Q_{c}}(R)C_{G}(R)/C_{G}(R)$
is a $p$-group.
Now let $(R,\bar{b}_{QR})$ be an essential $b_{Q}$-Brauer pair. Clearly, $(R,\bar{b}_{R})\leq(QR,\bar{b}_{QR})$ and $R\unlhd QR$. We have $\mathrm{Br}_{QR}(\bar{b}_{R})\bar{b}_{QR}=\bar{b}_{QR}$. Since $Z(R)$ is the defect group of the block $b_{R}$ of $C_{G}(R)$, $R$ is the defect group of the block $b_{R}$ viewed as a block of $RC_{G}(R)$. Note that $QR\leq RC_{G}(R)$. We have $QR\leq R$. In particular, $Q\leq R$ and then $C_{G}(QR)=C_{G}(R)$. On the other hand, by the argument above, we have $N_{D}(R)C_{G}(R)/C_{G}(R)$ is a normal $p$-subgroup of $N_{G}(R,\bar{b}_{R})/C_{G}(R)$. Therefore, $N_{D}(R)$ is contained in $RC_{G}(R)$. Set $T=RN_{D}(R)$ and so $R\unlhd T$. We have $\mathrm{Br}_{T}(\bar{b}_{R})\bar{b}_{T}=\bar{b}_{T}$. Similarly, we can get $T\leq R$ which forces $D\leq R$.
In conclusion, when $(R,\bar{b}_{QR})$ is either essential or equal to $(Q_{c},\bar{b}_{Q_{c}})$, we both have $D\leq R$. We are done by the fourth paragraph under the assumption that $C_{P}(Q)$ is a defect group of the block $b_{Q}$.
In general, it is well-known that there is some element $x$ of $G$ such that $(Q,\bar{b}_{Q})\leq(P,\bar{b}_{P})^{x}$ and $C_{P^{x}}(Q)$ is a defect group of the block $b_{Q}$. Since the hyperfocal subgroup $D$ is abelian, we can assume that $x$ is in $N_{G}(D,\bar{b}_{D})$ by [33, Theorem 2]. Therefore, $C_{P}({{}^{x}}Q)$ is a defect group of the block ${{}^{x}}b_{Q}$ of $C_{G}({{}^{x}}Q)$. Since $x$ belongs to $N_{G}(D,\bar{b}_{D})$, ${{}^{x}}Q$ is still a nontrivial subgroup of $D$ and $({{}^{x}}Q,{{}^{x}}\bar{b}_{Q})$ is contained in $(D,\bar{b}_{D})$. By the uniquness of $\bar{b}_{{{}^{x}}Q}$, we get ${{}^{x}}b_{Q}=b_{{{}^{x}}Q}$. On the other hand, by the argument for the special case, the block $b_{{{}^{x}}Q}$ of $C_{G}({{}^{x}}Q)={{}^{x}}C_{G}(Q)$ is nilpotent. Obviously, the conjugation action of $x$ induces an isomorphism between the Brauer categories of the block $b_{Q}$ and $b_{{{}^{x}}Q}$. Hence, the block $b_{Q}$ of $C_{G}(Q)$ is also nilpotent. ∎
4 The bimodule structure and the Brauer quotients of $\mathbb{A}_{\delta}$
Keep the notation in the last section. We will first describe the $\mbox{$\mathcal{O}$}(D\times D)$-module structure of $\mathbb{A}_{\delta}$ without assuming that the block $b$ is a hyperfocal abelian Frobenius block. This is analogous to the $\mbox{$\mathcal{O}$}(P\times P)$-module structure of the source algebra $A$ (see [29, Theorem (44.3)]). Let us collect some basic results about the $\mbox{$\mathcal{O}$}(D\times D)$-module structure of $\mathbb{A}_{\delta}$.
Lemma 4.1.
The direct summand of $\mathbb{A}_{\delta}$ as an $\mbox{$\mathcal{O}$}(D\times D)$-module has the form $\mbox{$\mathcal{O}$}DgD$ for some $g\in G$, which is isomorphic to $\mathrm{Ind}_{\Delta_{g}(D)}^{D\times D}(\mbox{$\mathcal{O}$})$. Here, $\Delta_{g}(D)=\{(u,g^{-1}ug)\in D\times D\mid u\in D\cap{{}^{g}}D\}$ is a subgroup of $D\times D$. Moreover, setting $Q=D\cap{{}^{g}}D$, then $(Q,\bar{b}_{Q})^{g}\subseteq(D,\bar{b}_{D})$.
Proof.
Obviously, as $\mbox{$\mathcal{O}$}(D\times D)$-modules, $\mathbb{A}$ is a direct summand of $A$ and $\mathbb{A}_{\delta}$ is a direct summand of $\mathbb{A}$. Therefore, $\mathbb{A}_{\delta}$ is a direct summand of $A$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. The statements in this lemma can follow from [17, Theorem 8.7.1]. ∎
Lemma 4.2.
There exists a $(D\times D)$-invariant $\mathcal{O}$-basis $\mathcal{B}$ of $\mathbb{A}_{\delta}$ containing the unit element $j$ of $\mathbb{A}_{\delta}$.
Proof.
The existence of a $(D\times D)$-invariant $\mathcal{O}$-basis $\mathcal{B}$ follows from Lemma 4.1. In order to show that the $\mathcal{O}$-basis $\mathcal{B}$ can be chosen to contain $j$, it suffices to demonstrate that this $\mathcal{B}$ contains an invertible element of $\mathbb{A}_{\delta}^{D}$. So we can assume that $\mbox{$\mathcal{O}$}=k$. Set $j=\sum\limits_{x\in\mathcal{B}}\lambda_{x}x$ with $\lambda_{x}\in k$. Then $\mathrm{Br}_{D}(j)=\sum_{x\in\mathcal{B}^{D}}\lambda_{x}\mathrm{Br}_{D}(x)$ is the unit element of $\mathbb{A}_{\delta}(D)$ which is isomorphic to the local algebra $kZ(D)$ by Lemma 3.6. In particular, there is at least one element $x_{0}\in\mathcal{B}^{D}$ with $\lambda_{x_{0}}\neq 0$ and $\mathrm{Br}_{D}(x_{0})$ being invertible in $\mathbb{A}_{\delta}(D)$. Therefore, $x_{0}$ is invertible in $\mathbb{A}_{\delta}^{D}$ by [29, Theorem (3.2)]. ∎
Lemma 4.3.
The multiplicity of $\mbox{$\mathcal{O}$}D$ in any decomposition of $\mathbb{A}_{\delta}$ as a direct sum of indecomposable $\mbox{$\mathcal{O}$}(D\times D)$-modules is $1$.
Proof.
Denote by $\Delta(D)$ the diagonal group of $D\times D$. Then by the Mackey decomposition formula, the direct summand of $\mathrm{Res}^{D\times D}_{\Delta(D)}(\mbox{$\mathcal{O}$}DgD)$ has the form $\mathrm{Ind}_{\Delta(D)\cap{{}^{(u,v)}}\Delta_{g}(D)}^{\Delta(D)}(\mbox{$\mathcal{O}$})$. It is easy to check that $\Delta(D)\cap{{}^{(u,v)}}\Delta_{g}(D)=\Delta(D)$ if and only if $g\in DC_{G}(D)$ and when $g\in DC_{G}(D)$, $\mbox{$\mathcal{O}$}DgD=\mbox{$\mathcal{O}$}Dg$ is isomorphic to $\mbox{$\mathcal{O}$}D$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. Hence, $(\mbox{$\mathcal{O}$}DgD)(D)$ is isomorphic to $kZ(D)$ when $g\in DC_{G}(D)$. By Lemma 3.6, we have $\mathbb{A}_{\delta}(D)\cong kZ(D)$. Therefore, there is only one direct summand of $\mathbb{A}_{\delta}$ isomorphic to $\mbox{$\mathcal{O}$}D$. ∎
$4.$
Recall that $N_{G}(D_{\delta})/DC_{G}(D)$ is a $p$-nilpotent group with normal $p$-complement $E_{\mathfrak{h}}$. By the proof of Proposition 3.7, the isomorphism from $N_{G}(D_{\delta})/DC_{G}(D)$ to $F_{A}(D)$ maps $E_{\mathfrak{h}}$ onto $N_{\mathbb{A}_{\delta}}(D)(A_{\delta}^{D})^{*}/D(A_{\delta}^{D})^{*}$, which is isomorphic to $N_{\mathbb{A}_{\delta}}(D)/D(\mathbb{A}_{\delta}^{D})^{*}$. For any $g\in N_{G}(D_{\delta})$ with $gDC_{G}(D)\in E_{\mathfrak{h}}$, let $\mathfrak{a}_{g}\in\mathbb{A}_{\delta}^{*}$ such that $\mathfrak{a}_{g}(uj)\mathfrak{a}_{g}^{-1}=(gug^{-1})j$ for any $u\in D$.
Lemma 4.5.
As $\mbox{$\mathcal{O}$}(D\times D)$-modules, $\mbox{$\mathcal{O}$}D\mathfrak{a}_{g}$ is isomorphic to $\mbox{$\mathcal{O}$}Dg$ and it is a direct summand of $\mathbb{A}_{\delta}$ with multiplicity $1$.
Proof.
It is easy to check that mapping $u\mathfrak{a}_{g}$ to $ug$ gives an $\mbox{$\mathcal{O}$}(D\times D)$-module isomorphism from $\mbox{$\mathcal{O}$}D\mathfrak{a}_{g}$ to $\mbox{$\mathcal{O}$}Dg$. By Lemma 4.2, we can write $\mathbb{A}_{\delta}=\mbox{$\mathcal{O}$}Dj\oplus Y$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules for some $\mbox{$\mathcal{O}$}(D\times D)$-submodule $Y$. Then we have $\mathbb{A}_{\delta}=\mathbb{A}_{\delta}\mathfrak{a}_{g}=\mbox{$\mathcal{O}$}D\mathfrak{a}_{g}\oplus Y\mathfrak{a}_{g}$. Hence, $\mbox{$\mathcal{O}$}D\mathfrak{a}_{g}$ is a direct summand of $\mathbb{A}_{\delta}$. By Lemma 4.3, we can get the multiplicity of $\mbox{$\mathcal{O}$}D\mathfrak{a}_{g}$ is $1$. ∎
Proposition 4.6.
There is a decomposition of $\mathbb{A}_{\delta}$ as an $\mbox{$\mathcal{O}$}(D\times D)$-module
$\mathbb{A}_{\delta}=(\mathop{\bigoplus}\limits_{gDC_{G}(D)\in E_{\mathfrak{h}}}\mbox{$\mathcal{O}$}D\mathfrak{a}_{g})\oplus X,$ (4.6.1)
where $X$ is isomorphic to a direct sum of modules of the form $\mbox{$\mathcal{O}$}DhD$ for some $h\in G-N_{G}(D)$. The module $\mbox{$\mathcal{O}$}D\mathfrak{a}_{g}$ is independent of the choice of $g$ up to isomorphism and has multiplicity $1$ for any $gDC_{G}(D)\in E_{\mathfrak{h}}$. Moreover, there is a unitary subalgebra $\mathbb{B}$ of $\mathbb{A}_{\delta}$ isomorphic to the twisted group algebra $\mbox{$\mathcal{O}$}_{\alpha}(D\rtimes E_{\mathfrak{h}})$ for some $\alpha$ of $H^{2}(E_{\mathfrak{h}};k^{*})$ such that it contains $\mbox{$\mathcal{O}$}D$ and $\mathbb{A}_{\delta}=\mathbb{B}\oplus Y$ as $\mathbb{B}$-$\mathbb{B}$-bimodules. In particular, $\mathbb{B}$ is isomorphic to $\mathop{\bigoplus}\limits_{gDC_{G}(D)\in E_{\mathfrak{h}}}\mbox{$\mathcal{O}$}D\mathfrak{a}_{g}$ and $Y$ is isomorphic to $X$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules.
Proof.
By Lemma 4.5 and [29, Lemma (44.7)], we have $\mathbb{A}_{\delta}=(\mathop{\bigoplus}\limits_{gDC_{G}(D)\in E_{\mathfrak{h}}}\mbox{$\mathcal{O}$}D\mathfrak{a}_{g})\oplus X$ such that $\mathop{\bigoplus}\limits_{gDC_{G}(D)\in E_{\mathfrak{h}}}\mbox{$\mathcal{O}$}D\mathfrak{a}_{g}$ and $X$ have no common direct summand. Suppose that $X$ has a direct summand isomorphic $\mbox{$\mathcal{O}$}DhD$ with $h\in N_{G}(D)$. This means that $\mathbb{A}_{\delta}$ has a direct summand $\mbox{$\mathcal{O}$}D\mathfrak{a}_{h}$ such that $u\mathfrak{a}_{h}w^{-1}=\mathfrak{a}_{h}$ if and only if $w=h^{-1}uh$ for any $(u,w)\in D\times D$.
On the other hand, by Lemma 4.1, the element $h$ has to belong to $N_{G}(D_{\delta})$. Then $h=vx$ for some $v\in P$ and $xDC_{G}(D)\in E_{\mathfrak{h}}$. Since there is an element $\mathfrak{a}_{x}$ of $\mathbb{A}_{\delta}^{*}$ such that $\mathfrak{a}_{x}(uj)\mathfrak{a}_{x}^{-1}=(xux^{-1}j)$ for any $u\in D$, we can assume that $h=v$ is in $P$. Recall from the proof of Lemma 3.6, for $v\in P$, there is an element $a_{v,\delta}$ of $A_{\delta}^{*}\cap j\mathbb{A}vj$ such that $a_{v,\delta}(uj)a_{v,\delta}^{-1}=(vuv^{-1})j$ for any $u\in D$. Therefore, $\mathfrak{a}_{v}a_{v,\delta}^{-1}$ is in $j\mathbb{A}v^{-1}j$ and $u\mathfrak{a}_{v}a_{v,\delta}^{-1}w^{-1}=\mathfrak{a}_{v}a_{v,\delta}^{-1}$ if and only if $u=w$ for any $(u,w)\in D\times D$. In particular, $\mbox{$\mathcal{O}$}D\mathfrak{a}_{v}a_{v,\delta}^{-1}$ is isomorphic to $\mbox{$\mathcal{O}$}D$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules, which implies that $(j\mathbb{A}v^{-1}j)(D)\neq 0$. By Lemma 3.6, we have $v$ is in $DC_{P}(D)$. Then $\mbox{$\mathcal{O}$}Dv$ is isomorphic to $\mbox{$\mathcal{O}$}D$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules, from which we can deduce that $X$ has a direct summand isomorphic to $\mbox{$\mathcal{O}$}D\mathfrak{a}_{x}$. This contradicts Lemma 4.5. So the direct summand of $X$ has the form $\mbox{$\mathcal{O}$}DhD$ with $h\in G-N_{G}(D)$. The second statement follows from Lemma 4.5.
Let us consider the remaining statements of this proposition. For any $g\in N_{G}(D_{\delta})$, we denote $gDC_{G}(D)\in N_{G}(D_{\delta})/DC_{G}(D)$ by $\tilde{g}$. By the proof of Proposition 3.7, we have $N_{\mathbb{A}_{\delta}}(D)/D(\mathbb{A}_{\delta}^{D})^{*}$ isomorphic to $E_{\mathfrak{h}}$. In particular, we have the following short exact sequence
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${D(\mathbb{A}_{\delta}^{D})^{*}/(D\times k^{*})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N_{\mathbb{A}_{\delta}}(D)/(D\times k^{*})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{\mathfrak{h}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
It is clear that $D(\mathbb{A}_{\delta}^{D})^{*}/(D\times k^{*})$ is isomorphic to $j+J(\mathbb{A}_{\delta})^{D}$. Since $E_{\mathfrak{h}}$ is a $p^{\prime}$-group, by [29, Lemma (45.6)], this sequence splits. So there is a subgroup $\hat{E}_{\mathfrak{h}}$ of $N_{\mathbb{A}_{\delta}}(D)$ containing $D\times k^{*}$ such that $\hat{E}_{\mathfrak{h}}\cap(D(\mathbb{A}_{\delta}^{D})^{*})=D\times k^{*}$ and $\hat{E}_{\mathfrak{h}}/(D\times k^{*})\cong E_{\mathfrak{h}}$. This yields another short exact sequence
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(D\times k^{*})/k^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\hat{E}_{\mathfrak{h}}/k^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{\mathfrak{h}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1,}$
which still splits by the Schur-Zassenhaus Theorem. Now let $\mathbb{M}$ be a subgroup of $\hat{E}_{\mathfrak{h}}$ containing $k^{*}$ such that $\mathbb{M}/k^{*}$ is a complement of $(D\times k^{*})/k^{*}$ in $\hat{E}_{\mathfrak{h}}/k^{*}$. Set $\mathbb{B}$ to be the $\mathcal{O}$-submodule of $\mathbb{A}_{\delta}$ generated by $D\rtimes\mathbb{M}$. It is routine to check that $\mathbb{B}$ is a unitary subalgebra containing $\mbox{$\mathcal{O}$}D$ of $\mathbb{A}_{\delta}$ which is isomorphic to $\mbox{$\mathcal{O}$}_{\alpha}(D\rtimes E_{\mathfrak{h}})$ as $D$-interior algebras for the element $\alpha\in H^{2}(E_{\mathfrak{h}};k^{*})$ determined the second short exact sequence above and isomorphic to $\mathop{\bigoplus}\limits_{gDC_{G}(D)\in E_{\mathfrak{h}}}\mbox{$\mathcal{O}$}D\mathfrak{a}_{g}$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. Therefore, $\mathrm{Br}_{D}(\mathbb{B})$ is isomorphic to $kZ(D)$. By [21, Lemma 5.5 and Proposition 5.15], there is a finite subgroup $\mathbb{M}^{\prime}$ of $\mathbb{M}$ such that $\mathbb{B}$ is isomorphic to $\mbox{$\mathcal{O}$}(D\rtimes\mathbb{M}^{\prime})\mathfrak{b}^{\prime}$ for some central idempotent $\mathfrak{b}^{\prime}$ of $\mbox{$\mathcal{O}$}(D\rtimes\mathbb{M}^{\prime})$ as $D$-interior algebras. Then $\mathrm{Br}_{D}(\mbox{$\mathcal{O}$}(D\rtimes\mathbb{M}^{\prime})\mathfrak{b}^{\prime})$ is isomorphic to the local $k$-algebra $kZ(D)$ which implies $\mathfrak{b}^{\prime}$ is a block idempotent since $D$ is normal in $D\rtimes\mathbb{M}^{\prime}$. Hence, by [17, Proposition 6.7.15], the inclusion map $\mathbb{B}$ to $\mathbb{A}_{\delta}$ splits as $\mathbb{B}$-$\mathbb{B}$-bimodules. We are done. ∎
By the decomposition (4.6.1), we can directly get the information about the rank of $\mathbb{A}_{\delta}$ and the dimension of simple modules as follows.
Corollary 4.7.
We have $\frac{\mathrm{rank}_{\mathcal{O}}(\mathbb{A}_{\delta})}{|D|}\equiv|E_{\mathfrak{h}}|(\mathrm{mod}~|D|)$. In particular, there is at least one simple $k\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$-module of dimension coprime to $p$.
Another consequence of Proposition 4.6, we can show that a special automorphism of $\mathbb{A}_{\delta}$ is an inner automorphism. This result is similar to [21, Proposition 14.9]
Corollary 4.8.
With the notation above, let $f$ be an $\mathcal{O}$-algebra automorphism of $\mathbb{A}_{\delta}$ such that its restriction to the subalgebra $\mathbb{B}$ is trivial. Then $f$ is an inner automorphism induced by an element of $C_{\mathbb{A}_{\delta}}(\mathbb{B})^{*}$.
Proof.
Since $\mathbb{A}_{\delta}=\mathbb{B}\oplus Y$ as $\mathbb{B}$-$\mathbb{B}$-bimodules, we have $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathbb{B}}\mathbb{A}_{\delta}$ being isomorphic to $\mathbb{A}_{\delta}\oplus Y\oplus(Y\mathop{\otimes}\limits_{\mathbb{B}}Y)$ as $\mathbb{B}$-$\mathbb{B}$-bimodules. Since $Y$ is isomorphic to $X$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules, we have $\mathrm{Br}_{D}(Y)\cong\mathrm{Br}_{D}(X)=0$. For $Y\mathop{\otimes}\limits_{\mathbb{B}}Y$, it is clear that it is isomorphic to a direct summand of $Y\mathop{\otimes}\limits_{\mathcal{O}D}Y$ as $\mathbb{B}$-$\mathbb{B}$-bimodules and then it is isomorphic to $X\mathop{\otimes}\limits_{\mathcal{O}D}X$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. By the same argument in the proof of [10, Lemma 3.3], we have $\mathrm{Br}_{D}(X\mathop{\otimes}\limits_{\mathcal{O}D}X)=0$. So is $\mathrm{Br}_{D}(Y\mathop{\otimes}\limits_{\mathbb{B}}Y)$. Since $\mathbb{B}$ contains $\mbox{$\mathcal{O}$}D$ and $\mathbb{A}_{\delta}$ is relatively $\mbox{$\mathcal{O}$}D$-separable, it is easy to check that $\mathbb{A}_{\delta}$ is also relatively $\mathbb{B}$-separable. In particular, $\mathbb{A}_{\delta}$ is isomorphic to a direct summand of $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathbb{B}}\mathbb{A}_{\delta}$ as $\mathbb{A}_{\delta}$-$\mathbb{A}_{\delta}$-bimodules. Then $\mathbb{A}_{\delta}$ is the unique direct summand of the $\mathbb{A}_{\delta}$-$\mathbb{A}_{\delta}$-bimodule $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathbb{B}}\mathbb{A}_{\delta}$, up to isomorphisms, with respect to the Brauer quotients at $D$ being nonzero.
On the other hand, since $\mathbb{A}_{\delta}$ is relatively $\mathbb{B}$-separable, then the multiplication map $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathbb{B}}\mathbb{A}_{\delta}\longrightarrow\mathbb{A}_{\delta}$ splits. This means that there is an element $\sum\limits_{l}\mathfrak{a}_{l}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}\mathfrak{a}^{\prime}_{l}$ of $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathbb{B}}\mathbb{A}_{\delta}$ such that $\sum\limits_{l}\mathfrak{a}_{l}\mathfrak{a}^{\prime}_{l}=j$ and $\sum\limits_{l}\mathfrak{a}\mathfrak{a}_{l}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}\mathfrak{a}^{\prime}_{l}=\sum\limits_{l}\mathfrak{a}_{l}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}\mathfrak{a}^{\prime}_{l}\mathfrak{a}$ for any $\mathfrak{a}\in\mathbb{A}_{\delta}$. Let $(\mathbb{A}_{\delta})_{f}$ be the $\mathbb{A}_{\delta}$-$\mathbb{A}_{\delta}$-bimodule such that it equals $\mathbb{A}_{\delta}$ as left $\mathbb{A}_{\delta}$-modules with $f(\mathfrak{a})$ acting by right multiplication on $\mathfrak{a}^{\prime}\in(\mathbb{A}_{\delta})_{f}$ as $\mathfrak{a}$ on $\mathfrak{a}^{\prime}\in\mathbb{A}_{\delta}$. With the same arguments in the proof of [10, Proposition 3.11], $(\mathbb{A}_{\delta})_{f}$ is isomorphic to a direct summand of $(\mathbb{A}_{\delta})_{f}\mathop{\otimes}\limits_{\mathbb{B}}\mathbb{A}_{\delta}$ as $\mathbb{A}_{\delta}$-$\mathbb{A}_{\delta}$-bimodules. Since the restriction of $f$ to $\mathbb{B}$ is trivial, it is easy to check that $(\mathbb{A}_{\delta})_{f}\mathop{\otimes}\limits_{\mathbb{B}}\mathbb{A}_{\delta}$ is isomorphic to $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathbb{B}}\mathbb{A}_{\delta}$ as $\mathbb{A}_{\delta}$-$\mathbb{A}_{\delta}$-bimodules. Obviously, $(\mathbb{A}_{\delta})_{f}$ is isomorphic to $\mathbb{A}_{\delta}$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. By the uniqueness of $\mathbb{A}_{\delta}$, $(\mathbb{A}_{\delta})_{f}$ is also isomorphic to $\mathbb{A}_{\delta}$ as $\mathbb{A}_{\delta}$-$\mathbb{A}_{\delta}$-bimodules. The image of $j$ under this isomorphism is the element we want. ∎
4.9.
Now we assume that the block $b$ is a hyperfocal abelian Frobenius block. In the following, we will give a more explicit description of the $\mbox{$\mathcal{O}$}(D\times D)$-module structure of $X$ in the decomposition (4.6.1) under this assumption. To do this, we need to investigate the Brauer quotients of $\mathbb{A}_{\delta}$ at a class of special nontrivial subgroups of $D$.
Lemma 4.10.
Keep the assumption as above. Let $Q$ be a nontrivial subgroup of $D$. Assume that $C_{P}(Q)$ is a defect group of the block $b_{Q}$ of $C_{G}(Q)$. Then there is an indecomposable endopermutation $k(D/Q)$-module $V_{Q}$ with vertex $D/Q$ such that setting $\bar{S}_{Q}=\mathrm{End}_{k}(V_{Q})$, we have the following $D$-interior algebra isomorphism
$A_{\delta}(Q)\cong\bar{S}_{Q}\mathop{\otimes}\limits_{k}kC_{P}(Q)$ (4.10.1)
with its restriction to $\mathbb{A}_{\delta}(Q)$ giving the following $D$-interior algebra isomorphism
$\mathbb{A}_{\delta}(Q)\cong\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD.$ (4.10.2)
In particular, $(j\mathbb{A}vj)(Q)\neq 0$ if and only if $v$ is in $C_{P}(Q)$.
Proof.
Denote $N_{P}(Q)$ by $P_{Q}$. It is clear that $b_{Q}$ is still a block of $P_{Q}C_{G}(Q)$ over $\mathcal{O}$. We denote it by $c_{Q}$ when the block $b_{Q}$ is viewed as a block of $P_{Q}C_{G}(Q)$. By assumption, $P_{Q}$ is a defect group of the block $c_{Q}$. Since the block $b_{Q}$ is nilpotent by Proposition 3.12, the block $c_{Q}$ of $P_{Q}C_{G}(Q)$ is also nilpotent. Since $P_{Q}$ contains $D$, there is a unique local pointe $\varepsilon$ of $P_{Q}$ on $\mbox{$\mathcal{O}$}Gb$ such that $(P_{Q})_{\varepsilon}\leq P_{\gamma}$ by [25, Proposition 4.2]. Now we can take an element $l$ of $\varepsilon$ such that $l$ is in $\mathbb{A}$ by [25, Proposition 2.8] and $lj=jl=j$.
By [29, Lemma (40.2)], $\mathrm{Br}_{Q}(l)$ belongs to a local point of $P_{Q}$ on $kC_{G}(Q)$. Since $P_{Q}$ is a defect group of the block $c_{Q}$, $\mathrm{Br}_{Q}(l)kP_{Q}C_{G}(Q)\mathrm{Br}_{Q}(l)$ is a source algebra of the block $\bar{c}_{Q}$ over $k$. Since the block $c_{Q}$ is nilpotent, by [17, Theorem 8.12.3], there is a an indecomposable endopermutation $kP_{Q}$-module $V_{P_{Q}}$ with vertex $P_{Q}$ such that, setting $\bar{T}_{Q}=\mathrm{End}_{k}(V_{P_{Q}})$, we have an isomorphism of $P_{Q}$-interior algebras
$\mathrm{Br}_{Q}(l)kP_{Q}C_{G}(Q)\mathrm{Br}_{Q}(l)\cong\bar{T}_{Q}\mathop{\otimes}\limits_{k}kP_{Q}$
which restricts to a $C_{P}(Q)$-interior algebra isomorphism
$\mathrm{Br}_{Q}(l)kC_{G}(Q)\mathrm{Br}_{Q}(l)\cong\bar{T}_{Q}\mathop{\otimes}\limits_{k}kC_{P}(Q).$ (4.10.3)
Since $Q$ is normal in $P_{Q}C_{G}(Q)$, it is clear that the action of $Q$ on $V_{P_{Q}}$ is trivial. Then $V_{P_{Q}}$ can be viewed as a $k(P_{Q}/Q)$-module. Note that $\mathrm{Br}_{Q}(l)kC_{G}(Q)\mathrm{Br}_{Q}(l)=(lAl)(Q)$ and $(l\mathbb{A}l)(Q)$ is a unitary subalgebra of $(lAl)(Q)$. By the proof of [25, Proposition 4.3], the restriction of the isomorphism (4.10.3) to $(l\mathbb{A}l)(Q)$ gives a $D$-interior algebra isomorphism
$(l\mathbb{A}l)(Q)\cong\bar{T}_{Q}\mathop{\otimes}\limits_{k}kD.$ (4.10.4)
It is well-known that there is a unique local point,denoted by $\tilde{\delta}$, of $D$ on $\bar{T}_{Q}$. Take $f\in\tilde{\delta}$ and set $V_{Q}=fV_{P_{Q}}$. Then $V_{Q}$ is an indecomposable endopermutation $kD$-module with vertex $D$. Clearly, the action of $Q$ on $V_{Q}$ is also trivial. So $V_{Q}$ can be viewed as a $k(D/Q)$-module with vertex $D/Q$. Set $\bar{S}_{Q}=\mathrm{End}_{k}(V_{Q})\cong f\bar{T}_{Q}f$. By [17, Proposition 7.4.4], the $D$-interior algebras $\bar{S}_{Q}\mathop{\otimes}\limits_{k}kC_{P}(Q)$ and $\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD$ are both primitive. Note that $\mathrm{Br}_{Q}:\mathbb{A}^{D}\rightarrow\mathbb{A}(Q)^{D/Q}$ is surjective and $\mathrm{Br}_{D/Q}\circ\mathrm{Br}_{Q}$ is just the Brauer homomorphism $\mathrm{Br}_{D}$ from $\mathbb{A}^{D}$ to $\mathbb{A}(D)$. Then $\mathrm{Br}_{Q}(j)$ belongs to a local point of $D$ on $((l\mathbb{A}l)(Q))^{D}$. Comparing the isomorphism (4.10.3) with the isomorphism (4.10.4), we get $(lAl)(Q)=\mathop{\bigoplus}\limits_{uD\in C_{P}(Q)/D}(l\mathbb{A}l)(Q)\cdot u$. Therefore, by [25, Proposition 2.8], $\mathrm{Br}_{Q}(j)$ also belongs to a local point of $D$ on $((lAl)(Q))^{D}$. Note that by [22, 5.3], local points of $D$ on $(lAl)(Q)$ and $(l\mathbb{A}l)(Q)$ are both unique. Therefore, we can assume that the isomorphism (4.10.3) maps $\mathrm{Br}_{Q}(j)$ to $f\mathop{\otimes}\limits_{k}1_{C_{P}(Q)}$. Then the isomorphisms (4.10.1) and (4.10.2) are obtained.
Borrowing the notation from the paragraph 3.1, we have $A_{\delta}=\mathop{\bigoplus}\limits_{uD\in P/D}\mathbb{A}_{\delta}a_{u,\delta}=\mathop{\bigoplus}\limits_{uD\in P/D}j\mathbb{A}uj$ with $a_{u,\delta}\in A_{\delta}^{*}\cap\mathbb{A}u$ satisfying $a_{u,\delta}(wj)a_{u,\delta}^{-1}=(uwu^{-1})j$ for any $w\in D$. So $(j\mathbb{A}uj)(Q)=\mathbb{A}_{\delta}(Q)\cdot\mathrm{Br}_{Q}(a_{u,\delta})$ when $u$ is in $C_{P}(Q)$. In particular, $\mathrm{dim}_{k}(\mathbb{A}_{\delta}(Q))=\mathrm{dim}_{k}((j\mathbb{A}uj)(Q))$ when $u$ is in $C_{P}(Q)$. By isomorphisms (4.10.3) and (4.10.4), we have $\mathrm{dim}_{k}(A_{\delta}(Q))=\frac{|C_{P}(Q)|}{|D|}\mathrm{dim}_{k}(\mathbb{A}_{\delta}(Q))$. We are done. ∎
In general, the strucutre of $\mathbb{A}_{\delta}(Q)$ above holds for arbitrary nontrivial subgroup $Q$ of $D$. And this general structure will help us construct a stable equivalence of Morita type between $\mathbb{A}_{\delta}$ and $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$ in the next section.
Proposition 4.11.
For any nontrivial subgroup $Q$ of $D$, there is an indecomposable endopermutation $k(D/Q)$-module $V_{Q}$ with vertex $D/Q$ such that, setting $\bar{S}_{Q}=\mathrm{End}_{k}(V_{Q})$, we have a $D$-interior algebra isomorphism
$\mathbb{A}_{\delta}(Q)\cong\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD.$ (4.11.1)
Moreover, if $V_{Q}^{\prime}$ is another indecomposable endopermutation $k(D/Q)$-module with vertex $D/Q$ such that setting $\bar{S}_{Q}^{\prime}=\mathrm{End}_{k}(V_{Q}^{\prime})$, $\mathbb{A}_{\delta}(Q)\cong\bar{S}_{Q}^{\prime}\mathop{\otimes}\limits_{k}kD$ as $D$-interior algebras, then $V_{Q}$ is isomorphic to $V_{Q}^{\prime}$ as $kD/Q$-modules.
Proof.
It is well-known that there is some $x\in G$ such that ${{}^{x}}(Q,\bar{b}_{Q})\leq(P,\bar{b}_{P})$ and $C_{P}({{}^{x}}Q)$ is a defect group of the block ${{}^{x}}b_{Q}$ of $C_{G}({{}^{x}}Q)$ over $\mathcal{O}$. Equivalently, $C_{P^{x}}(Q)$ is a defect group of the block $b_{Q}$. Since $D$ is abelian, by [33, Theomre 2], we can assume that $x$ is in $N_{G}(D_{\delta})$. Furthermore, since $N_{G}(D_{\delta})/C_{G}(D)=PC_{G}(D)/C_{G}(D)\cdot E_{\mathfrak{h}}$, we can assume that $xC_{G}(D)$ is in $E_{\mathfrak{h}}$. Then we have $\mathfrak{a}_{x}\in\mathbb{A}_{\delta}^{*}$ such that $\mathfrak{a}_{x}(wj)\mathfrak{a}_{x}^{-1}=(xwx^{-1})j$ for any $w\in D$.
Clearly, $A^{x}$ is also a source algebra of the block $b$ determined by the defect pointed group $P_{\gamma}^{x}$. Since $x$ is in $N_{G}(D_{\delta})$, we have $\mathbb{A}^{x}$ is a hyperfocal subalgebra of $A^{x}$ with hyperfocal subgroup $D$. The conjugation action of $x$ induces an $\mathcal{O}$-algebra isomorphism $\mathbb{A}_{\delta}^{{{}^{x}}Q}\longrightarrow(\mathbb{A}^{x})_{\delta^{x}}^{Q}$. At the same time, the conjugation action of $\mathfrak{a}_{x}$ induces an $\mathcal{O}$-algebra isomorphism $\mathbb{A}_{\delta}^{Q}\longrightarrow\mathbb{A}_{\delta}^{{{}^{x}}Q}$. Composing these two isomorphisms together, we get an $\mathcal{O}$-algebra isomorphism $\mathbb{A}_{\delta}^{Q}\longrightarrow(\mathbb{A}^{x})_{\delta^{x}}^{Q}$, which is in fact a $D$-interior algebra isomorphism. By Lemma 4.10, we get this general structure of $\mathbb{A}_{\delta}(Q)$.
Let $V_{Q}^{\prime}$ be another indecomposable endopermutation $k(D/Q)$-module with vertex $D/Q$ fulfilling the property in this lemma. Then we have $\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD\cong\bar{S}_{Q}^{\prime}\mathop{\otimes}\limits_{k}kD$ as $D$-interior algebras. Tensoring with $\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}$- on the both sides of this isomorphism yields a $D$-interior algebra isomorphism $\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD\cong\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}^{\prime}\mathop{\otimes}\limits_{k}kD$. By [22, 5.3], local points of $D$ on $\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}^{\prime}$ and $\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD$ are both unique. We take two elements $e^{\prime}$ and $e$ belonging to these two unqiue local points, respectively. Then by [17, Theorem 7.4.2], $e(\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD)e$ is isomorphic to $kD$ as $D$-interior algebras. It is clear that $e^{\prime}(\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}^{\prime})e^{\prime}$ is also a primitive Dade $D$-interior algebra. By [17, Proposition 7.4.4], we have $e^{\prime}(\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}^{\prime})e^{\prime}\mathop{\otimes}\limits_{k}kD$ is a primitive $D$-interior algebra. Hence, we have $e^{\prime}(\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}^{\prime})e^{\prime}\mathop{\otimes}\limits_{k}kD\cong kD$ as $D$-interior algebras. This implies that $e^{\prime}(\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}^{\prime})e^{\prime}$ is a trivial $D$-interior algebra. Then the elements in the Dade group of $D$ corresponding to $\bar{S}_{Q}$ and $\bar{S}_{Q}^{\prime}$ are inverses of each other. In particular, $V_{Q}\cong V_{Q}^{\prime}$ as $k(D/Q)$-modules. ∎
Now we can refine the $\mbox{$\mathcal{O}$}(D\times D)$-structure of $\mathbb{A}_{\delta}$ in the case where the block $b$ is a hyperfocal abelian Frobenius block.
Proposition 4.12.
Assume that the block $b$ is a hyperfocal abelian Frobenius block. Then every direct summand of the $\mbox{$\mathcal{O}$}(D\times D)$-module $X$ appearing in the decomposition (4.6.1) has the form $\mathrm{Ind}_{\Delta_{\varphi}(Q)}^{D\times D}(\mbox{$\mathcal{O}$})$ with $Q$ being a proper subgroup of $D$ and $\varphi:Q\longrightarrow D$ induced by an element of $E_{\mathfrak{h}}$. Moreover, there is an injective group homomorphism from $E_{\mathfrak{h}}$ to $N_{\mathbb{A}_{\delta}}(D)$ sending $\tilde{g}=gC_{G}(D)$ to $\mathfrak{e}_{\tilde{g}}$ satisfying that $(gwg^{-1})j=\mathfrak{e}_{\tilde{g}}(wj)\mathfrak{e}_{\tilde{g}}^{-1}$ for any $w\in D$. Identifying $E_{\mathfrak{h}}$ with a subgroup of $N_{\mathbb{A}_{\delta}}(D)$ through this homomorphism, then $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$ is a unitary subalgebra of $\mathbb{A}_{\delta}$ that is stable under the conjugation of $a_{u,\delta}$ and $\mathbb{A}_{\delta}=\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})\oplus Y$ as $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$-$\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$-bimodules with the restriction of $Y$ to $\mbox{$\mathcal{O}$}D$-$\mbox{$\mathcal{O}$}D$ being isomorphic to $X$.
Proof.
By Proposition 4.6, every direct summand of the $\mbox{$\mathcal{O}$}(D\times D)$-module $X$ has the form $\mbox{$\mathcal{O}$}DhD$ for some $h\in G-N_{G}(D)$. Set $Q=D\cap{{}^{h}}D$. So $Q$ is a proper subgroup of $D$. Moreover, we can assume that $Q$ is nontrivial. For in this case, we can set $\varphi=\mathrm{id}$. Let $\varphi_{h}$ be the group homomorphism from $Q$ to $D$, sending $v$ to $hvh^{-1}$. Then as $\mbox{$\mathcal{O}$}(D\times D)$-modules, $\mbox{$\mathcal{O}$}DhD$ is isomorphism to $\mathrm{Ind}_{\Delta_{\varphi_{h}}(Q)}^{D\times D}(\mbox{$\mathcal{O}$})$, which is isomorphic to $\mbox{$\mathcal{O}$}D\mathop{\otimes}\limits_{\mathcal{O}Q}{{}_{\varphi_{h}}}\mbox{$\mathcal{O}$}D$. By [17, Theorem 8.7.1], we have ${{}^{h}}(Q,\bar{b}_{Q})\leq(D,\bar{b}_{D})$. Since $D$ is abelian, by [33, Theorem 2], we can assume that $h=gz$ for some $g\in N_{G}(D_{\delta})=N_{G}(D,b_{D})$ and some $z\in C_{G}(Q)$. Therefore, $\varphi_{h}=\varphi_{g}:Q\rightarrow D,~~u\mapsto gug^{-1}$. Furthermore, since $N_{G}(D_{\delta})/C_{G}(D)=PC_{G}(D)/C_{G}(D)\cdot E_{\mathfrak{h}}$, we can write $g$ as $vx$ for some $v\in P$ and some $x\in N_{G}(D_{\delta})$ with $xC_{G}(D)\in E_{\mathfrak{h}}$. By the proof of Proposition 4.6, we can deduce that $\mathbb{A}_{\delta}$ has a direct summand isomorphic to $\mathrm{Ind}_{\Delta_{\varphi_{v}}(Q)}^{D\times D}(\mbox{$\mathcal{O}$})$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. Here, $\varphi_{v}$ is the group homomorphism from $Q$ to $D$, sending $u$ to $vuv^{-1}$. Equivalently, there is an element $\mathfrak{a}_{v}$ in $\mathbb{A}_{\delta}$ such that for any $(w,w^{\prime})\in D\times D$, $w\mathfrak{a}_{v}(w^{\prime})^{-1}=\mathfrak{a}_{v}$ if and only if $w^{\prime}=vwv^{-1}$ and $w\in Q$. At the same time, by the proof of Proposition 4.11, there is an element $y$ in $N_{G}(D_{\delta})$ such that $yC_{G}(D)$ is in $E_{\mathfrak{h}}$ and $C_{P}({{}^{y}}Q)$ is a defect group of the block $b_{{{}^{y}}Q}$ of $C_{G}({{}^{y}}Q)$. For this element $y$, there is an element $\mathfrak{a}_{y}\in\mathbb{A}_{\delta}^{*}$ such that $\mathfrak{a}_{y}(wj)\mathfrak{a}_{y}^{-1}=(ywy^{-1})j$ for any $w\in D$. Since $E_{\mathfrak{h}}$ is normal in $N_{G}(D_{\delta})/C_{G}(D)$, $(vy^{-1}v^{-1})C_{G}(D)$ still belongs to $E_{\mathfrak{h}}$. Hence, we have an element $\mathfrak{a}_{vy^{-1}v^{-1}}\in\mathbb{A}_{\delta}^{*}$ such that $\mathfrak{a}_{vy^{-1}v^{-1}}(wj)\mathfrak{a}_{vy^{-1}v^{-1}}^{-1}=(vy^{-1}v^{-1}wvyv^{-1})j$ for any $w\in D$. Then we can check that the $\mbox{$\mathcal{O}$}(D\times D)$-module generated by $\mathfrak{a}_{y}\mathfrak{a}_{v}\mathfrak{a}_{vy^{-1}v^{-1}}$ is isomorphic to $\mathrm{Ind}_{\Delta_{\varphi_{v}}({{}^{y}}Q)}^{D\times D}(\mbox{$\mathcal{O}$})$. This implies that the $\mbox{$\mathcal{O}$}(D\times D)$-module $\mathbb{A}_{\delta}$ has a direct summand isomorphic to $\mathrm{Ind}_{\Delta_{\varphi_{v}}({{}^{y}}Q)}^{D\times D}(\mbox{$\mathcal{O}$})$. Without loss of generality, we can assume that $C_{P}(Q)$ is a defect group of the block $b_{Q}$ of $C_{G}(Q)$.
Borrowing the notation from the paragraph 3.1, there is an element $a_{v,\delta}$ in $A_{\delta}^{*}\cap\mathbb{A}v$ such that $a_{v,\delta}(uj)a_{v,\delta}^{-1}=(vuv^{-1})\cdot j$. Then $\mathfrak{a}_{v}a_{v,\delta}$ is in $j\mathbb{A}vj$ such that $w\mathfrak{a}_{v}a_{v,\delta}(w^{\prime})^{-1}=\mathfrak{a}_{v}a_{v,\delta}$ if and only if $w^{\prime}=w$ and $w\in Q$ for any $(w,w^{\prime})\in D\times D$. In particular, there is a $\mbox{$\mathcal{O}$}(D\times D)$-submodule of $j\mathbb{A}vj$ isomorphic to $\mathrm{Ind}_{\Delta(Q)}^{D\times D}(\mbox{$\mathcal{O}$})$. This implies that $(j\mathbb{A}vj)(Q)\neq 0$. By Lemma 4.10, we have $v$ being in $C_{P}(Q)$. Then the forms of indecomposable direct summand of $X$ follow from this fact.
It is well-known that $N_{\mathbb{A}_{\delta}}(D)/(\mathbb{A}_{\delta}^{D})^{*}$ is isomorphic to a subgroup of $N_{A_{\delta}}(D)/(A_{\delta}^{D})^{*}$, which is isomorphic to $N_{G}(D_{\delta})/C_{G}(D)$. By the proof of Lemma 3.7, we have $N_{\mathbb{A}_{\delta}}(D)/(\mathbb{A}_{\delta}^{D})^{*}$ is in fact isomorphic to $E_{\mathfrak{h}}$. We identify these two groups with each other. At the same time, we have the following short exact sequence
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${k^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N_{\mathbb{A}_{\delta}}(D)/(j+J(\mathbb{A}_{\delta})^{D})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{\mathfrak{h}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
Since $D\rtimes E_{\mathfrak{h}}$ is a Frobenius group, we have $H^{2}(E_{\mathfrak{h}};k^{*})=1$. In particular, the short exact sequence above splits. So there is a subgroup $\hat{E}_{\mathfrak{h}}$ of $N_{\mathbb{A}_{\delta}}(D)$ containing $j+J(\mathbb{A}_{\delta})^{D}$ such that $k^{*}\cap\hat{E}_{\mathfrak{h}}=j$ and $\hat{E}_{\mathfrak{h}}/(j+J(\mathbb{A}_{\delta})^{D})\cong E_{\mathfrak{h}}$. In particular, we get another short exact sequence
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${j+J(\mathbb{A}_{\delta})^{D}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\hat{E}_{\mathfrak{h}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{\mathfrak{h}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
Since $E_{\mathfrak{h}}$ is a $p^{\prime}$-group, by [29, Lemma (45.6)], this sequence also splits. Therefore, we can view $E_{\mathfrak{h}}$ as a subgroup $\{\mathfrak{e}_{\tilde{g}}\mid\tilde{g}\in E_{\mathfrak{h}}\}$ of $N_{\mathbb{A}_{\delta}}(D)$ intersecting with $(\mathbb{A}_{\delta}^{D})^{*}$ trivially and acting on $D$ as same as $E_{\mathfrak{h}}$ acting on $D$.
As $\mbox{$\mathcal{O}$}(D\times D)$-modules, by Lemma 4.5, $\mathbb{A}_{\delta}=(\mathop{\bigoplus}\limits_{\tilde{g}\in E_{\mathfrak{h}}}\mbox{$\mathcal{O}$}D\mathfrak{e}_{\tilde{g}})\oplus X$ for some $\mbox{$\mathcal{O}$}(D\times D)$-submodule $X$. Clearly, $\mathop{\bigoplus}\limits_{\tilde{g}\in E_{\mathfrak{h}}}\mbox{$\mathcal{O}$}D\mathfrak{e}_{\tilde{g}}$ is a unitary subalgebra of $\mathbb{A}_{\delta}$ isomorphic to $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$. Then by [17, Proposition 6.15.2], the inclusion map $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})\hookrightarrow\mathbb{A}_{\delta}$ splits as a homomorphism of $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$-$\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$-bimodules. Therefore, $\mathbb{A}_{\delta}=\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})\oplus Y$ for some $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$-$\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$-bimodule $Y$. By the Krull-Schmidt Theorem, $Y$ is isomorphic to $X$ as $\mbox{$\mathcal{O}$}D$-$\mbox{$\mathcal{O}$}D$-bimodules. ∎
5 A stable equivalence of Morita type
We still use the notation in the last two sections. Throughout this section, we assume that the block is a hyperfocal abelian Frobenius block. The main result in this section is the construction of a stable equivalence of Morita type between $\mathbb{A}_{\delta}$ and $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$ under this assumption. To accomplish this, we need to glue the indecomposable endopermutation $k(D/Q)$-modules $V_{Q}$ for any $1\neq Q\leq D$ appearing in Proposition 4.11 together. The main technique is the result due to Puig [23, Proposition 3.6] (see also [17, Theorem 7.8.2]). For convenience, we present this result here.
Lemma 5.1.
([23, Proposition 3.6]) Let $Z$ be a finite abelian $p$-group and $E$ a subgroup of $\mathrm{Aut}(Z)$. Let $\mathcal{C}$ be a nonempty upwardly closed $E$-stable set of subgroups of $Z$, which means that if $Q\leq R\leq Z$ with $Q$ in $\mathcal{C}$, then $R$ and $y(Q)$ belong to $\mathcal{C}$ fro all $y\in E$. For any subgroup $Q$ in $\mathcal{C}$ let $V_{Q}$ be an indecomposable endopermutation $k(Z/Q)$-module with vertex $Z/Q$. Suppose that if $Q,R$ belong to $\mathcal{C}$ such that $Q\leq R$, then $\mathrm{Defres}_{Z/R}^{Z/Q}(V_{Q})\cong V_{R}$, where $Z/R$ is identified with $(Z/Q)/(R/Q)$. Then there is an indecomposable endopermutation $kZ$-module $V$ with vertex $Z$ such that $\mathrm{Defres}_{Z/Q}^{Z}(V)\cong V_{Q}$ for any subgroup $Q$ in $\mathcal{C}$. In addition, if ${{}_{y}}(V_{Q})\cong V_{y(Q)}$ as $k(Z/y(Q))$-modules for all $y\in E$ and all $Q$ in $\mathcal{C}$, then ${{}_{y}}V\cong V$ as $kZ$-modules for all $y\in E$.
5.2.
In our setting, we can identify $N_{G}(D_{\delta})/C_{G}(D)$ with a subgroup of $\mathrm{Aut}(D)$. Denote $N_{G}(D_{\delta})/C_{G}(D)$ by $\tilde{N}$ and the image of $g\in N_{G}(D_{\delta})$ in $\tilde{N}$ by $\tilde{g}$. Let $\mathcal{C}$ be the set of subgroups of $D$ consisting of all nontrivial subgroups of $D$. Obviously, $\mathcal{C}$ is a nonempty upwardly $\tilde{N}-$stable set. Take two subgroups $Q$ and $R$ of $D$ with $Q\leq R$. We denote $\mathrm{Defres}_{D/R}^{D/Q}(V_{Q})$ by $W_{Q,R}$. Clearly, $W_{Q,R}$ is up to isomorphism a unique endopermutation $k(D/R)$-module with vertex $D/R$ such that $\bar{S}_{Q}(R)$ is isomorphic to $\mathrm{End}_{k}(W_{Q,R})$ as $D/R$-algebras. Here, $\bar{S}_{Q}$ is $\mathrm{End}_{k}(V_{Q})$.
Lemma 5.3.
Keep the notation as above. We have $W_{Q,R}\cong V_{R}$ as $k(D/R)$-modules. Moreover, for all $\tilde{g}\in\tilde{N}$, we have ${{}_{g}}V_{Q}\cong V_{{{}^{g}}Q}$ as $kD$-modules.
Proof.
By Proposition 4.11, we have $\mathbb{A}_{\delta}(Q)\cong\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD$ and $\mathbb{A}_{\delta}(R)\cong\bar{S}_{R}\mathop{\otimes}\limits_{k}kD$ as $D$-interior algebras. Since $D$ is abelian and $Q\leq R$, we have $\mathbb{A}_{\delta}(R)\cong(\mathbb{A}_{\delta}(Q))(R)$ as $D$-interior algebras. It is clear that $(\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD)(R)\cong\bar{S}_{Q}(R)\mathop{\otimes}\limits_{k}(kD)(R)\cong\bar{S}_{Q}(R)\mathop{\otimes}\limits_{k}kD$. Then we have $\bar{S}_{Q}(R)\mathop{\otimes}\limits_{k}kD\cong\bar{S}_{R}\mathop{\otimes}\limits_{k}kD$ as $D$-interior algebras. By the uniqueness of $V_{R}$, we have $W_{Q,R}\cong V_{R}$ as $k(D/R)$-modules since $\bar{S}_{Q}(R)\cong\mathrm{End}_{k}(W_{Q,R})$.
Take an element $\tilde{g}$ of $\tilde{N}$. Clearly, $\mathrm{End}_{k}(_{g}V_{Q})\cong{{}_{g}}\bar{S}_{Q}$. Here, ${{}_{g}}\bar{S}_{Q}$ is the $D/{{}^{g}}Q$-interior algebra with being equal to $\bar{S}_{Q}$ as $k$-algebra and the group homomorphism from $D/{{}^{g}}Q$ to $({{}_{g}}\bar{S}_{Q})^{*}$ being just the composition of the group homomorphism from $D/{{}^{g}}Q$ to $D/Q$ induced by the conjugation of $g$ and the structural homomorphism from $D/Q$ to $\bar{S}_{Q}^{*}$. We use the similar notation for any $D$-interior algebra. Since $\mathbb{A}_{\delta}(Q)\cong\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD$ as $D$-interior algebras, we have
${{}_{g}}(\mathbb{A}_{\delta}(Q))\cong{{}_{g}}\bar{S}_{Q}\mathop{\otimes}\limits_{k}{{}_{g}}kD$ (5.3.1)
as $D$-interior algebras. On the other hand, obviously, we have
${{}_{g}}(\mathbb{A}_{\delta}(Q))\cong({{}_{g}}\mathbb{A}_{\delta})({{}^{g}}Q)$ (5.3.2)
as $D$-interior algebras. For $\tilde{g}\in\tilde{N}$, it can be written as $ux$ for some $u\in P$ and some $x\in N_{G}(D_{\delta})$ with $\tilde{x}\in E_{\mathfrak{h}}$. Borrowing the notation from the paragraph 3.1 and Proposition 4.12, we have $a_{u,\delta}\in A_{\delta}^{*}\cap j\mathbb{A}uj$ and $\mathfrak{e}_{x}\in N_{\mathbb{A}_{\delta}}(D)$ such that $a_{u,\delta}(wj)a_{u,\delta}^{-1}=(uwu^{-1})j$ and $\mathfrak{e}_{x}(wj)\mathfrak{e}_{x}^{-1}=(xwx^{-1})j$ for any $w\in D$. Denote $a_{u,\delta}\mathfrak{e}_{x}$ by $a_{ux,\delta}$. Then $a_{g,\delta}$ is in $N_{A_{\delta}}(D)\cap j\mathbb{A}uj$ and $a_{g,\delta}(wj)a_{g,\delta}^{-1}=(gwg^{-1})j$ for any $w\in D$. In particular, the conjugation action of $a_{g,\delta}^{-1}$ on $\mathbb{A}_{\delta}$ induces a $D$-interior algebra isomorphism $\mathbb{A}_{\delta}\cong{{}_{g}}\mathbb{A}_{\delta}$. Therefore,
$({{}_{g}}\mathbb{A}_{\delta})({{}^{g}}Q)\cong\mathbb{A}_{\delta}({{}^{g}}Q)\cong\bar{S}_{{{}^{g}}Q}\mathop{\otimes}\limits_{k}kD$ (5.3.3)
as $D$-interior algebras. By the isomorphisms (5.3.1) and (5.3.2) and (5.3.3), we can get
$\bar{S}_{{{}^{g}}Q}\mathop{\otimes}\limits_{k}kD\cong{{}_{g}}\bar{S}_{Q}\mathop{\otimes}\limits_{k}{{}_{g}}kD$ (5.3.4)
as $D$-interior algebras. Clearly, the conjugation action of $g^{-1}$ induces a $D$-interior algebra isomorphism $kD\cong{{}_{g}}kD$. In conclusion, we obtain a $D$-interior algebra isomorphism $\bar{S}_{{{}^{g}}Q}\mathop{\otimes}\limits_{k}kD\cong{{}_{g}}\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD$. By the uniqueness of $V_{{{}^{g}}Q}$ and the isomorphism $\mathrm{End}_{k}({{}_{g}}V_{Q})\cong{{}_{g}}\bar{S}_{Q}$, we can get ${{}_{g}}V_{Q}\cong V_{{{}^{g}}Q}$ as $kD$-modules. ∎
5.4.
By Lemmas 5.1 and 5.3, we can get an indecomposable endopermutation $kD$-module $\bar{V}$ with vertex $D$ such that $(\mathrm{End}_{k}(\bar{V}))(Q)\cong\bar{S}_{Q}$ as $D$-interior algebras for any nontrivial subgroup $Q$ of $D$ and ${{}_{g}}\bar{V}\cong\bar{V}$ as $kD$-modules for any $\tilde{g}\in\tilde{N}$. Furthermore, by [30, Theorem 14.2], there is an indecomposable endopermutation $\mbox{$\mathcal{O}$}D$-module $V$ with vertex $D$ such that $k\mathop{\otimes}\limits_{\mathcal{O}}V\cong\bar{V}$. Setting $S=\mathrm{End}_{\mathcal{O}}(V)$, then $S(Q)\cong\bar{S}_{Q}$ as $D$-interior algebras for any nontrivial subgroup $Q$ of $D$. In the following, we will show that $V$ can be chose to be $\tilde{N}-$stable, namely, ${{}_{g}}V\cong V$ as $kD$-modules for any $\tilde{g}\in\tilde{N}$.
Lemma 5.5.
Keep the notation as above. Then the indecomposable endopermutation $\mbox{$\mathcal{O}$}D$-module $V$ can be chosen such that $\mathrm{det}(w\cdot 1_{S})=1$ for any $w\in D$. Here, $\mathrm{det}$ means the determinant map from $S$ to $\mathcal{O}$. Moreover, for this module $V$, we have ${{}_{g}}V\cong V$ as $\mbox{$\mathcal{O}$}D$-modules for any $\tilde{g}\in\tilde{N}$. In particular, ${{}_{u}}V\cong V$ as $kD$-modules for any $u\in P$.
Proof.
Take an element $\tilde{g}\in\tilde{N}$. Denote the structural homomorphism from $D$ to $S^{*}$ by $\phi$. Clearly, $\mathrm{End}_{\mathcal{O}}({{}_{g}}V)\cong{{}_{g}}S$ and the the structural homomorphism from $D$ to $({{}_{g}}S)^{*}$, denoted by ${{}_{g}}\phi$, mapping $w\in D$ to $\phi(gwg^{-1})$. Since $k\mathop{\otimes}\limits_{\mathcal{O}}{{}_{g}}V\cong{{}_{g}}\bar{V}$, we get $k\mathop{\otimes}\limits_{\mathcal{O}}{{}_{g}}V\cong k\mathop{\otimes}\limits_{\mathcal{O}}V$ as $kD$-modules. By [17, Proposition 7.3.12], $S$ is isomorphic to ${{}_{g}}S$ as $D$-algebras. Denote this isomorphism by $\psi$. Obviously, $\psi$ is an $\mathcal{O}$-automorphism of $S$. By the Skolem-Noether Theorem, $\psi$ is in fact an inner automorphism. In particular, $\psi$ preserves the determinants. Set $\mathrm{rank}_{\mathcal{O}}(V)=n$. By [17, Proposition 7.3.10 (i)], $n$ is coprime to $p$. Then the statements in this lemma follows from [29, Proposition (21.5)]. ∎
5.6.
Now we fix an indecomposable endopermutation $\mbox{$\mathcal{O}$}D$-module $V$ with vertex $D$ such that setting $S=\mathrm{End}_{\mathcal{O}}(V)$, $S(Q)\cong\bar{S}_{Q}$ for any nontrivial subgroup $Q$ of $D$ and $\mathrm{det}(w\cdot 1_{S})=1$ for any $w\in D$. Then by Lemma 5.5, we have ${{}_{g}}V\cong V$ as $\mbox{$\mathcal{O}$}D$-modules for any $\tilde{g}\in\tilde{N}$. In particular, $V$ is $E_{\mathfrak{h}}$-stable. By [17, Corollary 7.8.4], $V$ can be extended to an indecomposable $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$-module $U$ such that $U$ has a vertex $D$ and a source $V$ and $\mathrm{Res}_{D}^{D\rtimes E_{\mathfrak{h}}}(U)\cong V$.
5.7.
Set $\tilde{G}=D\rtimes\tilde{N}$ and $L=D\rtimes E_{\mathfrak{h}}$. Then $L\unlhd\tilde{G}$ and $D$ is a Sylow $p$-subgroup of $L$. Therefore, $P$ can act on $L$ by conjugation through the canonical map $P\longrightarrow P/C_{P}(D)$. In this setting, ${{}_{v}}U$ makes sense and is still an indecomposable $\mbox{$\mathcal{O}$}L$-module for any $v\in P$. In fact, the module $U$ can be chosen in such a way that $U$ is $P$-stable.
Lemma 5.8.
Keep the notation as above. There is an indecomposable $P$-stable $\mbox{$\mathcal{O}$}L$-module $W$ such that $\mathrm{Res}_{D}^{L}(W)\cong V$ and has a vertex $D$ and a source $V$.
Proof.
[…]
5.9.
We still denote $W$ by $V$ whenever this does not introduce any confusion. Now $S=\mathrm{End}_{\mathcal{O}}(V)$ is a $P$-stable primitive $L$-interior algebra. This means that for any $u\in P$ there exists an element $s_{u}$ in $S$ such that $s_{u}(x\cdot\mathrm{v})=(uxu^{-1})\cdot s_{u}(\mathrm{v})$ for any $x\in L$ and any $\mathrm{v}\in V$. In particular $s_{u}$ belongs to $S^{L}$ when $u$ is in $C_{P}(D)$. Since taking the inverse gives a group isomorphism from $S^{*}$ to $(S^{\circ})^{*}$, the opposite algebra $S^{\circ}$ of $S$ is also a $P$-stable primitive $L$-algebra.
5.10.
Recall from Proposition 4.12 that $\mbox{$\mathcal{O}$}L$ is a subalgebra of $\mathbb{A}_{\delta}$. So we can consider the $L$-interior algebra $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$. The following result gives the $\mbox{$\mathcal{O}$}(D\times D)$-module structure of $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$, which is similar to the one of $\mathbb{A}_{\delta}$ in Proposition 4.12.
Lemma 5.11.
There is a decomposition of $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$ as an $\mbox{$\mathcal{O}$}(D\times D)$-module
$S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}\cong(\mathop{\bigoplus}\limits_{\tilde{g}\in E_{\mathfrak{h}}}\mbox{$\mathcal{O}$}Dg)\oplus\mathbb{X},$ (5.11.1)
where $\mathbb{X}$ is isomorphic to a direct sum of modules of the form $\mathrm{Ind}_{\Delta_{\tilde{h}}(Q)}^{D\times D}(\mbox{$\mathcal{O}$})$ for some proper subgroup $Q$ of $D$ and some $\tilde{h}\in E_{\mathfrak{h}}$. In particular, $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$ has a finite $D\times D$-stable $\mathcal{O}$-basis such that $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$ is projective as a left $\mbox{$\mathcal{O}$}D$-module and as a right $\mbox{$\mathcal{O}$}D$-module.
Proof.
By Proposition 4.12, we have
$S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}\cong(\mathop{\bigoplus}\limits_{\tilde{g}\in E_{\mathfrak{h}}}\mathrm{Ind}_{\Delta_{\tilde{g}}(D)}^{D\times D}(\mathrm{Res}_{\Delta_{\tilde{g}}(D)}^{D\times D}(S^{\circ})))\oplus(S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{X}),$
where $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{X}$ is isomorphic to a direct sum of modules of the form $\mathrm{Ind}_{\Delta_{\tilde{h}}(Q)}^{D\times D}(\mathrm{Res}_{\Delta_{\tilde{h}}(Q)}^{D\times D}(S^{\circ}))$ for some proper subgroup $Q$ of $D$ and some $\tilde{h}\in E_{\mathfrak{h}}$. For any $\tilde{g}\in E_{\mathfrak{h}}$, we denote by $s^{\circ}_{\tilde{g}}$ its image in $(S^{\circ})^{*}$. Let $\Omega^{\circ}$ be a $D$-stable $\mathcal{O}$-basis of $S^{\circ}$ containing $1_{S^{\circ}}$ as the unique $D$-invariant element. Then $\Omega^{\circ}\cdot(s^{\circ}_{\tilde{g}})^{-1}$ is a $\Delta_{\tilde{g}}(D)$-stable $\mathcal{O}$-basis of $S^{\circ}$ containing $(s^{\circ}_{\tilde{g}})^{-1}$ as the unique $\Delta_{\tilde{g}}(D)$-invariant element. We are done. ∎
5.12.
Let $e$ be a primitive idempotent of $(S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})^{L}$ with $\mathrm{Br}_{D}(e)\neq 0$. Set $\mathbb{A}^{\prime}=e(S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})e$. So $\mathbb{A}^{\prime}$ is a primitive $L$-interior algebra. In fact, $e$ is still primitive in $(\mathbb{A}^{\prime})^{D}$. Indeed, by [16, Corollary 5.5.9], every point of $D$ on $\mathbb{A}^{\prime}$ is local. Since $\mathbb{A}_{\delta}$ is a primitive $D$-interior algebra with $\mathbb{A}_{\delta}(D)\neq 0$, by [17, Theorem 7.4.1 (i)], there is a unique local point of $D$ on $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$ with multiplicity $1$. Therefore, $e$ remains primitive in $(\mathbb{A}^{\prime})^{D}$. In the following, we collect some properties of $\mathbb{A}^{\prime}$.
Proposition 5.13.
For the algebra $\mathbb{A}^{\prime}$, the following hold.
(i) $\mathbb{A}^{\prime}$ is a primitive relatively $\mbox{$\mathcal{O}$}D$-separable $D$-interior algebra.
(ii) The $(\mathbb{A}^{\prime},\mathbb{A}_{\delta})$-bimodule $e\cdot(V^{*}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})$ induces a Morita equivalence between $\mathbb{A}^{\prime}$ and $\mathbb{A}_{\delta}$. Here, $V^{*}=\mathrm{Hom}(V,\mbox{$\mathcal{O}$})$ is the $\mathcal{O}$-dual module of $V$.
(iii) For any nontrivial subgroup $Q$ of $D$, $\mathbb{A}^{\prime}(Q)\cong kD$ as $D$-interior algebras.
Proof.
Since the rank of $V$ is coprime to $p$, by [16, Proposition 5.1.18] and Corollary 3.4, $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$ is a relatively $\mbox{$\mathcal{O}$}D$-separable $D$-interior algebra. Clearly, the $\mathcal{O}$-rank of $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$ equals $\mathrm{rank}_{\mathcal{O}}(S^{\circ})\cdot\mathrm{rank}_{\mathcal{O}}(\mathbb{A}_{\delta})$. By Corollary 4.7, $\frac{\mathrm{rank}_{\mathcal{O}}(S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})}{|D|}$ is coprime to $p$. By Lemma 5.11 and argument in the paragraph 5.12. applying Lemma 3.3 to $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$, we can get the statements (i) and (ii).
For any nontrivial subgroup $Q$ of $D$, by [22, Proposition 5.6], there is a $D$-interior algebra isomorphism $(S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})(Q)\cong S^{\circ}(Q)\mathop{\otimes}\limits_{k}\mathbb{A}_{\delta}(Q)$, which is isomorphic to $(S(Q))^{\circ}\mathop{\otimes}\limits_{k}\mathbb{A}_{\delta}(Q)$. By Proposition 4.11, $\mathbb{A}_{\delta}(Q)$ is isomorphic to $\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD$ as $D$-interior algebras. Note that $S(Q)\cong\bar{S}_{Q}$ as $D$-interior algebras. Therefore, $(S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})(Q)$ is isomorphic to $\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD$ as $D$-interior algebras. Since $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$ has a finite $D\times D$-stable $\mathcal{O}$-basis by Lemma 5.11, we have $((S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})(Q))(D)\cong(S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})(D)$. This implies that $\mathrm{Br}_{Q}(e)$ belongs to a local point of $D$ on $(S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})(Q)$, which is isomorphic to $\bar{S}_{Q}^{\circ}\mathop{\otimes}\limits_{k}\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD$. Then the statement (iii) follows from [17, Proposition 7.3.10 (iv)]. ∎
Now we can state the main result in this section and give a proof of it.
Theorem 5.14.
Assume that the block $b$ is a hyperfocal abelian Frobenius block. Then there is a stable equivalence of Morita type between $\mathbb{A}_{\delta}$ and $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$. Consequently, there is a stable equivalence of Morita type between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$.
Proof.
By [17, Proposition 6.15.2], the structural map $\mbox{$\mathcal{O}$}L\longrightarrow\mathbb{A}^{\prime}$ splits as $\mbox{$\mathcal{O}$}L$-$\mbox{$\mathcal{O}$}L$-bimodules. We identify $L$ with a subgroup of $(\mathbb{A}^{\prime})^{*}$. Therefore, as $\mbox{$\mathcal{O}$}L$-$\mbox{$\mathcal{O}$}L$-bimodules, $\mathbb{A}^{\prime}=\mbox{$\mathcal{O}$}L\oplus\mathbb{Y}$ for some $\mbox{$\mathcal{O}$}L$-$\mbox{$\mathcal{O}$}L$-subbimodules $\mathbb{Y}$ of $\mathbb{A}^{\prime}$. In particular, $\mathbb{Y}$ is a direct summand of $\mathbb{A}^{\prime}$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. It is easy to check that $C_{L}(Q)=D$ for any nontrivial subgroup $Q$ of $D$. Then $(\mbox{$\mathcal{O}$}L)(Q)\cong kD$ for any nontrivial subgroup $Q$ of $D$. By Proposition 5.13 (iii), we have $\mathbb{Y}(Q)=0$ for any nontrivial subgroup $Q$ of $D$.
On the other hand, as $\mbox{$\mathcal{O}$}(D\times D)$-modules, $\mathbb{A}^{\prime}$ is a direct summand of $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$. Hence, $\mathbb{Y}$ is also a direct summand of $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. By the $\mbox{$\mathcal{O}$}(D\times D)$-isomorphism (5.11.1), $\mathbb{Y}$ has to be isomorphic to a direct summand of $\mathbb{X}$. Therefore, every direct summand of $\mathbb{Y}$ has the form $\mathrm{Ind}_{\Delta_{\tilde{h}}(Q)}^{D\times D}(\mbox{$\mathcal{O}$})$ for some proper subgroup $Q$ of $D$ and some $\tilde{h}\in E_{\mathfrak{h}}$. Suppose that $\mathbb{Y}$ has a direct summand $U$ isomorphic to $\mathrm{Ind}_{\Delta_{\tilde{h}}(Q)}^{D\times D}(\mbox{$\mathcal{O}$})$ with proper subgroup $1\neq Q$ of $D$. Then there is an element $y^{\prime}_{U}$ of $\mathbb{Y}$ such that for any $(w_{1},w_{2})\in D\times D$, $w_{1}y^{\prime}_{U}w_{2}^{-1}=y^{\prime}_{U}$ if and only if $w_{1}\in Q$ and $w_{2}=hw_{1}h^{-1}$ and $U=\mbox{$\mathcal{O}$}Dy^{\prime}_{U}D$. Then we have for any $(w_{1},w_{2})\in D\times D$, $w_{1}y^{\prime}_{U}\tilde{h}w_{2}^{-1}=y^{\prime}_{U}\tilde{h}$ if and only if $w_{1}\in Q$ and $w_{1}=w_{2}$ and then $\mbox{$\mathcal{O}$}Dy^{\prime}_{U}\tilde{h}D$ is isomorphic to $\mathrm{Ind}_{\Delta(Q)}^{D\times D}(\mbox{$\mathcal{O}$})$. Since $\mathbb{Y}$ is an $\mbox{$\mathcal{O}$}L$-$\mbox{$\mathcal{O}$}L$-subbimodule of $\mathbb{A}^{\prime}$, $\mbox{$\mathcal{O}$}Dy^{\prime}_{U}\tilde{h}D$ is contained in $\mathbb{Y}$ and also a direct summand of $\mathbb{Y}$ as $\mbox{$\mathcal{O}$}(D\times D)$-modules. But it is easy to check that $(\mathrm{Ind}_{\Delta(Q)}^{D\times D}(\mbox{$\mathcal{O}$}))(Q)\neq 0$. This is impossible since $\mathbb{Y}(Q)=0$ when $Q\neq 1$. In particular, as an $\mbox{$\mathcal{O}$}(D\times D)$-module, $\mathbb{Y}$ is projective.
Clearly, $\mbox{$\mathcal{O}$}L$ is a relatively $\mbox{$\mathcal{O}$}D$-separable $D$-interior algebra. So as $\mbox{$\mathcal{O}$}L$-$\mbox{$\mathcal{O}$}L$-bimodules, $\mbox{$\mathcal{O}$}L$ is isomorphic to a direct summand of $\mbox{$\mathcal{O}$}L\mathop{\otimes}\limits_{\mathcal{O}D}\mbox{$\mathcal{O}$}L$. Obviously, as $\mbox{$\mathcal{O}$}L$-$\mbox{$\mathcal{O}$}L$-bimodules, $\mathbb{Y}$ is isomorphic to $\mbox{$\mathcal{O}$}L\mathop{\otimes}\limits_{\mathcal{O}L}\mathbb{Y}\mathop{\otimes}\limits_{\mathcal{O}L}\mbox{$\mathcal{O}$}L$. Therefore, as $\mbox{$\mathcal{O}$}L$-$\mbox{$\mathcal{O}$}L$-bimodules, $\mathbb{Y}$ is isomorphic to a direct summand of $\mbox{$\mathcal{O}$}L\mathop{\otimes}\limits_{\mathcal{O}D}\mathbb{Y}\mathop{\otimes}\limits_{\mathcal{O}D}\mbox{$\mathcal{O}$}L$. Since $\mbox{$\mathcal{O}$}L$ is a projective $\mbox{$\mathcal{O}$}D$-$\mbox{$\mathcal{O}$}D$-bimodule, $\mbox{$\mathcal{O}$}L$ is also a projective $\mbox{$\mathcal{O}$}L$-$\mbox{$\mathcal{O}$}L$-bimodule. Denote the $\mathbb{A}^{\prime}$-$\mbox{$\mathcal{O}$}L$ bimodule $\mathbb{A}^{\prime}$ by $\mathbb{A}^{\prime}_{\mathcal{O}L}$ and the $\mbox{$\mathcal{O}$}L$-$\mathbb{A}^{\prime}$-bimodule $\mathbb{A}^{\prime}$ by ${{}_{\mathcal{O}L}}\mathbb{A}^{\prime}$. Then by [16, Proposition 4.14.12], the bimodules $\mathbb{A}^{\prime}_{\mathcal{O}L}$ and ${{}_{\mathcal{O}L}}\mathbb{A}^{\prime}$ induce a stable equivalence of Mortia type between $\mathbb{A}^{\prime}$ and $\mbox{$\mathcal{O}$}L$. By Proposition 5.13 (ii), the $\mathbb{A}_{\delta}$-$\mbox{$\mathcal{O}$}L$-bimodule $(V\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})\cdot e$ and the $\mbox{$\mathcal{O}$}L$-$\mathbb{A}_{\delta}$-bimodule $e\cdot(V^{*}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})$ induce a stable equivalence of Mortia type between $\mathbb{A}_{\delta}$ and $\mbox{$\mathcal{O}$}L$. Moreover, by Corollary 3.4, the $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule $\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}((V\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})\cdot e)$ and the $\mbox{$\mathcal{O}$}L$-$\mathbb{A}$-bimodule $e\cdot(V^{*}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}j\mathbb{A}$ induce a stable equivalence of Mortia type between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$. ∎
Corollary 5.15.
The Cartan matrix of the hyperfocal subalgebra $\mathbb{A}$ is nonsingular, the absolute value of which determinant equals $|D|$ when the block $b$ is a hyperfocal abelian Frobenius block.
Proof.
This can be deduced easily from [16, Proposition 4.14.3] and the fact that the Cartan matrix of $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{hyp}})$ has determinant $|D|$. ∎
5.16.
Keep the notation as above. Denote the $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule $\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}((V\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})\cdot e)$ by $M_{\mathrm{st}}$. It is clear that $M_{\mathrm{st}}$ is an indecomposable $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule. In order to get the information about source algebras, we need to consider the $\Delta(P)$-equivariant property of the bimodule $M_{\mathrm{st}}$. Here, $\Delta(P)=\{(u,u)\mid u\in P\}$ is the diagonal subgroup of $P\times P$. The following is an investigation of this issue.
5.17.
Borrowing the notation from Proposition 4.12, there is a unitary subalgebra of $\mathbb{A}_{\delta}$ isomorphic to $\mbox{$\mathcal{O}$}L$ as $D$-interior algebras, which is still denoted by $\mbox{$\mathcal{O}$}L$ for simplicity. At the same time, there is an element $a_{u,\delta}$ in $N_{A_{\delta}}(D)$ such that $a_{u,\delta}(vj)a_{u,\delta}^{-1}=(uvu^{-1})j$ and $\mathbb{A}_{\delta}^{a_{u,\delta}}=\mathbb{A}_{\delta}$ for any $u\in P$ and any $v\in D$. In particular, the subgroup $j+J(\mathbb{A}_{\delta})^{D}$ is stable under the conjugation action of $a_{u,\delta}$ for any $u\in P$. Recall from the proof of Proposition 4.12 that we have the following split short exact sequence
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${k^{*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N_{\mathbb{A}_{\delta}}(D)/(j+J(\mathbb{A}_{\delta})^{D})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{\mathfrak{h}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
Then we have a complement $\hat{E}_{\mathfrak{h}}/(j+J(\mathbb{A}_{\delta})^{D})$ of $k^{*}$ in $N_{\mathbb{A}_{\delta}}(D)/(j+J(\mathbb{A}_{\delta})^{D})$. Denote by $\pi$ the sujective homomorphism of the above sequence, which induces the following another split short exact sequence
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${j+J(\mathbb{A}_{\delta})^{D}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\hat{E}_{\mathfrak{h}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{\mathfrak{h}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
For simplicity, we still denote by $E_{\mathfrak{h}}$ and $\pi$ a complement of $j+J(\mathbb{A}_{\delta})^{D}$ in $\hat{E}_{\mathfrak{h}}$ and the surjective homomorphism of this sequence, respectively.
Lemma 5.18.
Let $u$ be an element of $P$ such that it acts trivially on $E_{\mathfrak{h}}$ through the canonical map $P\longrightarrow P/C_{P}(D)$. Then the element $a_{u,\delta}$ can be chosen such that it acts trivially on $E_{\mathfrak{h}}$ through the conjugation action.
Proof.
Denote $\hat{E}_{\mathfrak{h}}/(j+J(\mathbb{A}_{\delta})^{D})$ by $\bar{\hat{E}}_{\mathfrak{h}}$. Then for any $\mathfrak{a}\in\hat{E}_{\mathfrak{h}}$, denoting by $\bar{\mathfrak{a}}$ its image in $\bar{\hat{E}}_{\mathfrak{h}}$, we have $\pi(a_{u,\delta}\bar{\mathfrak{a}}a_{u,\delta}^{-1})={{}^{u}}\pi(\bar{\mathfrak{a}})$, which is just $\pi(\bar{\mathfrak{a}})$ by the assumption of $u$. Therefore, we have $a_{u,\delta}\bar{\mathfrak{a}}a_{u,\delta}^{-1}=\lambda_{\mathfrak{a}}\bar{\mathfrak{a}}$ for some $\lambda_{\mathfrak{a}}\in k^{*}$. Furthermore, it is clear that some $p$-th power of $a_{u,\delta}$ has to be in $(\mathbb{A}_{\delta}^{D})^{*}$, which is isomorphic to $k^{*}\times(j+J(\mathbb{A}_{\delta})^{D})$. Then this implies some $p$-th power of $\lambda_{\mathfrak{a}}$ being $1$ and hence $\lambda_{\mathfrak{a}}$ itself is trivial. Therefore, we have $a_{u,\delta}\hat{E}_{\mathfrak{h}}a_{u,\delta}^{-1}=\hat{E}_{\mathfrak{h}}$. In particular, $a_{u,\delta}E_{\mathfrak{h}}a_{u,\delta}^{-1}$ is also a complement of $j+J(\mathbb{A}_{\delta})^{D}$ in $\hat{E}_{\mathfrak{h}}$.
Since $E_{\mathfrak{h}}$ is a $p^{\prime}$-group, by [29, Lemma (45.6)], $a_{u,\delta}E_{\mathfrak{h}}a_{u,\delta}^{-1}$ and $E_{\mathfrak{h}}$ are conjugate to each other by an element of $j+J(\mathbb{A}_{\delta})^{D}$. Up to some adjustment, we can assume that $a_{u,\delta}E_{\mathfrak{h}}a_{u,\delta}^{-1}=E_{\mathfrak{h}}$. For any $\mathfrak{a}\in E_{\mathfrak{h}}$, it is clear that $\pi(a_{u,\delta}\mathfrak{a}a_{u,\delta}^{-1})=\pi(\mathfrak{a})$. Then we have $a_{u,\delta}\mathfrak{a}a_{u,\delta}^{-1}=\mathfrak{a}$. Then the lemma holds. ∎
5.19.
Fix an element $u$ of $P$ satisfying that it acts trivially on $E_{\mathfrak{h}}$ through the canonical map $P\longrightarrow P/C_{P}(D)$. By Lemma 5.18, we can assume that $a_{u,\delta}$ acts trivially on the subgroup $E_{\mathfrak{h}}$ of $N_{\mathbb{A}_{\delta}}(D)$ through the conjugation action. Then we have the twisted $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule ${{}_{(u,u)}}M_{\mathrm{st}}$, which is equal to $M_{\mathrm{st}}$ as $\mathcal{O}$-modules, with $(uau^{-1},uxu^{-1})$ acting on $m\in{{}_{(u,u)}}M_{\mathrm{st}}$ as $(a,x)$ on $m\in M_{\mathrm{st}}$.
Proposition 5.20.
Keep the notation and assumption as above. Then the two modules $M_{\mathrm{st}}$ and ${{}_{(u,u)}}M_{\mathrm{st}}$ are isomorphic to each other as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules.
Proof.
Recall that from the paragraph 5.9 that we have an element $s_{u}$ of $S=\mathrm{End}_{\mathcal{O}}(V)$ such that $s_{u}(x\cdot\mathrm{v})=(uxu^{-1})\cdot s_{u}(\mathrm{v})$ for any $x\in L$ and any $\mathrm{v}\in V$. At the same time, we have a unitary subalgebra $\mathop{\bigoplus}\limits_{\tilde{g}\in E_{\mathfrak{h}}}\mbox{$\mathcal{O}$}D\mathfrak{e}_{\tilde{g}}$ of $\mathbb{A}_{\delta}$ isomorphic to $\mbox{$\mathcal{O}$}L$ (see the proof of Proposition 4.12). By Lemma 5.18, we can assume that $a_{u,\delta}\mathfrak{e}_{\tilde{g}}a_{u,\delta}^{-1}=\mathfrak{e}_{\tilde{g}}$ for any $\tilde{g}\in E_{\mathfrak{h}}$. Note that $M_{\mathrm{st}}=\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}((V\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})\cdot e)$. We can define the following map from $M_{\mathrm{st}}$ to ${{}_{(u,u)}}M_{\mathrm{st}}$
$\tau:M_{\mathrm{st}}\longrightarrow{{}_{(u,u)}}M_{\mathrm{st}},~~~~\mathfrak{a}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}((\mathrm{v}\mathop{\otimes}\limits_{\mathcal{O}}\mathfrak{a}_{\delta})\cdot e)\mapsto u\mathfrak{a}a_{u,\delta}^{-1}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}((s_{u}(\mathrm{v})\mathop{\otimes}\limits_{\mathcal{O}}a_{u,\delta}\mathfrak{a}_{\delta}a_{u,\delta}^{-1})\cdot e),$
where $\mathfrak{a}$ is in $\mathbb{A}$ and $\mathrm{v}$ is in $V$ and $\mathfrak{a}_{\delta}$ is in $\mathbb{A}_{\delta}$. It is easy to check that $\tau$ is well-defined and a bijection. Let $\mathfrak{a}^{\prime}$ and $w$ and $\mathfrak{e}_{\tilde{g}}$ be in $\mathbb{A}$ and $D$ and $E_{\mathfrak{h}}$, respectively.
$)\displaystyle\tau(\mathfrak{a}^{\prime}\cdot(\mathfrak{a}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}((\mathrm{v}\mathop{\otimes}\limits_{\mathcal{O}}\mathfrak{a}_{\delta})\cdot e))\cdot(w\tilde{g}$))
$=$~ $\tau(\mathfrak{a}^{\prime}\mathfrak{a}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}(((w\tilde{g})^{-1}(\mathrm{v})\mathop{\otimes}\limits_{\mathcal{O}}\mathfrak{a}_{\delta}\cdot w\mathfrak{e}_{\tilde{g}})\cdot e))$
$=$~ $u\mathfrak{a}^{\prime}\mathfrak{a}a_{u,\delta}^{-1}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}((s_{u}((w\tilde{g})^{-1}(\mathrm{v}))\mathop{\otimes}\limits_{\mathcal{O}}a_{u,\delta}\mathfrak{a}_{\delta}\cdot w\mathfrak{e}_{\tilde{g}}a_{u,\delta}^{-1})\cdot e)$
$=$~ $(u\mathfrak{a}^{\prime}u^{-1})\cdot u\mathfrak{a}a_{u,\delta}^{-1}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}((uw\tilde{g}u^{-1})^{-1}\cdot s_{u}(\mathrm{v})\mathop{\otimes}\limits_{\mathcal{O}}a_{u,\delta}\mathfrak{a}_{\delta}a_{u,\delta}^{-1}\cdot(a_{u,\delta}w\mathfrak{e}_{\tilde{g}}a_{u,\delta}^{-1})\cdot e)$
[…]
where the multiplications in the last equation is taken in ${{}_{(u,u)}}M_{\mathrm{st}}$. We complete the proof of this proposition. ∎
Corollary 5.21.
(i) Let $u$ be an element of $C_{P}(D)$. Then the two modules $M_{\mathrm{st}}$ and ${{}_{(u,u)}}M_{\mathrm{st}}$ are isomorphic to each other as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules.
(ii) Suppose that $N_{G}(P_{\gamma})$ controls fusion of the block $b$. Then the two modules $M_{\mathrm{st}}$ and ${{}_{(u,u)}}M_{\mathrm{st}}$ are isomorphic to each other as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules for any $u\in P$.
Proof.
Statement (i) follows from Proposition 5.18 directly. Now we assume that $N_{G}(P_{\gamma})$ controls fuison of the block $b$. Let $(P,\bar{b}_{P})$ and $(D,\bar{b}_{D})$ be the $b$-Brauer pairs associated with $P_{\gamma}$ and $D_{\delta}$, respectively. By [25, Proposition 3.3], the block $\bar{b}_{D}$ is nilpotent and so $N_{G}(D,\bar{b}_{D})=N_{G}(D_{\delta})$. Since $N_{G}(P_{\gamma})$ controls fuison of the block $b$, we have
$N_{G}(D_{\delta})=N_{G}(D,\bar{b}_{D})=N_{G}(P,\bar{b}_{D})C_{G}(D)=N_{G}(P_{\gamma})C_{G}(D)$
due to the well-known fact that $N_{G}(P,\bar{b}_{D})=N_{G}(P_{\gamma})$. As a block of $PC_{G}(D)$, it is clear that $b_{D}$ is still nilpotent with a defect group $P$. Therefore, $(P,\bar{b}_{P})$ is still a maximal $b_{D}$-Brauer pair of the block $b_{D}$ of $PC_{G}(D)$. Then
$N_{G}(P_{\gamma})\cap C_{G}(D)=$ $~N_{G}(P,\bar{b}_{P})\cap C_{G}(D)\cap PC_{G}(D)$
$=$ $~N_{PC_{G}(D)}(P,\bar{b}_{D})\cap C_{G}(D)$
$=$ $~PC_{G}(P)\cap C_{G}(D)$
$=$ $~C_{P}(D)C_{G}(P).$
Here, the third equation is due to the definition of nilpotent blocks. Hence, $N_{G}(D_{\delta})/C_{G}(D)$ is isomorphic to $N_{G}(P_{\gamma})/C_{P}(D)C_{G}(P)$ and this isomorphism is just induced by restriction. Then we can identify $E_{\mathfrak{h}}$ with the inertial quotient $E=N_{G}(P_{\gamma})/PC_{G}(P)$ of the block $b$.
On the other hand, by [11, Lemma 4.28], we have $P=[P,E]\cdot C_{P}(E)$, where $[P,E]$ is the commutator subgroup of $P$ and $E$. By the definition of the hyperfocal subgroup, it is clear that $[P,E]$ is contained in $D$. By the assumption that the block $b$ is a hyperfocal abelian Frobenius block, $D\cap C_{P}(E)=C_{D}(E)$ is trivial. Therefore, we have $D=[P,E]$ and $P=D\rtimes C_{P}(E)$. Thus every element $u$ of $P$ satisfies that it acts trivially on $E_{\mathfrak{h}}$ through the canonical map $P\longrightarrow P/C_{P}(D)$. Now statement (ii) follows from Proposition 5.18. ∎
6 Applications
6.1 Klein four group case
Keep the notation from the last three sections. In this subsection, we will assume that the prime number $p$ is equal to $2$ and the hyperfocal subgroup $D$ is a Klein four group. In this case the hyperfocal quotient inertial $E_{\mathfrak{h}}$ is a cyclic group of order $3$ and acts freely on $D-\{1\}$. Then we can apply Theorem 5.14 to this case and get a stable equivalence of Morita type between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$. Here $L=D\rtimes E_{\mathfrak{h}}$ which is just $\mathrm{A}_{4}$ in this section. Here, $\mathrm{A}_{n}$ denotes the alternating group on $n$ letters for any positive integer $n$. In fact, we can classify the Morita equivalent classes and the derived equivalent classes of the hyperfocal subalgebras in this case, which is similar to the case where the defect group is a Klein four group.
First, we need to calculate the numbers of simple $\mbox{$\mathcal{K}$}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}$-modules and simple $k\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}$-modules.
6.1.
Denote $\mbox{$\mathcal{K}$}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}$ and $k\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}$ by $\hat{\mathbb{A}}$ and $\bar{\mathbb{A}}$, respectively. We have analogue notation for the source algebra $A$. It is clear that $\hat{\mathbb{A}}$ is a semi-simple $\mathcal{K}$-algebra (see [13]). By [3, (29.20)], there is a finite extension $\mbox{$\mathcal{K}$}^{\prime}$ of $\mathcal{K}$ such that the semi-simple $\mbox{$\mathcal{K}$}^{\prime}$-algebra $\mbox{$\mathcal{K}$}^{\prime}\mathop{\otimes}\limits_{\mathcal{K}}\hat{\mathbb{A}}$ is split. Since $k$ is algebraically closed, the field $\mbox{$\mathcal{K}$}^{\prime}$ determines a complete $p$-modular system $(\mbox{$\mathcal{K}$}^{\prime},\mbox{$\mathcal{O}$}^{\prime},k)$ such that $\mbox{$\mathcal{O}$}^{\prime}$ is an extension of $\mathcal{O}$ by [28, Chapter 2,§2, Proposition 3]. Replacing $\mathcal{O}$ by $\mbox{$\mathcal{O}$}^{\prime}$, we can assume that the $\mathcal{K}$-algebra $\hat{\mathbb{A}}$ is a split semi-simple $\mathcal{K}$-algebra. Then we can borrow the notation from the paragraphs 2.7-2.8 for these three $\mathcal{O}$-algebras $A$ and $\mathbb{A}$ and $\mathbb{A}_{\delta}$.
Lemma 6.2.
We have $\mathrm{k}(\mathbb{A})=4$ and $l(\mathbb{A})=3$.
Proof.
By Theorem 5.14, there is a stable equivalence of Morita type between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$. It is clear that the Cartan matrix of $\mbox{$\mathcal{O}$}L$ is nonsingular. Then by [16, Proposition 4.14.13], the Cartan matrix of $\bar{\mathbb{A}}$ is also nonsingular. Note that the hyperfocal subalgebra $\mathbb{A}$ is a symmetric $\mathcal{O}$-algebra (see [10, Theorem 1.1]). Then by [12, Proposition 3.1] the stable equivalence between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$ induces an isometry between $L^{0}(\mathbb{A})$ and $L^{0}(\mbox{$\mathcal{O}$}L)$. It is clear that $L^{0}(\mbox{$\mathcal{O}$}L)$ is a free $\mathbb{Z}$-module of rank one and setting $\zeta$ to be a generator, $\langle\zeta,\zeta\rangle_{\mathcal{O}L}=4$. Then $L^{0}(\mathbb{A})$ is also a free $\mathbb{Z}$-module of rank one which implies $\mathrm{k}(\mathbb{A})-l(\mathbb{A})=1$. Suppose that $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})=\{U_{1},U_{2},\cdots,U_{t}\}$ and $\sum_{a=1}^{t}n_{a}[U_{a}]\in L^{0}(\mathbb{A})$ is a generator, where $n_{1},n_{2},\cdots,n_{t}$ are integers. Then we have $\sum\limits_{a=1}^{t}n_{a}^{2}=4$. The only possibility is that $t=4$ and $n_{a}=\pm 1$ for $a=1,2,3,4$. We are done. ∎
6.3.
Let $\{T_{1},T_{2},T_{3}\}$ be a set of representatives of the isomorphism classes of simple $kL$-modules. Denote by $\Omega=\Omega_{kL}$ the Heller operator of $kL$. By [17, Corollary 7.2.11], there are six isomorphism classes of indecomposable $kL$-modules of dimension $2$, denoted by $T_{\mathrm{i}}^{\mathrm{j}}$ with distinct $\mathrm{i}$ and $\mathrm{j}$ in $\{1,2,3\}$. Furthermore, the indecomposable $kL$-module $T_{\mathrm{i}}^{\mathrm{j}}$ has composition series $T_{\mathrm{j}}$, $T_{\mathrm{i}}$ and $\Omega(T_{\mathrm{i}}^{\mathrm{j}})=T_{\mathrm{j}}^{\mathrm{l}}$, where $\mathrm{l}\in\{1,2,3\}-\{\mathrm{i,\mathrm{j}}\}$. Since the two algebras $\mathbb{A}$ and $\mathbb{A}_{\delta}$ are Morita equivalent, by Lemma 6.2, we can set $\{\bar{W}_{1},\bar{W}_{2},\bar{W}_{3}\}$ to be a set of representatives of the isomorphism classes of simple $\bar{\mathbb{A}}_{\delta}$-modules. Now we can state the following main theorem in this section and give a proof of it, which is similar to the one of [17, Theorem 12.1.2].
Theorem 6.4.
Keep the notation as above. Assume that the hyperfocal subgroup $D$ is a Klein four group. Then the hyperfocal subalgebra $\mathbb{A}$ is Morita equivalent to either $\mbox{$\mathcal{O}$}\mathrm{A}_{4}$ or the principal block algebra of $\mbox{$\mathcal{O}$}\mathrm{A}_{5}$. In particular, the hyperfocal subalgebra $\mathbb{A}$ is Rickard equivalent to $\mbox{$\mathcal{O}$}(D\rtimes C_{3})$, where $C_{3}$ denotes a cyclic group of order $3$.
Proof.
For any nontrivial subgroup $Q$ of $D$, it is clear that the quotient group $D/Q$ is either trivial or a cyclic group of order $2$. In each case, the primitive Dade $D$-interior algebra $S(Q)$ is just $k$. Then the primitive Dade $D$-interior algebra $S$ has to be $k$ and the primitive $L$-interior algebra $\mathbb{A}^{\prime}$ in the paragraph 5.12 is just $\mathbb{A}_{\delta}$. By applying Theorem 5.14, we get the stable equivalence of Morita type between $\mathbb{A}_{\delta}$ and $\mbox{$\mathcal{O}$}L$ is just given by the restriction. By Corollary 4.7, there is at least one of the simple $\bar{\mathbb{A}}_{\delta}$-modules has odd dimension. Now we can apply the same arguments of the proof of [17, Proposition 12.2.9] to our setting and obtain that exactly one of the following statments holds.
(i) There is an integer $n$ such that for $1\leq\mathrm{i}\leq 3$
$\mathrm{Res}_{kL}^{\bar{\mathbb{A}}_{\delta}}(\bar{W}_{\mathrm{i}})\cong\Omega^{n}(T_{\mathrm{i}}).$
(ii) There is an integer $n$ such that
$\mathrm{Res}_{kL}^{\bar{\mathbb{A}}_{\delta}}(\bar{W}_{1})\cong\Omega^{n}(T_{1}),$
$\mathrm{Res}_{kL}^{\bar{\mathbb{A}}_{\delta}}(\bar{W}_{2})\cong\Omega^{n}(T_{3}^{2}),$
$\mathrm{Res}_{kL}^{\bar{\mathbb{A}}_{\delta}}(\bar{W}_{3})\cong\Omega^{n}(T_{2}^{3}).$
In the case (i), we have $\bar{\mathbb{A}}_{\delta}\mathop{\otimes}\limits_{kL}T_{\mathrm{i}}\cong\Omega^{-n}_{\bar{\mathbb{A}}_{\delta}}(\bar{W}_{\mathrm{i}})$ for any $1\leq\mathrm{i}\leq 3$. Therefore by [16, Theorem 4.14.10], the $\mathbb{A}_{\delta}$-$\mbox{$\mathcal{O}$}L$-bimodule $M=\Omega_{\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathcal{O}}(\mathcal{O}L)^{\circ}}^{n}(\mathbb{A}_{\delta})$ and its dual induce a Morita equivalence between $\mathbb{A}_{\delta}$ and $\mbox{$\mathcal{O}$}L$. Hence, the $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule $\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}M$ and its dual $M^{*}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}j\mathbb{A}$ induce a Morita equivalence between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$, here $j\in\delta$ (see the paragraph 3.1).
In the case (ii), we denote by $B_{0}(\mathrm{A}_{5})$ the principal block algebra of $\mbox{$\mathcal{O}$}\mathrm{A}_{5}$. The same argument in the last paragraph of the proof of [17, Theorem 12.1.2] shows that there is an indecomposable summand $N^{\prime}$ of the $\mathbb{A}_{\delta}$-$B_{0}(\mathrm{A}_{5})$-bimodule $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathcal{O}L}B_{0}(\mathrm{A}_{5})$ such that it and its dual induce a stable equivalence of Morita type between $\mathbb{A}_{\delta}$ and $B_{0}(\mathrm{A}_{5})$ and the $\mathbb{A}_{\delta}$-$B_{0}(\mathrm{A}_{5})$-bimodule $N=\Omega_{\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathcal{O}}B_{0}(\mathrm{A}_{5})^{\circ}}^{n}(N^{\prime})$ and its dual induce a Morita equivalence between $\mathbb{A}_{\delta}$ and $B_{0}(\mathrm{A}_{5})$. Hence, the $\mathbb{A}$-$B_{0}(\mathrm{A}_{5})$-bimodule $\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}N$ and its dual $N^{*}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}j\mathbb{A}$ induce a Morita equivalence between $\mathbb{A}$ and $B_{0}(\mathrm{A}_{5})$.
Finally, by [17, Theorem 12.4.1], $\mbox{$\mathcal{O}$}L$ and $B_{0}(\mathrm{A}_{5})$ are Rickard equivalent and then the hyperfocal subalgebra $\mathbb{A}$ is Rickard equivalent to $\mbox{$\mathcal{O}$}L$, which is just $\mbox{$\mathcal{O}$}(D\rtimes C_{3})$. ∎
6.5.
[…]
Proposition 6.6.
[…]
Proof.
By the structure theory of source algebras with normal defect group, the source algebra of the block $b_{0}$ is isomorphic to $\mbox{$\mathcal{O}$}(L\times R)$. By [7, Theorem 1.1 (i)], we have $l(b):=|\mathrm{IBr}(b)|=3=|E_{\mathfrak{h}}|$. Then by [31, Theorem], there is an isotypy between the blocks $b$ and $c$. Recall that $A$ and $\mathbb{A}$ are a source algebra and a hyperfocal subalgebra of the block $b$, respectively. Then by [9, Proposition 4.8], there is an $\mathcal{O}$-algebra isomorphism $A\cong\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}\mbox{$\mathcal{O}$}R$. It is clear that as $\mathcal{O}$-algebras, $\mbox{$\mathcal{O}$}(\mathrm{A}_{4}\times R)$ is isomorphic to $\mbox{$\mathcal{O}$}\mathrm{A}_{4}\mathop{\otimes}\limits_{\mathcal{O}}\mbox{$\mathcal{O}$}R$ and the principal block algebra of $\mbox{$\mathcal{O}$}(\mathrm{A}_{5}\times R)$ is isomorphic to $B_{0}(\mathrm{A}_{5})\mathop{\otimes}\limits_{\mathcal{O}}\mbox{$\mathcal{O}$}R$, where $B_{0}(\mathrm{A}_{5})$ denotes the principal block algebra of $\mbox{$\mathcal{O}$}\mathrm{A}_{5}$. Then the statements of the proposition follows from Theorem 6.4. ∎
On the other hand, we can use Theorem 6.4 to obtain some information about $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})$ and $\mathrm{Irr}_{\mathcal{K}}(A)$. We will first verify the forward direction of the KLN conjecture for the hyperfocal subgroup $D$ being Klein four group. We need the following two lemmas.
Lemma 6.7.
The dimension of every simple $\hat{\mathbb{A}}_{\delta}$-module is coprime to $2$.
Proof.
We set $\{\hat{U}_{1},\hat{U}_{2},\hat{U}_{3},\hat{U}_{4}\}$ to be a set of representatives of the isomorphism classes of simple $\hat{\mathbb{A}}_{\delta}$-modules.
Suppose that $\mathbb{A}_{\delta}$ is Morita equivalent to $\mbox{$\mathcal{O}$}L$. Then we are in the case (i) of Theorem 6.4. In particular, the dimension of every simple $\bar{\mathbb{A}}_{\delta}$-module $\bar{W}_{\mathrm{i}}$ is coprime to $2$ by [17, Theorem 7.2.10]. Since $\mathbb{A}_{\delta}$ is Morita equivalent to $\mbox{$\mathcal{O}$}L$, we can assume that $d_{\mathbb{A}_{\delta}}([\hat{U}_{4}])=\sum\limits_{\mathrm{i}=1}^{3}[\bar{W}_{\mathrm{i}}]$ and $d_{\mathbb{A}_{\delta}}([\hat{U}_{\mathrm{i}}])=[\bar{W}_{\mathrm{i}}]$ for any $\mathrm{i}\in\{1,2,3\}$. By the definition of the decomposition map $d_{\mathbb{A}_{\delta}}$, we have $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{4})=\sum\limits_{\mathrm{i}=1}^{3}\mathrm{dim}_{k}\bar{W}_{\mathrm{i}}$ and $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\mathrm{i}})=\mathrm{dim}_{k}\bar{W}_{\mathrm{i}}$ for for any $\mathrm{i}\in\{1,2,3\}$. In particular, every simple $\hat{A}_{\delta}$-module has dimension coprime to $2$ in this case.
Suppose that $\mathbb{A}_{\delta}$ is Morita equivalent to the principal block $B_{0}(\mathrm{A}_{5})$ of $\mathrm{A}_{5}$. Then we are in the case (ii) of Theorem 6.4. In particular, $\mathrm{dim}_{k}(\bar{W}_{1})$ is coprime to $2$ and $\mathrm{dim}_{k}(\bar{W}_{2})=\mathrm{dim}_{k}(\bar{W}_{3})=2$ by [17, Corollary 7.2.11]. Since $\mathbb{A}_{\delta}$ is Morita equivalent to the principal block $B_{0}(\mathrm{A}_{5})$ of $\mathrm{A}_{5}$, of which the decomposition matrix is well-known, we can assume that $d_{\mathbb{A}_{\delta}}([\hat{U}_{1}])=[\bar{W}_{1}]$ and $d_{\mathbb{A}_{\delta}}([\hat{U}_{4}])=[\bar{W}_{1}]+[\bar{W}_{2}]+[\bar{W}_{3}]$ and $d_{\mathbb{A}_{\delta}}([\hat{U}_{\mathrm{i}}])=[\bar{W}_{1}]+[\bar{W}_{\mathrm{i}}]$ for any $\mathrm{i}\in\{2,3\}$. Hence, we still obtain that every simple $\hat{A}_{\delta}$-module has dimension coprime to $2$ in this case. We are done. ∎
Lemma 6.8.
Let $U$ be a finitely generated $\mathbb{A}_{\delta}$-module which is free as an $\mathcal{O}$-module. Then we have $\mathrm{rank}_{\mathcal{O}}(U)\equiv\mathrm{rank}_{\mathcal{O}}(\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U)~(\mathrm{mod}~2)$. In particular, $\mathrm{rank}_{\mathcal{O}}(U)$ is coprime to $2$ if and only if $\mathrm{rank}_{\mathcal{O}}(\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U)$ is coprime to $2$.
Proof.
First, note that $\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U$ is $\mathcal{O}$-free when $U$ is $\mathcal{O}$-free. So $\mathrm{rank}_{\mathcal{O}}(\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U)$ makes sense. Denote by $m_{\delta}$ the multiplicity of $\mathbb{A}$ at $\delta$ (see [19, Definition 2.2]). Then by [25, Propositions 3.3 and 3.5], we have $m_{\delta}$ coprime to $2$. Let $j^{\prime}$ be another element of $\delta$. Obviously, we have $j^{\prime}\mathbb{A}j$ isomorphic to $\mathbb{A}_{\delta}$ as $\mbox{$\mathcal{O}$}D$-$\mathbb{A}_{\delta}$-bimodules. Since $D_{\delta}$ is the unique local pointed group of $D$ on $\mathbb{A}$, there is an $\mbox{$\mathcal{O}$}D$-module decomposition
$\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U\cong U^{m_{\delta}}\oplus(e\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U)$ (6.8.1)
for some idempotent $e$ of $\mathbb{A}^{D}$ with $\mathrm{Br}_{D}(e)=0$. Here, $U^{m_{\delta}}$ denotes a direct sum of $m_{\delta}$-copies of $U$. Suppose that $e$ is primitive. By Rosenberg’s lemma, we have $e\in\mathrm{Tr}_{Q}^{D}(\mathbb{A}^{Q})$ for some proper subgroup $Q$ of $D$. By [29, Theorem 23.1], there is a primitive idempotent $e^{\prime}$ of $\mathbb{A}^{Q}$ such that $e=\mathrm{Tr}_{Q}^{D}(e^{\prime})$ and $e^{\prime}\cdot(e^{\prime})^{v}=(e^{\prime})^{v}\cdot e^{\prime}=0$ for any $v\in D-Q$. For any $\mathfrak{a}\in e\mathbb{A}j$ and any $x\in U$, the map sending $\mathfrak{a}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}x$ to $\sum\limits_{uD\in D/Q}u\mathop{\otimes}\limits_{\mathcal{O}Q}(e^{\prime}u^{-1}\mathfrak{a}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}x)$ determines an isomorphism of $\mbox{$\mathcal{O}$}D$-modules $e\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}W_{\delta}\cong\mbox{$\mathcal{O}$}D\mathop{\otimes}\limits_{\mathcal{O}Q}(e^{\prime}\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U)$. In particular, $\mathrm{rank}_{\mathcal{O}}(e\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U)$ is divided by $2$. In general, by decomposing $e$ as a sum of pairwise orthogonal primitive idempotents of $\mathbb{A}^{D}$, we still have $\mathrm{rank}_{\mathcal{O}}(e\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}W_{\delta})$ divided by $2$. Hence, by the decomposition (6.8.1), we have $m_{\delta}\mathrm{rank}_{\mathcal{O}}(U)\equiv\mathrm{rank}_{\mathcal{O}}(\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}U)~(\mathrm{mod}~2)$. Then the lemma follows from $m_{\delta}$ coprime to $2$. ∎
Combining Lemmas 6.7 and 6.8 with the Morita equivalence between $\mathbb{A}$ and $\mathbb{A}_{\delta}$, we can get the following result about the dimensions of simple $\hat{\mathbb{A}}-$modules.
Proposition 6.9.
The dimension of every simple $\hat{\mathbb{A}}-$module is coprime to $2$. In particular, the forward direction of the KLN conjecture is true in this case.
6.10.
Set $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})=\{\hat{U}_{1},\hat{U}_{2},\hat{U}_{3},\hat{U}_{4}\}$ and $\mathrm{Irr}_{k}(\mathbb{A})=\{\bar{W}_{1},\bar{W}_{2},\bar{W}_{3}\}$ if no confusion arises. By the proof of Lemma 6.7, we can describe the decomposition map $d_{\mathbb{A}}$ of $\mathbb{A}$ as follows. When $\mathbb{A}$ is Morita equivalent to $\mbox{$\mathcal{O}$}L$, then for any $\mathrm{i}\in\{1,2,3\}$,
$d_{\mathbb{A}}([\hat{U}_{\mathrm{i}}])=[\bar{W}_{\mathrm{i}}]~\mathrm{and}~d_{\mathbb{A}}([\hat{U}_{4}])=\sum_{\mathrm{i}=1}^{3}[\bar{W}_{\mathrm{i}}]$ (6.10.1)
When $\mathbb{A}$ is Morita equivalent to the principal block $B_{0}(\mathrm{A}_{5})$ of $\mathrm{A}_{5}$, then
$d_{\mathbb{A}}([\hat{U}_{1}])=[\bar{W}_{1}]~\mathrm{and}~d_{\mathbb{A}}([\hat{U}_{4}])=\sum_{\mathrm{i}=1}^{3}[\bar{W}_{\mathrm{i}}]~\mathrm{and}~d_{\mathbb{A}}([\hat{U}_{\mathrm{i}}])=[\bar{W}_{1}]+[\bar{W}_{\mathrm{i}}]$ (6.10.2)
for any $\mathrm{i}\in\{2,3\}$. It is clear that the conjugation action of $P$ on $\mathbb{A}$ can induce actions of $P$ on $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})$ and $\mathrm{Irr}_{k}(\mathbb{A})$, respectively. First, we demonstrate that $C_{P}(D)$ acts trivially on these two sets. This follows from the equivariant property of the Morita equivalences occurring in Theorem 6.4. For simplicity, we denote $C_{P}(D)$ by $P_{0}$.
Lemma 6.11.
Suppose that $P_{0}$ acts trivially on $L$ and $\mathrm{A}_{5}$. Then the Morita equivalences occurring in Theorem 6.4 are both $\Delta(P_{0})$-equivariant.
Proof.
We will borrow the notation from the proof of Theorem 6.4. Fix an element $u$ of $P_{0}$.
First, we assume that the hyperfocal subalgebra $\mathbb{A}$ is Morita equivalent to $\mbox{$\mathcal{O}$}L$. Then the $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule $\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}M$ and its dual induce a Morita equivalence between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$, where $M=\Omega_{\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathcal{O}}(\mathcal{O}L)^{\circ}}^{n}(\mathbb{A}_{\delta})$. By Lemma 5.18, there is an element $a_{u,\delta}$ of $A_{\delta}^{*}$ such that it acts trivially on the subalgebra $\mbox{$\mathcal{O}$}L$ of $\mathbb{A}_{\delta}$ through the conjugation action on $\mathbb{A}_{\delta}$. For any $\mathbb{A}_{\delta}$-$\mbox{$\mathcal{O}$}L$-bimodule $U$, we can define the twisted bimodule ${{}_{(u,u)}}U$ by setting ${{}_{(u,u)}}U=U$ as an $\mathcal{O}$-module, with $a_{u,\delta}\mathfrak{a}a_{u,\delta}^{-1}\otimes_{\mathcal{O}}y$ acting on $\mathrm{u}\in{{}_{(u,u)}}U$ as $\mathfrak{a}\otimes_{\mathcal{O}}y$ on $\mathrm{u}\in U$, where $\mathfrak{a}$ belongs to $\mathbb{A}_{\delta}$ and $y$ belongs to $L$. Then it is easy to check that its conjugation action on $\mathbb{A}_{\delta}$ can induce an $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule isomorphism between $\mathbb{A}_{\delta}$ and ${{}_{(u,u)}}\mathbb{A}_{\delta}$. By the construction of the Heller operator $\Omega$, we can get an $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule isomorphism $\vartheta:M\longrightarrow{{}_{(u,u)}}M$. Now we can define a map
$\varTheta:\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}M\longrightarrow{{}_{(u,u)}}(\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}M),~~~~aj\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}m\mapsto uaja_{u,\delta}^{-1}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}\vartheta(m),$
for any $aj\in\mathbb{A}j$ and any $m\in M$. Then it is routine to check that this map is well-defined and gives an $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule isomorphism. In particular, this Morita equivalence is $\Delta(P_{0})$-equivariant.
Consier the second case where $\mathbb{A}$ is Morita equivalent to the principal block $B_{0}(\mathrm{A}_{5})$ of $\mathrm{A}_{5}$. We have an analogue notation ${{}_{(u,u)}}U^{\prime}$ for any $\mathbb{A}_{\delta}$-$B_{0}(\mathrm{A}_{5})$-bimodule $U^{\prime}$. By [16, Theorem 4.14.2], the $\mathbb{A}_{\delta}$-$B_{0}(\mathrm{A}_{5})$-bimodule $N^{\prime}$, which is an indecomposable direct summand of $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathcal{O}L}B_{0}(\mathrm{A}_{5})$, is unique up to isomorphisms. It is clear that $\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathcal{O}L}B_{0}(\mathrm{A}_{5})$ is isomorphic to $){{}_{(u,u)}}(\mathbb{A}_{\delta}\mathop{\otimes}\limits_{\mathcal{O}L}B_{0}(\mathrm{A}_{5}$)) as $\mathbb{A}_{\delta}$-$B_{0}(\mathrm{A}_{5})$-bimodules. Therefore, we have an $\mathbb{A}_{\delta}$-$B_{0}(\mathrm{A}_{5})$-bimodule isomorphism $\vartheta^{\prime}:N\longrightarrow{{}_{(u,u)}}N$ with the same argument above. Similar to the construction of map $\varTheta$, we can get an $\mathbb{A}_{\delta}$-$B_{0}(\mathrm{A}_{5})$-bimodule isomorphism $\varTheta^{\prime}:\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}N\longrightarrow{{}_{(u,u)}}\mathbb{A}j\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}N$, sending $aj\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}n$ to $uaja_{u,\delta}^{-1}\mathop{\otimes}\limits_{\mathbb{A}_{\delta}}\vartheta^{\prime}(n)$ for any $aj\in\mathbb{A}j$ and any $n\in N$. We are done. ∎
Corollary 6.12.
The actions of $P_{0}$ on $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})$ and $\mathrm{Irr}_{k}(\mathbb{A})$ are trivial.
Proof.
This follows from Lemma 6.11 easily. ∎
6.13.
[…]
6.14.
There are two possibilities on $|\mathrm{Irr}_{k}(A)|$ (see [7, Theorem 1.1]). Let us recall some notation and results from [7]. Suppose that $P_{0}=P$, namely, $D$ is in the center of $P$. Then $N_{G}(P_{\gamma})$ controls fusion of the block $b$ and $|\mathrm{Irr}_{k}(A)|=3=|\mathrm{Irr}_{k}(\mathbb{A})|$. Furthermore, by the proof of Corollary 5.21, we can identify $E$ with the inertial quotient $N_{G}(P_{\gamma})/PC_{G}(P)$ and then $P=D\times C_{P}(E)$. We denote $C_{P}(E)$ by $R$ in this case. The other case is that $P_{0}<P$. In this case, $|\mathrm{Irr}_{k}(A)|=2$ and $|P:P_{0}|=2$. Furthermore, $E$ can be identified with a subgroup of $\mathrm{Aut}(P_{0})$ and then $P_{0}=D\times C_{P_{0}}(E)$ with this identification. We denote $C_{P_{0}}(E)$ by $R_{0}$ in this case.
Now we have the following structures about $\mathrm{Irr}_{\mathcal{K}}(A)$ and $\mathrm{Irr}_{k}(A)$.
Proposition 6.15.
Keep the notation as above. Then exactly one of the two following statements holds.
(i) Suppose that $P_{0}=P$. Then every simple $\hat{\mathbb{A}}-$module and every simple $\bar{\mathbb{A}}$-module are both $P$-stable and can be extended to $\hat{A}$ and $\bar{A}$, respectively. Therefore, we have
$\mathrm{Irr}_{\mathcal{K}}(A)=\{\hat{U}_{\mathrm{i}}\mu\mid\mathrm{i}=1,2,3,4~\mathrm{and}~\mu\in\mathrm{Irr}(R)\}$
and
$\mathrm{Irr}_{k}(A)=\{\bar{W}_{\mathrm{i}}\mid\mathrm{i}=1,2,3\}.$
(ii) Suppose that $P_{0}<P$. Then we can choose notation such that $\bar{W}_{1}$ is $P$-stable and the stabilizer of $\bar{W}_{2}$ in $P$ is $P_{0}$. So $P$ can permute $\bar{W}_{2}$ and $\bar{W}_{3}$. Moreover, $\bar{W}_{1}$ can be extended to $\bar{A}$ and $\bar{W}_{\mathrm{i}}$ can be extended to $\bar{\mathbb{A}}_{P_{0}}$ for $\mathrm{i}=1,2$. For $\mathrm{Irr}_{\mathcal{K}}(A)$, we can obtain that $\hat{U}_{1}$ and $\hat{U}_{4}$ are $P$-stable and the stabilizer of $\hat{U}_{2}$ in $P$ is $P_{0}$. So $P$ can permute $\hat{U}_{2}$ and $\hat{U}_{3}$. Moreover, $\hat{U}_{1}$ and $\hat{U}_{4}$ can be extended to $\hat{A}$ and $\hat{U}_{\mathrm{i}}$ can be extended to $\hat{\mathbb{A}}_{P_{0}}$ for $\mathrm{i}=1,2$. In conclusion, we have
$\mathrm{Irr}_{\mathcal{K}}(A)=\{\hat{U}_{\mathrm{i}}\mu\mid\mathrm{i}=1,4~\mathrm{and}~\mu\in\mathrm{Irr}(P/D)\}\bigcup\{\mathrm{Ind}_{P_{0}}^{P}(\hat{U}_{\mathrm{i}}\mu)\mid\mathrm{i}=2~\mathrm{or}~3~\mathrm{and}~\mu\in\mathrm{Irr}(R_{0})\}$
and
$\mathrm{Irr}_{k}(A)=\{\bar{W}_{1},\mathrm{Ind}_{P_{0}}^{P}(\bar{W}_{2})\cong\mathrm{Ind}_{P_{0}}^{P}(\bar{W}_{3})\}.$
Proof.
For the case (i), since $|\mathrm{Irr}_{k}(A)|=3=|\mathrm{Irr}_{k}(\mathbb{A})|$ and $P$ is a $2$-group, it is easy to see that every simple $\bar{\mathbb{A}}$-module is $P$-stable. Then its extendibility is well known since $P$ is a $2$-group. The description of $\mathrm{Irr}_{k}(A)$ can be deduced from this extendibility. Consider simple $\hat{A}-$modules. Obviously, the decomposition map $d_{\mathbb{A}}$ commutes with the action of $P$. Then by equations (6.10.1) and (6.10.2), we can easily obtain that every simple $\hat{\mathbb{A}}-$module is $P$-stable. By Proposition 6.9, $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\mathrm{i}})$ is coprime to $2$ for any $\mathrm{i}\in\{1,2,3,4\}$. Then by [13, Proposition 2.4], every simple $\hat{\mathbb{A}}-$module can be extended to $\hat{A}$. Hence, the description of $\mathrm{Irr}_{\mathcal{K}}(A)$ follows from Clifford Theory and the fact that $P/D$ is isomorphic to $R$.
For the case (ii), by Corollary 6.12, every simple $\hat{\mathbb{A}}-$module and every simple $\bar{\mathbb{A}}$-module are both $P_{0}$-stable. So the actions of $P$ on these two sets are essentially equivalent to the actions of $P/P_{0}$, which has order $2$. Since in this case $|\mathrm{Irr}_{k}(A)|=2$ and $|\mathrm{Irr}_{k}(\mathbb{A})|=3$, then one of simple $\bar{\mathbb{A}}$-modules is $P$-stable and $P/P_{0}$ permutes the remaining two simple modules. When $\mathbb{A}$ is Morita equivalent to the principal block $B_{0}(\mathrm{A}_{5})$ of $\mathrm{A}_{5}$, by the equations (6.10.2), $\bar{W}_{1}$ has to be the unique $P$-stable simple $\bar{\mathbb{A}}$-module. When $\mathbb{A}$ is Morita equivalent to $\mbox{$\mathcal{O}$}L$, we have to choose notation such that $\bar{W}_{1}$ is $P$-stable since we do not known whether the Morita equivalence in this case is $\Delta{P}$-equivariant. Then $P/P_{0}$ permutes $\bar{W}_{2}$ and $\bar{W}_{3}$. Similarly, we can obtain that $\bar{W}_{1}$ can be extended to $\bar{A}$ and $\bar{W}_{\mathrm{i}}$ can be extended to $\bar{\mathbb{A}}_{P_{0}}$ for $\mathrm{i}=2,3$. Now the description of $\mathrm{Irr}_{k}(A)$ follows from the extendibility and Clifford Theory. Consider simple $\hat{A}-$modules. By equations (6.10.1) and (6.10.2), we can obtain that $\hat{U}_{1}$ and $\hat{U}_{4}$ are $P$-stable and $P/P_{0}$ can permute $\hat{U}_{2}$ and $\hat{U}_{3}$. With the same argument above, $\hat{U}_{1}$ and $\hat{U}_{4}$ can be both extended to $\hat{A}$. Note that the proof of [13, Proposition 2.4] is still valid if we replace $P$ and $\hat{A}$ there by $P_{0}$ and $\hat{\mathbb{A}}_{P_{0}}$. Then $\hat{U}_{2}$ and $\hat{U}_{3}$ can be both extended to $\hat{\mathbb{A}}_{P_{0}}$. Similarly, the description of $\mathrm{Irr}_{\mathcal{K}}(A)$ follows from Clifford Theory and the fact that $P_{0}/D$ is isomorphic to $R_{0}$. ∎
6.2 Cyclic group case
In this subsection, we will assume the hyperfocal subgroup $D$ is a nontrivial cyclic group. In this case the hyperfocal quotient inertial $E_{\mathfrak{h}}$ is a cyclic group of order divides p1p-1 and acts freely on $D-\{1\}$. Then we can apply Theorem 5.14 to this case and get a stable equivalence of Morita type between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$. Here $L=D\rtimes E_{\mathfrak{h}}$. It is well-known that this stable equivalence of Morita type can be lifted to a Rickard equivalence by adopting the method in the proof of [17, Theorem 11.12.1]. In the following, we will investigate the equivariant property of this Rickard equivalence.
6.16.
Since the hyperfocal subgroup $D$ is cyclic, by [33, Theorem 3], $N_{G}(P_{\gamma})$ controls fusion of the block $b$. So we can identify $E_{\mathfrak{h}}$ with the inertial quotient $E=N_{G}(P_{\gamma})/PC_{G}(P)$ and then $P=D\rtimes C_{P}(E)$. We denote $C_{P}(E)$ by $R$. Then the conjutation action of $R$ on $E$ is trivial. Moreover, by [33, Theorem 1], $l(b)=|\mathrm{IBr}(b)|=|E|$. Then $l(A)$ also equals $|E|$. On the other hand, since $A$ is symmetric and stably equivalent to $\mbox{$\mathcal{O}$}L$, which is a serial algebra in this case, by [17, Proposition 11.6.1], $l(\mathbb{A})=|\mathrm{IBr}(L)|$, which is just $|E|$. Hence, we have $l(\mathbb{A})=l(A)$ equal to the inertial index $|E|$ of the block. It can be deduced from this equality that $P$ acts trivially on $\mathrm{Irr}_{k}(\mathbb{A})$ and then every simple $\bar{\mathbb{A}}$-module can be extended to $\bar{A}$.
6.17.
Let $\mathfrak{i}$ and $\varsigma$ be primitive idempotents of $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$, respectively. For any $u\in P$, we have the twisted $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule ${{}_{(u,u)}}(\mathbb{A}\mathfrak{i}\mathop{\otimes}\limits_{\mathcal{O}}\varsigma\mbox{$\mathcal{O}$}L)$. It is clear that the map sending $\mathfrak{a}\mathfrak{i}^{u}\mathop{\otimes}\limits_{\mathcal{O}}\varsigma^{u}x$ to ${{}^{u}}\mathfrak{a}\mathfrak{i}\mathop{\otimes}\limits_{\mathcal{O}}\varsigma{{}^{u}}x$ defines an isomorphism of $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules from $\mathbb{A}\mathfrak{i}^{u}\mathop{\otimes}\limits_{\mathcal{O}}\varsigma^{u}\mbox{$\mathcal{O}$}L$ to ${{}_{(u,u)}}(\mathbb{A}\mathfrak{i}\mathop{\otimes}\limits_{\mathcal{O}}\varsigma\mbox{$\mathcal{O}$}L)$, where $\mathfrak{a}\in\mathbb{A}$ and $x\in L$. By the paragraph above, $P$ acts trivially on $\mathrm{Irr}_{k}(\mathbb{A})$. It is clear that $\mathbb{A}\mathfrak{i}^{u}$ is isomorphic to $\mathbb{A}\mathfrak{i}$ as left $\mathbb{A}$-modules. Similarly, $P$ acts on $\mathrm{IBr}(L)$ trivially and then $\varsigma^{u}\mbox{$\mathcal{O}$}L$ is isomorphic to $\varsigma\mbox{$\mathcal{O}$}L$ as right $\mbox{$\mathcal{O}$}L$-modules. In conclusion, ${{}_{(u,u)}}(\mathbb{A}\mathfrak{i}\mathop{\otimes}\limits_{\mathcal{O}}\varsigma\mbox{$\mathcal{O}$}L)$ is isomorphic to $\mathbb{A}\mathfrak{i}\mathop{\otimes}\limits_{\mathcal{O}}\varsigma\mbox{$\mathcal{O}$}L$ for any $u\in P$. Hence, we have the following equivariant property for projective $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules.
Lemma 6.18.
Every finitely generated projetive $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule is $\Delta(P)$-stable. Namely, if $\mathbb{P}$ is a finitely generated projective $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule and $u$ belongs to $P$, then $_{(u,u)}$$\mathbb{P}$ is isomorphic to $\mathbb{P}$ as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules.
6.19.
Let $M_{\cdot}=(M_{n},\partial_{n})$ be a bounded complex of $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules. For any $u\in P$, we can define the twisted complex ${{}_{(u,u)}}(M_{\cdot})$ with ${{}_{(u,u)}}(M_{\cdot})_{n}={{}_{(u,u)}}(M_{n})$ and the morphism from ${{}_{(u,u)}}(M_{\cdot})_{n}$ to ${{}_{(u,u)}}(M_{\cdot})_{n-1}$ is just $\partial_{n}$. The complex $M_{\cdot}$ is called $P$-invariant if $M_{\cdot}$ is isomorphic to ${{}_{(u,u)}}(M_{\cdot})$ in $\mathcal{D}^{b}(\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ})$. Furthermore, if the complex $M_{\cdot}$ induces a derived equivalence between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$, then this derived equivalence is called $P$-equivariant. We refer to [18, Definition 2.5] for these two definitions.
Theorem 6.20.
There is a $P$-invariant Rickard $2$-term complex $\mathbb{M}_{\cdot}$ of $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules. In particular, there is a $P$-equivariant Rickard equivalence between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$.
Proof.
Recall from the paragraph 5.16 that we have an indecomposable $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule $M_{\mathrm{st}}$ that induces a stable equivalence of Morita type between $\mathbb{A}$ and $\mbox{$\mathcal{O}$}L$. Moreover, since the hyperfocal subgroup $D$ is cyclic, $N_{G}(P_{\gamma})$ controls fusion of the block $b$ (see the paragraph 6.16). By Corollary 5.21, the two modules $M_{\mathrm{st}}$ and ${{}_{(u,u)}}M_{\mathrm{st}}$ are isomorphic to each other as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules for any $u\in P$. Fix an isomoprhism $\varphi_{\mathrm{st}}$ between these two bimodules.
Let $\mathbb{P}$ be a projective cover of the bimodule $M_{\mathrm{st}}$ as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules and $\pi$ a corresponding surjective homomorphism from $\mathbb{P}$ to $M_{\mathrm{st}}$. By [10, Theorem 1.1], the hyperfocal subalgebra $\mathbb{A}$ is symmetric. Then by a slight generalization of the proof of [17, Theorem 11.12.1], there is a direct summand $\mathbb{Q}$ of $\mathbb{P}$, uniquely determined by $M_{\mathrm{st}}$ up to isomorphism, such that the following $2$-term complex $\mathbb{M}_{\cdot}$ of $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules
$\mathbb{Q}\ignorespaces\ignorespaces\ignorespaces\ignorespaces$$\scriptstyle{\mathrm{Res}^{\mathbb{P}}_{\mathbb{Q}}(\pi)}$${M_{\mathrm{st}}}$
is a Rickard complex of $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules (see the proof of [9, Theorem 5.24]). For simplicity, we still denote $\mathrm{Res}^{\mathbb{P}}_{\mathbb{Q}}(\pi)$ by $\pi$. Obviously, $\pi:{{}_{(u,u)}}\mathbb{P}\rightarrow{{}_{(u,u)}}M_{\mathrm{st}}$ is still a surjective homomorphism of $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules and then $_{(u,u)}$$\mathbb{P}$ is a projective cover of ${{}_{(u,u)}}M_{\mathrm{st}}$. Due to $M_{\mathrm{st}}$ isomorphic to ${{}_{(u,u)}}M_{\mathrm{st}}$ and the uniqueness of projective covers, there is an isomorphism $\psi_{\mathrm{st}}$ of $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules between $\mathbb{P}$ and $_{(u,u)}$$\mathbb{P}$ such that the following diagram commutes
[…]
By the construction of $\mathbb{P}$ (see [16, Theorem 11.9.1]), it is clear that every indecomposable direct summand of $\mathbb{P}$ has multiplicity one and $\mathbb{P}$ is isomorphic to $(\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ})\mathfrak{f}$ as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules for some idempotent $\mathfrak{f}$ of $\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ}$. We can write $_{(u,u)}$$\mathbb{P}$ as $_{(u,u)}$$\mathbb{Q}\oplus\mathbb{Q}^{\prime}$ for some projective $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodule $\mathbb{Q}^{\prime}$. By Lemma 6.18, $\mathbb{Q}$ is isomorphic to $_{(u,u)}$$\mathbb{Q}$ as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules. Then we have $\psi_{\mathrm{st}}(\mathbb{Q})$ isomorphic to $_{(u,u)}$$\mathbb{Q}$ as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules. Therefore, by [29, Lemma 44.7], we have ${{}_{(u,u)}}\mathbb{P}=\psi_{\mathrm{st}}(\mathbb{Q})\oplus\mathbb{Q}^{\prime}$ as $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules. We can assume that ${{}_{(u,u)}}\mathbb{P}=(\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ})\mathfrak{f}$. Then the two decompositions ${{}_{(u,u)}}\mathbb{P}={{}_{(u,u)}}\mathbb{Q}\oplus\mathbb{Q}^{\prime}$ and ${{}_{(u,u)}}\mathbb{P}=\psi_{\mathrm{st}}(\mathbb{Q})\oplus\mathbb{Q}^{\prime}$ yields two orthogonal decompositions of the idempotent $\mathfrak{i}$ as follows
$\mathfrak{f}=\mathfrak{f}_{u}+\mathfrak{f}^{\prime}_{u}~\mathrm{and}~\mathfrak{f}=\mathfrak{f}_{\mathrm{st}}+\mathfrak{f}^{\prime}_{\mathrm{st}}$
with ${{}_{(u,u)}}\mathbb{Q}=(\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ})\mathfrak{f}_{u}$ and $\psi_{\mathrm{st}}(\mathbb{Q})=(\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ})\mathfrak{f}_{\mathrm{st}}$ and $(\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ})\mathfrak{f}_{u}^{\prime}=\mathbb{Q}^{\prime}=(\mathbb{A}\mathop{\otimes}\limits_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)^{\circ})\mathfrak{f}_{\mathrm{st}}^{\prime}$. It is easy to check that $\mathfrak{f}_{u}\mathfrak{f}_{\mathrm{st}}^{\prime}=0=\mathfrak{f}_{\mathrm{st}}^{\prime}\mathfrak{f}_{u}$ and $\mathfrak{f}_{\mathrm{st}}\mathfrak{f}_{u}^{\prime}=0=\mathfrak{f}_{u}^{\prime}\mathfrak{f}_{\mathrm{st}}$. Then
$\mathfrak{f}_{u}=\mathfrak{f}\mathfrak{f}_{u}=\mathfrak{f}_{\mathrm{st}}\mathfrak{f}_{u}=\mathfrak{f}_{\mathrm{st}}(\mathfrak{f}_{u}+\mathfrak{f}_{u}^{\prime})=\mathfrak{f}_{\mathrm{st}}.$
So we have $\psi_{\mathrm{st}}(\mathbb{Q})={{}_{(u,u)}}\mathbb{Q}$. By restricting from $\mathbb{P}$ to $\mathbb{Q}$, we can deduce from the commutative diagram (6.20.6) the following diagram commuting
[…]
which implies that the complex $\mathbb{M}_{\cdot}$ is isomorphic to ${{}_{(u,u)}}\mathbb{M}_{\cdot}$ as complexes of $\mathbb{A}$-$\mbox{$\mathcal{O}$}L$-bimodules. In particular, the complex $\mathbb{M}_{\cdot}$ is $P$-invariant. We are done. ∎
6.21.
[…]
$\mathrm{End}_{\mathcal{D}^{b}(\hat{\mathbb{A}})}(\hat{U})\cong\mathrm{End}_{\mathcal{D}^{b}(\hat{\mathbb{A}})}(\hat{U}[m])\cong\mathrm{End}_{\mathcal{D}^{b}(\mathcal{K}L)}(\Xi)$
as $\mathcal{K}$-algebras. But it is well-known that
$\mathrm{End}_{\mathcal{D}^{b}(\hat{\mathbb{A}})}(\hat{U})\cong\mathrm{End}_{\hat{\mathbb{A}}}(\hat{U})~\mathrm{and}~\mathrm{End}_{\mathcal{D}^{b}(\mathcal{K}L)}(\Xi)\cong\mathrm{End}_{\mathcal{K}L}(\Xi)$
as $\mathcal{K}$-algebras. In conclusion, we can get $\mathrm{End}_{\hat{\mathbb{A}}}(\hat{U})\cong\mathrm{End}_{\mathcal{K}L}(\Xi)$, which is just $\mathcal{K}$. This implies that the field $\mathcal{K}$ is a splitting field for $\hat{\mathbb{A}}$. Therefore, the set $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})$ is an orthonormal basis of $R_{\mathcal{K}}(\mathbb{A})$.
6.22.
Next, we will investigate the structures of $\mathrm{Irr}_{\mathcal{K}}(A)$ and $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})$. Set $|D|=p^{n}$ with $n\geq 1$ and $|E|=e$. Clearly, $e$ divides p1p-1. By Theorem 6.20, $\mathrm{k}(\mathbb{A})=|\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})|=|\mathrm{Irr}(L)|=e+\frac{p^{n}-1}{e}$ and $l(\mathbb{A})=|\mathrm{Irr}_{k}(\mathbb{A})|=|\mathrm{IBr}(L)|=e$. Denote by $\mathcal{M}$ a set of representatives of the $E$-conjugacy classes of nontrivial irreducible characters of $D$. Clearly, $|\mathcal{M}|=\frac{p^{n}-1}{e}$. We can identify $\mathrm{Irr}(E)$ with a subset of $\mathrm{Irr}(L)$ through the canonical way. Then $\mathrm{Irr}(L)=\{\mathrm{Ind}_{D}^{L}(\chi)\mid\chi\in\mathcal{M}\}\cup\{\lambda\mid\lambda\in\mathrm{Irr}(E)\}$. So we can set $\{\hat{U}_{\chi},\hat{U}_{\lambda}\mid\chi\in\mathcal{M},\lambda\in\mathrm{Irr}(E)\}$ to be a set of representatives of the isomorphism classes of simple $\hat{\mathbb{A}}-$modules and $\{\bar{W}_{\lambda}\mid\lambda\in\mathrm{Irr}(E)\}$ to be a set of representatives of the isomorphism classes of simple $\bar{\mathbb{A}}$-modules. By [12, Proposition 3.2], the decomposition map $d_{\mathbb{A}}$ is surjective. At the same time, by Corollary 5.15, the Cartan matrix of $\mathbb{A}$ is non-singular. Therefore, by [12, Proposition 2.2], we have $L^{0}(\mathbb{A})^{\bot}=\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})$. For the group algebra $\mbox{$\mathcal{O}$}L$, we denote by $R_{\mathcal{K}}(\mbox{$\mathcal{O}$}L)$ the set of all generalized characters of $L$. We adopt the similar notation for $R_{k}(\mbox{$\mathcal{O}$}L)$.
6.23.
By Theorem 6.20, the Rickard equivalence induced by the complex $\mathbb{M}_{\cdot}$ can determine two bijections $\Phi_{\mathcal{K}}:R_{\mathcal{K}}(\mathbb{A})\rightarrow R_{\mathcal{K}}(\mbox{$\mathcal{O}$}L)$ and $\Phi_{k}:R_{k}(\mathbb{A})\rightarrow R_{k}(\mbox{$\mathcal{O}$}L)$ such that $\Phi_{\mathcal{K}}$ is an isometry and the following the diagram commutes
${R_{\mathcal{K}}(\mathbb{A})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{\mathcal{K}}}$$\scriptstyle{d_{\mathbb{A}}}$${R_{\mathcal{K}}(\mbox{$\mathcal{O}$}L)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{d_{\mathcal{O}L}}$${R_{k}(\mathbb{A})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{k}}$${R_{k}(\mbox{$\mathcal{O}$}L).}$ (6.23.5)
Then the restricting of $\Phi_{\mathcal{K}}$ to $L^{0}(\mathbb{A})$ induces a bijective isometry, still denoted by $\Phi_{\mathcal{K}}$ between $L^{0}(\mathbb{A})$ and $L^{0}(\mbox{$\mathcal{O}$}L)$. Furthermore, since $\Phi_{\mathcal{K}}$ is an isometry, we can get an bijective isometry $\Phi_{\mathcal{K}}:\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})\rightarrow\mathrm{Pr}_{\mathcal{O}}(\mbox{$\mathcal{O}$}L)$. Then we can obtain $\mathbb{Z}$-bases of $\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})$ and $L^{0}(\mathbb{A})$ as follows.
Proposition 6.24.
For any $\chi\in\mathcal{M}$ and any $\lambda\in\mathrm{Irr}(E)$, there are signs $\epsilon_{\chi}\in\{\pm 1\}$ and $\epsilon_{\lambda}\in\{\pm 1\}$ such that the following holds.
(i) $\Phi_{\mathcal{K}}([\hat{U}_{\chi}])=\epsilon_{\chi}\chi$ and $\Phi_{\mathcal{K}}([\hat{U}_{\lambda}])=\epsilon_{\lambda}\lambda$.
(ii) $\{\epsilon_{\chi}[\hat{U}_{\chi}]-\sum\limits_{\lambda\in\mathrm{Irr}(E)}\epsilon_{\lambda}[\hat{U}_{\lambda}]\mid\chi\in\mathcal{M}\}$ is a $\mathbb{Z}$-basis of $L^{0}(\mathbb{A})$. In particular, we have $\epsilon_{\chi}=\epsilon_{\chi^{\prime}}$, denoted by $\epsilon$ for any $\chi,\chi^{\prime}\in\mathcal{M}$.
(iii) $\{\epsilon_{\lambda}[\hat{U}_{\lambda}]+\sum\limits_{\chi\in\mathcal{M}}\epsilon_{\chi}[\hat{U}_{\chi}]\mid\lambda\in\mathrm{Irr}(E)\}$ is a $\mathbb{Z}$-basis of $\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})$.
Proof.
The statements in this proposition can be easily checked from the arguments in the paragraph above except the one $\epsilon_{\chi}=\epsilon_{\chi^{\prime}}$. We can assume that $|\mathcal{M}|\geq 2$. Take two distinct elements $\chi$ and $\chi^{\prime}$ of $\mathcal{M}$. Then $\epsilon_{\chi}[\hat{U}_{\chi}]-\epsilon_{\chi^{\prime}}[\hat{U}_{\chi^{\prime}}]$ belongs to $L^{0}(\mathbb{A})$. By the lemma below, we have $\epsilon_{\chi}\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\chi})-\epsilon_{\chi^{\prime}}\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\chi^{\prime}})=0$, which forces $\epsilon_{\chi}=\epsilon_{\chi^{\prime}}$. ∎
Lemma 6.25.
Let $\sum\limits_{\chi\in\mathcal{M},\lambda\in\mathrm{Irr}(E)}(m_{\chi}[\hat{U}_{\chi}]+m_{\lambda}[\hat{U}_{\lambda}])$ be an element of $L^{0}(\mathbb{A})$. Here, $m_{\chi}$ and $m_{\lambda}$ are integers. We have $\sum\limits_{\chi\in\mathcal{M},\lambda\in\mathrm{Irr}(E)}(m_{\chi}\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\chi})+m_{\lambda}\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\lambda}))=0$.
Proof.
It can be easily deduced from the well-known fact that $\mathrm{dim}_{\mathcal{K}}(\mathrm{Hom}_{\hat{\mathbb{A}}}(\hat{\mathbb{A}}\mathfrak{i},X))=\mathrm{dim}_{\mathcal{K}}(\mathfrak{i}X)$ for any primitive idempotent $\mathfrak{i}$ of $\mathbb{A}$ and any finite dimensional $\hat{\mathbb{A}}-$module $X$. ∎
6.26.
Now we can verify the forward direction of the KLN conjecture when the hyperfocal subgroup $D$ is nontrivial cyclic. We first recall a well-known fact that there is a simple $\hat{A}-$module of dimension coprime to $p$. Indeed, by [17, Proposition 6.11.11], the $p$-part of the dimension of a simple $\hat{A}-$module equals to the height of its corresponding irreducible ordinary character in the block $b$ through the Morita equivalence induced by the bimodule $i\mbox{$\mathcal{O}$}G$.
Lemma 6.27.
There is at least one simple $\hat{\mathbb{A}}-$module of dimension coprime to $p$.
Proof.
This can be easily deduced from the argument above and [13, Proposition 2.3 (iv)]. ∎
Proposition 6.28.
The dimension of every simple $\hat{\mathbb{A}}-$module is coprime to $p$ when the hyperfocal subgroup $D$ is nontrivial cyclic. In particular, the forward direction of the KLN conjecture is true in this case.
Proof.
Since the hyperfocal subgroup $D$ is assumed to be nontrivial, the order $e$ of $E$ can not be $1$. Then by Proposition 6.24 (iii), $\epsilon_{\lambda}[\hat{U}_{\lambda}]-\epsilon_{\mu}[\hat{U}_{\mu}]$ belongs to $\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})$ for any two distinct elements $\lambda$ and $\mu$ of $\mathrm{Irr}(E)$. Since $\mathbb{A}$ is projective as left $\mbox{$\mathcal{O}$}D$-module, we have $p$ dividing $\epsilon_{\lambda}\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\lambda})-\epsilon_{\mu}\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\mu})$. In particular, $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\lambda})$ is coprime to $p$ if and only if $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\mu})$ is coprime to $p$. Suppose that $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\lambda})$ is divided by $p$ for some $\lambda\in\mathrm{Irr}(E)$. So is $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\mu})$ for any $\mu\in\mathrm{Irr}(E)$. Now by Proposition 6.24 (ii) and Lemma 6.25, $\epsilon\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\chi})-\sum\limits_{\lambda\in\mathrm{Irr}(E)}\epsilon_{\lambda}\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\lambda})=0$. Hence, $\epsilon\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\chi})$ is divided by $p$ for any $\chi\in\mathcal{M}$. In conclusion, every simple $\hat{\mathbb{A}}-$module has dimension divided by $p$, which contradicts Lemma 6.27. So $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\lambda})$ is coprime to $p$ for any $\lambda\in\mathrm{Irr}(E)$. By Proposition 6.24 (iii), there is at least one element $\chi$ of $\mathcal{M}$ such that $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\chi})$ is coprime to $p$. We denote it by $\chi_{0}$. Now we can assume that $\mathcal{M}$ has at least two elements. For any other $\chi\in\mathcal{M}$ different from $\chi_{0}$, by Proposition 6.24 (ii), $\epsilon([\hat{U}_{\chi_{0}}]-[\hat{U}_{\chi}])$ belongs to $L^{0}(\mathbb{A})$. In particular, by Lemma 6.25, these two simple modules $\hat{U}_{\chi_{0}}$ and $\hat{U}_{\chi}$ has the same dimension. Then we are done. ∎
6.29.
Note that the structure of a $\mathbb{Z}$-basis of $\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})$ in Proposition 6.24 plays an important role in the proof of Proposition 6.28. Furthermore, by this structure, we can also calculate decomposition numbers of $\mathbb{A}$ and get that these numbers are either $1$ or 0 in the following. This result can be regarded as a ‘hyperfocal decomposition number’ version of the classical fact that decomposition numbers of blocks with cyclic defect groups are either $1$ or 0. By [12, §2], this is equivalent to showing that $\hat{\mathbb{A}}\mathfrak{i}]=\sum\limits_{\theta\in\mathrm{Irr}(L)}a_{\theta}[\hat{U}_{\theta}]$ with $a_{\theta}$ equal to $1$ or 0 for any primitive idempotent $\mathfrak{i}$ of $\mathbb{A}$. One key technique in the context of blocks with cyclic defect groups is making use of $\mathcal{O}$-pure submodules. An $\mathbb{A}$-submodule $U^{\prime}$ of an $\mathbb{A}$-module $U$ is $\mathcal{O}$-pure in $U$ if it is a direct summand of $U$ as an $\mathcal{O}$-module (see [16, Section 4.2]). Our proof is similar to the one of [17, Theorem 11.10.5]. We need the following two lemmas which are analogues of [17, Lemma 11.10.3] and [17, Proposition 11.10.4].
Lemma 6.30.
Let $U$ be a finitely generated $\mathcal{O}$-free $\mathbb{A}$-module such that $\bar{U}=k\mathop{\otimes}\limits_{\mathcal{O}}U$ is nonprojective indecomposable. Then the image $\hat{U}$] of $\hat{U}=\mathcal{K}\mathop{\otimes}\limits_{\mathcal{O}}U$ in $R_{\mathcal{K}}(\mathbb{A})$ is not contained in $\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})$.
Proof.
The key ingredient in the proof of [17, Lemma 11.10.3] is the existence of a stable equivalence induced by induction and restriction. Recall from Section $5$ that there is an $\mathcal{O}$-algebra $\mathbb{A}^{\prime}$ Morita equivalent to $\mathbb{A}$ such that it has a subalgebra isomorphic to $\mbox{$\mathcal{O}$}L$ and a stable equivalence of Morita type between $\mathrm{A}^{\prime}$ and this subalgebra induced by induction and restriction exists. Then the proof of [17, Lemma 11.10.3] can be applied to our case. ∎
Lemma 6.31.
Let $U$ be a projective indecomposable $\mathbb{A}$-module. Denote $\mathcal{K}\mathop{\otimes}\limits_{\mathcal{O}}U$ by $\hat{U}$. Suppose that there are nonzero finite dimensional $\hat{\mathbb{A}}-$modules $\hat{U}_{1}$ and $\hat{U}_{2}$ such that $\hat{U}]=[\hat{U}_{1}]+[\hat{U}_{2}]$. Then none of $\hat{U}_{1}]$ or $\hat{U}_{2}]$ is contained in $\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})$.
Proof.
Since $\hat{\mathbb{A}}$ is semi-simple, we have $\hat{U}\cong\hat{U}_{1}\oplus\hat{U}_{2}$ as $\hat{\mathbb{A}}-$modules. Then by [16, Theorem 4.16.4], there is an $\mathcal{O}$-pure submodule $U_{1}$ of $U$ such that $\hat{U}_{1}$ is isomorphic to $\mathcal{K}\mathop{\otimes}\limits_{\mathcal{O}}U_{1}$. By [16, Proposition 4.2.6], $k\mathop{\otimes}\limits_{\mathcal{O}}U_{1}$ can be viewed as a submodule of $k\mathop{\otimes}\limits_{\mathcal{O}}U$ which is a projective indecomposable $\bar{\mathbb{A}}$-module. Since $\mathbb{A}$ is a symmetric $\mathcal{O}$-algebra, the socle of $k\mathop{\otimes}\limits_{\mathcal{O}}U$ is simple. So is the socle of $k\mathop{\otimes}\limits_{\mathcal{O}}U_{1}$. In particular, $k\mathop{\otimes}\limits_{\mathcal{O}}U_{1}$ is indecomposable. By Lemma 6.30, $\hat{U}_{1}]$ is not contained in $\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})$. ∎
Proposition 6.32.
With the notation as in Proposition 6.24. Let $U$ be a projective indecomposable $\mathbb{A}$-module. Then either $\hat{U}]=[\hat{U}_{\lambda}]+[\hat{U}_{\lambda^{\prime}}]$ for some $\lambda,\lambda^{\prime}\in\mathrm{Irr}(E)$ such that $\epsilon_{\lambda}\neq\epsilon_{\lambda^{\prime}}$, or $\hat{U}]=[\hat{U}_{\lambda}]+\sum\limits_{\chi\in\mathcal{M}}[\hat{U}_{\chi}]$ for some $\lambda\in\mathrm{Irr}(E)$ such that $\epsilon_{\lambda}=\epsilon$. In particular, all decomposition numbers of $\mathbb{A}$ are either $1$ or 0.
Proof.
Due to the structure of $\mathrm{Pr}_{\mathcal{O}}(\mathbb{A})$ described in Proposition 6.24 (iii), by Lemmas 6.30 and 6.31, the proof is similar to the one of [17, Theorem 11.10.5]. ∎
6.33.
[…]
Proposition 6.34.
Keep the notation as above. The following statements hold.
(i) For any $\chi\in\mathcal{M}_{R}$, the stabilizer of $\hat{U}_{\chi}$ under the conjugation action of $P$ is just $R_{\chi}$. So $\hat{U}_{\chi}$ is $P$-stable if and only if $\chi$ is $P$-stable. Moreover, for any $\chi\in\mathcal{M}_{R}$, the simple $\hat{\mathbb{A}}-$module $\hat{U}_{\chi}$ can be extended to a simple $\hat{\mathbb{A}}_{R_{\chi}}$-module, denoted by $\hat{U}_{\chi}^{\mathrm{ext}}$.
(ii) For any $\lambda\in\mathrm{Irr}(E)$, the simple $\hat{\mathbb{A}}-$module $\hat{U}_{\lambda}$ is $P$-stable and then it can be extended to a simple $\hat{A}-$module, denoted by $\hat{U}_{\lambda}^{\mathrm{ext}}$.
(iii) $\mathrm{Irr}_{\mathcal{K}}(A)=\{\mathrm{Ind}_{R_{\chi}}^{R}(\hat{U}_{\chi}^{\mathrm{ext}}\mu_{\chi})\mid\chi\in\mathcal{M}_{R},\mu_{\chi}\in\mathrm{Irr}(R_{\chi})\}\bigcup\{\hat{U}_{\lambda}^{\mathrm{ext}}\mu\mid\lambda\in\mathrm{Irr}(E),\mu\in\mathrm{Irr}(R)\}$.
Proof.
By Theorem 6.20 and [18, Theorem 2.6], the isometry bijection $\Phi_{\mathcal{K}}$ in the commutative diagram (6.23.5) preserves the actions of $P$ on $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})$ and $\mathrm{Irr}(L)$. Then the statements on stabilizers of simple $\hat{\mathbb{A}}-$modules are clear. For any $\chi\in\mathcal{M}_{R}$, since $\mathrm{dim}_{\mathcal{K}}(\hat{U}_{\chi})$ is coprime to $p$, the same argument in the proof of Proposition 6.15 shows that $\hat{U}_{\chi}$ can be extended to $\hat{\mathbb{A}}_{R_{\chi}}$. For any $\lambda\in\mathrm{Irr}(E)$, since $\hat{U}_{\lambda}$ is $P$-stable and has dimension coprime to $p$, by [13, Proposition 2.4], $\hat{U}_{\lambda}$ can be extended to $\hat{A}$. This shows (i) and (ii). The statement (iii) just follows from [6, Theorem 1.6]. ∎
As a conesquence, we can get the following decomposition numbers of the block $b$ with nontrivial cyclic hyperfocal subgroup.
Corollary 6.35.
With the notation as above, decomposition numbers of the block $b$ are either $p^{l}$ or 0 with $1\leq p^{l}\leq|R/C_{R}(D)|$ when the hyperfocal subgroup $D$ is nontrivial cyclic. In particular, when the hyperfocal subgroup $D$ is nontrivial cyclic and central in $P$, decomposition numbers of the block $b$ are either $1$ or 0.
Proof.
Since the block algebra $\mbox{$\mathcal{O}$}Gb$ is Morita equivalent to the source algebra $A$, it suffices to show that decomposition numbers of the source algebra $A$ satisfy this property. As before, we can get these numbers by calculating the image $\hat{U}_{A}]$ of $\hat{U}_{A}$ in $R_{\mathcal{K}}(A)$ of any indecomposable projective $A$-module $U_{A}$ such that $\hat{U}_{A}=\mathcal{K}\mathop{\otimes}\limits_{\mathcal{O}}U_{A}$. Since every simple $\bar{\mathbb{A}}$-module can be extended to $\bar{A}$ (see the paragraph 6.16), then every indecomposable projective $A$-module has the form $A\mathfrak{i}$, which is isomorphic to $A\mathop{\otimes}\limits_{\mathbb{A}}\mathbb{A}\mathfrak{i}$ for some primitive idempotent $\mathfrak{i}$ of $\mathbb{A}$. Now fix a primitive idempotent $\mathfrak{i}$ of $\mathbb{A}$. By Proposition 6.32, we can assume that as $\hat{\mathbb{A}}-$modules, $\hat{\mathbb{A}}\mathfrak{i}\cong\hat{U}_{\lambda}\oplus\hat{U}_{\lambda^{\prime}}$ for some suitable $\lambda,\lambda^{\prime}\in\mathrm{Irr}(E)$ or $\hat{\mathbb{A}}\mathfrak{i}\cong\hat{U}_{\mu}\oplus(\mathop{\bigoplus}\limits_{\chi\in\mathcal{M}}\hat{U}_{\chi})$ for some suitable $\mu\in\mathrm{Irr}(E)$. Fix these notation. Then as $\hat{A}-$modules, $\hat{A}\mathfrak{i}\cong\hat{A}\mathop{\otimes}\limits_{\hat{\mathbb{A}}}\mathbb{A}\mathfrak{i}\cong\mathrm{Ind}_{1}^{R}(\hat{U}_{\lambda})\oplus\mathrm{Ind}_{1}^{R}(\hat{U}_{\lambda^{\prime}})$ or $\hat{A}\mathfrak{i}\cong\hat{A}\mathop{\otimes}\limits_{\hat{\mathbb{A}}}\mathbb{A}\mathfrak{i}\cong\mathrm{Ind}_{1}^{R}(\hat{U}_{\mu})\oplus(\mathop{\bigoplus}\limits_{\chi\in\mathcal{M}}\mathrm{Ind}_{1}^{R}(\hat{U}_{\chi}))$. By Proposition 6.34, as $\hat{A}-$modules, $\mathrm{Ind}_{1}^{R}(\hat{U}_{\nu})\cong\mathop{\bigoplus}\limits_{\vartheta\in\mathrm{Irr}(R)}\hat{U}_{\nu}^{\mathrm{ext}}\vartheta$ for any $\nu\in\mathrm{Irr}(E)$ and $\mathrm{Ind}_{1}^{R}(\hat{U}_{\chi})\cong\mathop{\bigoplus}\limits_{\mu_{\chi}\in\mathrm{Irr}(R_{\chi})}\mathrm{Ind}_{R_{\chi}}^{R}(\hat{U}_{\chi}^{\mathrm{ext}}\mu_{\chi})$ for any $\chi\in\mathcal{M}$. When two elements $\chi$ and $\chi^{\prime}$ of $\mathcal{M}$ belong to the same $R$-orbit, the two $\hat{A}-$modules $\mathrm{Ind}_{1}^{R}(\hat{U}_{\chi})$ and $\mathrm{Ind}_{1}^{R}(\hat{U}_{\chi^{\prime}})$ are isomorphic to each other. Since $C_{R}(D)$ acts trivially on $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})$, it is easy to get the information on decomposition numbers of $A$ as stated in this corollary. ∎